### v1

# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

Abstract
Neighborhood smoothing methods achieve minimax-optimal rates for estimating edge probabilities under graphon models, but their use for statistical inference has remained limited. The main obstacle is that classical neighborhood smoothers select data-driven neighborhoods and average edges using the same adjacency matrix, inducing complex dependencies that invalidate standard concentration and normal approximation arguments.
We introduce a leave-one-out modification of neighborhood smoothing for undirected simple graphs. When estimating a single entry $P_{ij}$, the neighborhood of node $i$ is constructed from an adjacency matrix in which the $j$th row and column are set to zero, thereby decoupling neighborhood selection from the edges being averaged. We show that this construction restores conditional independence of the centered summands, enabling the use of classical probabilistic tools for inference.
Under piecewise Lipschitz graphon assumptions and logarithmic degree growth, we derive variance-adaptive concentration inequalities based on Bousquet’s inequality and establish Berry-Esseen bounds with explicit rates for the normalized estimation error. These results yield both finite-sample and asymptotic confidence intervals for individual edge probabilities. The same leave-one-out structure also supports an honest cross-validation scheme for tuning parameter selection, for which we prove an oracle inequality. The proposed estimator retains the optimal row-wise mean-squared error rates of classical neighborhood smoothing while providing valid entrywise uncertainty quantification.
1 Introduction
Graphon models provide a canonical framework for nonparametric analysis of network data, representing latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2008; Lovász, 2012). Over the past decade, a substantial literature has established sharp minimax rates for estimating the probability matrix $P$ or the underlying graphon $h$ from a single observed adjacency matrix, under a variety of smoothness and structural assumptions (Bickel et al., 2011; Wolfe et al., 2013; Gao et al., 2015; Chan et al., 2014). From the perspective of point estimation, the theory is now mature.
What remains largely unresolved is statistical inference for individual edge probabilities. Given an estimator $\widehat{P}_{ij}$, it is unclear how to attach a valid measure of uncertainty-such as a confidence interval-to this estimate under nonparametric graphon models. Existing inferential theory for networks has focused almost exclusively on low-dimensional or global functionals, including eigenvalues, eigenvectors, and spectral embeddings (Lei and Rinaldo, 2015; Athreya et al., 2018; Tang et al., 2017; Capitaine et al., 2019). Entrywise inference for $P_{ij}$, despite its natural interpretation and relevance for downstream tasks, has received comparatively little attention.
This gap is not merely technical. It is structural.
Neighborhood smoothing estimators, such as those of Zhang et al. (2017), achieve optimal mean-squared error by averaging observed edges over data-driven neighborhoods of similar nodes. However, these procedures reuse the adjacency matrix twice: first to construct neighborhoods, and then to average edges within those neighborhoods. As a consequence, the random neighborhood weights depend on the very edges being averaged. Even after conditioning on the underlying probability matrix, the centered summands in the estimator are no longer independent. This dependence invalidates the assumptions underlying classical concentration inequalities and central limit theorems, and it is the fundamental reason why standard inferential tools fail for neighborhood smoothers.
This paper shows that this obstruction can be removed by a simple but principled modification.
We introduce leave-one-out (LOO) neighborhood smoothing. When estimating a single entry $P_{ij}$, we construct the neighborhood of node $i$ using a modified adjacency matrix in which the $j$th row and column are set to zero. Distances and neighborhood selection are computed using this modified matrix, after which the estimator averages the original edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$. By construction, conditional on the probability matrix, the random neighborhood $\mathcal{N}_{i}^{(-j)}$ is independent of the edges used to form the estimate.
This decoupling is the central insight of the paper. It restores conditional independence of the centered summands and makes classical probabilistic tools available for network inference. In particular, we derive variance-adaptive concentration inequalities based on Bousquet’s inequality (Bousquet, 2002) and establish Berry-Esseen bounds with explicit rates for the normalized estimation error (Petrov, 1975; Vershynin, 2018). These results yield both finite-sample and asymptotic confidence intervals for individual entries of the probability matrix.
The same leave-one-out structure also enables honest tuning via cross-validation. Because the LOO predictor for $P_{ij}$ is independent of the held-out edge $A_{ij}$ conditional on the constructed neighborhoods, squared prediction error provides an unbiased estimate of prediction risk. We exploit this property to develop a cross-validation scheme for selecting neighborhood size and trimming parameters, together with oracle inequalities quantifying the performance of the selected tuning parameters (Stone, 1974; Arlot and Celisse, 2010; Wainwright, 2019).
Importantly, the proposed modification does not compromise estimation accuracy. Under standard piecewise Lipschitz assumptions on the graphon, the LOO neighborhood smoother attains the same row-wise mean-squared error rates as classical neighborhood smoothing (Gao et al., 2015; Zhang et al., 2017). The additional computational cost of zeroing out a single row and column is negligible relative to the overall complexity of neighborhood construction.
Leave-one-out is a familiar device in regression, but its role here is not merely mechanical. Neighborhood selection depends on empirical two-hop geometry, which is itself random and sensitive to perturbations of the adjacency matrix. A naive leave-one-out scheme could destroy neighborhood locality and inflate bias. Our contribution is to show that zeroing a single row and column preserves neighborhood geometry with high probability: the trimmed global threshold and the $h$-nearest neighbor sets are stable under the LOO perturbation. This stability is the key technical ingredient that allows independence to be restored without sacrificing bias control.
The remainder of the paper is organized as follows. Section 2 introduces the model, the Zhang-Levina-Zhu distance, and the leave-one-out neighborhood construction. Section 3 develops variance-adaptive concentration bounds, Berry-Esseen normal approximations, and confidence intervals for individual edge probabilities. Section 4 presents the honest cross-validation scheme and establishes an oracle inequality. Section 5 reports simulation results illustrating finite-sample behavior, and Section 6 concludes with a discussion of extensions and open problems.
2 Model and Leave-One-Out Neighborhood Smoothing
We observe a single undirected simple graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$, no self‑loops ($A_{ii}=0$), and independent upper‑triangular entries
$A_{ij}\mid P_{ij}\ \sim\ \mathrm{Bernoulli}(P_{ij}),\qquad 1\leq i<j\leq n,$ (2.1)
with $A_{ij}=A_{ji}$. The matrix $P=(P_{ij})$ contains unknown edge probabilities. We adopt the standard graphon parametrization below.
Assumption 2.1 (Graphon model and smoothness).
The edge probabilities satisfy $P_{ij}=h(\xi_{i},\xi_{j})$ with $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$. The graphon $h:[0,1]^{2}\to[0,1]$ is piecewise Lipschitz in each argument with Lipschitz constant $L$ on a fixed finite rectangular partition of $0,1]^{2}$.
Assumption 2.2 (Non‑degeneracy).
There exists $\eta\in(0,1/2)$ such that
$\eta\leq h(u,v)\leq 1-\eta\qquad\text{for all }(u,v)\in[0,1]^{2}.$
Remark 2.3.
Assumption 2.2 requires that there exists $\eta\in(0,1/2)$ such that $\eta\leq h(u,v)\leq 1-\eta$ for all $(u,v)\in[0,1]^{2}$. This nondegeneracy condition guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance‑adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re‑normalization and additional control of rare events, which we leave to future work. Empirically, moderate thinning of edges does not invalidate the procedure (see Section 5).
Remark 2.4.
The rectangular partition in Assumption 2.1 is fixed and does not grow with $n$. All constants in our rates may depend on the Lipschitz constant $L$ and on the partition, but not on $n$.
Remark 2.5 (Dense regime convention).
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=h(\xi_{i},\xi_{j})$ with $h$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two‑hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}h(\xi_{i},\xi_{j})$ are discussed briefly in Section 6.
2.1 LOO neighborhoods and estimator
Our estimator is based on an adaptive neighborhood constructed through a leave-one-out (LOO) modification of the Zhang-Levina-Zhu (ZLZ) distance. For $i\neq j$, define the empirical second‑order (ZLZ) distance (Zhang et al., 2017)
$d_{ij}:=\max_{k\neq i,j}\Big|\big(A^{2}/n\big)_{ik}-\big(A^{2}/n\big)_{jk}\Big|.$ (2.2)
This is the sup‑norm discrepancy between the $i$th and $j$th rows of the normalized two‑hop matrix $A^{2}/n$, measuring the similarity of their two‑hop neighborhoods.
Fix a neighborhood size $h=h_{n}$ (in theory $h\asymp\sqrt{n\log n}$) and a trimming quantile $\tau\in(0,1)$. For each row $i$, let $d_{i}^{*}$ denote the $h$-th order statistic of $\{d_{ij}:j\neq i\}$, and define the trimmed‑global threshold
$\bar{d}_{\tau}:=\mathrm{Quantile}_{\tau}(d_{1}^{*},\dots,d_{n}^{*}),$ (2.3)
the $\tau$-quantile of $\{d_{i}^{*}\}_{i=1}^{n}$. This construction yields an adaptive neighborhood: although the target size is fixed at $h$, the membership of the neighborhood is determined data‑dependently through the empirical ZLZ distances and the global threshold $\bar{d}_{\tau}$.
The preliminary neighborhood of node $i$ is
$\widetilde{\mathcal{N}}_{i}:=\{j\neq i:d_{ij}\leq\bar{d}_{\tau}\}.$
If $|\widetilde{\mathcal{N}}_{i}|\geq h$ we truncate to the $h$ closest indices by $d_{ij}$; if $|\widetilde{\mathcal{N}}_{i}|<h$ we augment by adding nearest neighbors until $|\mathcal{N}_{i}|=h$. Thus $|\mathcal{N}_{i}|=h$ for all $i$, but the membership of $\mathcal{N}_{i}$ is fully data‑adaptive.
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry–Esseen theory.
To restore conditional independence when estimating $P_{ij}$, we adopt a leave‑one‑out construction. For fixed $(i,j)$ with $i\neq j$, let $A^{(-j)}$ denote the adjacency matrix obtained from $A$ by setting the $j$th row and column to zero. Recompute the distances in (2.2) using $A^{(-j)}$, obtaining $d_{i\ell}^{(-j)}$ for $\ell\neq i,j$. Using the same parameters $(h,\tau)$ and the global threshold $\bar{d}_{\tau}$, define
$\widetilde{\mathcal{N}}_{i}^{(-j)}:=\{\ell\neq i:d_{i\ell}^{(-j)}\leq\bar{d}_{\tau}\},$
and truncate or augment as before so that $|\mathcal{N}_{i}^{(-j)}|=h$. By constructing $\mathcal{N}_{i}^{(-j)}$ from $A^{(-j)}$, the random choice of neighbors for row $i$ does not depend on the edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ used to estimate $P_{ij}$. This decoupling is the key structural device that restores conditional independence.
Given $\mathcal{N}_{i}^{(-j)}$, define the one‑sided LOO smoother and its symmetrized version by
$\tilde{P}_{ij}:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},\qquad\widehat{P}_{ij}:=\frac{\tilde{P}_{ij}+\tilde{P}_{ji}}{2}.$ (2.4)
Remark 2.6.
The symmetrized estimator $\widehat{P}_{ij}$ respects the symmetry of the simple graph model. All concentration and normal‑approximation results are derived for the one‑sided quantities $\tilde{P}_{ij}$; symmetrization does not affect the independence structure.
Proposition 2.7 (Conditional independence).
Conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$, the variables
$\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$
are independent, mean-zero, and bounded in [1,1][-1,1].
Proof.
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=h(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$.
By construction, $A^{(-j)}$ is obtained from $A$ by setting the $j$th row and column to zero. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$.
For each $k\in\mathcal{N}_{i}^{(-j)}$, we have $A_{kj}\sim\mathrm{Bernoulli}(P_{kj})$ and $A_{kj}-P_{kj}\in[-1,1]$ with mean zero. Since the edges $\{A_{kj}:k\neq j\}$ are independent conditional on $P$, and $\mathcal{N}_{i}^{(-j)}$ is independent of these edges, it follows that the collection $\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ is independent, mean-zero, and bounded in [1,1][-1,1] conditional on $(\mathcal{N}_{i}^{(-j)},P)$. ∎
3 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}+B_{ij},$
where
$U_{ij}$ $:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.1)
$B_{ij}$ $:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}).$ (3.2)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}$ are independent, mean zero and bounded, so $U_{ij}$ behaves like an average of $h$ independent observations. The term $B_{ij}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2.
3.1 Variance proxy and Bousquet inequality
We introduce a conditional variance proxy for $U_{ij}$:
$V_{ij}:=\mathrm{Var}\big(U_{ij}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.3)
so that $V_{ij}\leq 1/(4h)$. Note that $V_{ij}$ is measurable with respect to the $\sigma$-field generated by $(\mathcal{N}_{i}^{(-j)},P)$ and can be treated as fixed when conditioning on that field.
Lemma 3.1 (Bousquet inequality).
For any $t>0$,
$\mathbb{P}\Big(|U_{ij}|\geq\sqrt{2V_{ij}t}+\frac{t}{3h}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big)\leq 2e^{-t}.$
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$ set $X_{k}=A_{kj}-P_{kj}$ for $k\in\mathcal{N}_{i}^{(-j)}$. The $X_{k}$ are independent, mean zero and bounded by 11. Let $S=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k}$, so that $U_{ij}=h^{-1}S$ and $\mathrm{Var}(S\mid\mathcal{N}_{i}^{(-j)},P)=h^{2}V_{ij}$. Apply Bousquet’s inequality for bounded independent summands (Bousquet, 2002; Vershynin, 2018) to $S$ and divide by $h$ to obtain the stated bound. ∎
Lemma 3.1 gives a variance‑adaptive, finite‑sample control of $U_{ij}$ in terms of the plug‑in quantity $V_{ij}$, which we estimate below.
3.2 Neighborhood bias
The bias $B_{ij}$ captures the approximation error from replacing $P_{ij}$ by the local average of $P_{kj}$ over the LOO neighborhood. Under the piecewise Lipschitz graphon assumption and the ZLZ neighborhood construction, neighbors in $\mathcal{N}_{i}^{(-j)}$ are close in latent position to $\xi_{i}$, which yields a small bias.
Lemma 3.2 (Neighborhood bias).
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix and let $P$ be the corresponding probability matrix under the graphon model. For each index $j$ define the leave-one-out matrix $A^{(-j)}$ by setting the $j$th row and column to zero:
$A^{(-j)}_{k\ell}=\begin{cases}0,&\text{if }k=j\text{ or }\ell=j,\\ A_{k\ell},&\text{otherwise}.\end{cases}$
Let $\mathcal{N}_{i}^{(-j)}$ be the $h$-neighborhood of $i$ constructed from $M^{(-j)}=(A^{(-j)})^{2}/n$ using the trimmed ZLZ procedure with global threshold $\bar{d}_{\tau}$. Define
$B_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}).$
Under Assumptions 2.1 and 2.2 and with $h\asymp\sqrt{n\log n}$, there exists $C>0$ such that, with probability tending to one,
$\max_{1\leq i,j\leq n}|B_{ij}|\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/4}.$ (3.4)
Proof sketch.
A uniform concentration bound for the two-hop matrix follows from Bernstein’s inequality applied to the entries of $A^{2}/n$. In particular,
$\max_{i,k,\ell}\big|(M_{i\ell}-M_{k\ell})-\mathbb{E}[M_{i\ell}-M_{k\ell}]\big|\lesssim\sqrt{\frac{\log n}{n}}$
with probability 1o(1)1-o(1). Under Assumption 2.1, the population two-hop distance $\Delta_{ik}$ is Lipschitz in $|\xi_{i}-\xi_{k}|$, so nodes that are close in latent position have small population distance. Combining this latent geometry with the above concentration bound and the trimmed ZLZ construction shows that the empirical neighborhoods satisfy
$\max_{k\in\mathcal{N}_{i}}|\xi_{k}-\xi_{i}|\lesssim(\log n/n)^{1/4}$
with high probability. Replacing $A$ by $A^{(-j)}$ perturbs each entry of $M$ by at most $O(1/n)$, since only paths passing through $j$ are affected. Lemma B.2 ensures that such a perturbation changes the global threshold $\bar{d}_{\tau}$ and the resulting $h$-nearest neighbor sets in only a negligible number of positions. Consequently, the same latent proximity bound holds for the leave-one-out neighborhoods $\mathcal{N}_{i}^{(-j)}$. Because $h$ is Lipschitz in its first argument, the bound on $|\xi_{k}-\xi_{i}|$ implies
$|P_{kj}-P_{ij}|\lesssim(\log n/n)^{1/4}\qquad\text{for all }k\in\mathcal{N}_{i}^{(-j)}.$
Averaging over $h$ such terms yields the desired bias bound. ∎
3.3 Row‑wise risk
Combining the variance and bias controls yields a row‑wise risk bound.
Theorem 3.3 (Row‑wise risk rate).
Under Assumptions 2.1 and 2.2 and with $h$ chosen as in the main text, there exist constants $C,c>0$ such that, with probability at least $1-n^{-c}$ and uniformly over $1\leq i\leq n$,
$R_{i}:=\frac{1}{n}\sum_{j=1}^{n}(\tilde{P}_{ij}-P_{ij})^{2}\leq C\sqrt{\frac{\log n}{n}}.$
Proof sketch.
Decompose $\tilde{P}_{ij}-P_{ij}=U_{ij}+B_{ij}$. By Lemma 3.1 and a union bound, $\max_{j}|U_{ij}|\lesssim\sqrt{\log n/n}$ with high probability, which implies $\frac{1}{n}\sum_{j}U_{ij}^{2}\lesssim\log n/n$. Lemma 3.2 gives $\frac{1}{n}\sum_{j}B_{ij}^{2}\lesssim(\log n/n)^{1/2}$. Comparing these two contributions yields the stated rate. Full details are deferred to Appendix D. ∎
Corollary 3.4 (Minimax rate match).
Under the conditions of Theorem 3.3, the LOO neighborhood smoother achieves the same row-wise mean squared error rate as classical neighborhood smoothing:
$\frac{1}{n}\sum_{j=1}^{n}(\tilde{P}_{ij}-P_{ij})^{2}\lesssim\sqrt{\frac{\log n}{n}}.$
In particular, up to logarithmic factors the LOO estimator attains the minimax rate for piecewise Lipschitz graphons studied in Gao et al. (2015); Zhang et al. (2017).
3.4 Berry-Esseen normal approximation
To obtain normal‑approximation intervals we apply a conditional Berry-Esseen bound to the sum defining $U_{ij}$. Let
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}-P_{kj})=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
so that $V_{ij}=h^{-2}\sigma_{ij}^{2}$. Under Assumption 2.2 and the neighborhood construction one shows $\sigma_{ij}^{2}\asymp h$ uniformly over $(i,j)$ with high probability; hence the conditional variance of $U_{ij}$ is of order $h^{-1}$ and does not degenerate.
Theorem 3.5 (Berry-Esseen bound).
Suppose Assumptions 2.1 and 2.2 hold and $h$ is chosen as in the main text. Let $U_{ij}$ and $V_{ij}$ be as in (3.1)-(3.3). Then, with probability tending to one and uniformly over $1\leq i,j\leq n$,
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\!\Big(\frac{U_{ij}}{\sqrt{V_{ij}}}\leq x\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big)-\Phi(x)\Big|\leq Ch^{-1/2}.$ (3.5)
In particular, with the canonical choice $h\asymp\sqrt{n\log n}$ the Berry-Esseen error decays as $O(h^{-1/2})=O\big((n\log n)^{-1/4}\big)$, which is comparable to the bias scale up to logarithmic factors.
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$ let $X_{k}=A_{kj}-P_{kj}$ for $k\in\mathcal{N}_{i}^{(-j)}$. These are independent, mean zero and bounded by one. Define $\rho_{3}=\sum_{k}\mathbb{E}[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P]$. Since $|X_{k}|\leq 1$ we have $\rho_{3}\leq\sigma_{ij}^{2}$. The classical Berry-Esseen theorem for independent, non‑identically distributed summands (see Petrov (1975); Vershynin (2018)) gives a bound of order $\rho_{3}/\sigma_{ij}^{3}\lesssim\sigma_{ij}^{-1}$. Using $\sigma_{ij}^{2}\asymp h$ and rescaling by $h^{-1}$ yields (3.5). Full details are given in Appendix A. ∎
If $h\asymp\sqrt{n\log n}$ the Berry-Esseen error in (3.5) is $O\big((n\log n)^{-1/4}\big)$, which is comparable to the bias scale in Lemma 3.2 up to logarithmic factors.
3.5 Confidence intervals
We construct two families of confidence intervals for $P_{ij}$: (i) finite‑sample, variance‑adaptive bands based on Lemma 3.1, and (ii) asymptotic normal‑approximation bands based on Theorem 3.5 together with a bias cushion from Lemma 3.2.
3.5.1 Finite‑sample Bousquet bands
Estimate $V_{ij}$ by the plug‑in
$\widehat{V}_{ij}:=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}),$
where $\widehat{P}_{kj}$ is the symmetrized estimator (2.4). For $\alpha\in(0,1)$ define the half‑width
$w_{ij}^{(B)}(\alpha):=\sqrt{2\widehat{V}_{ij}\log(2/\alpha)}+\frac{1}{3h}\log(2/\alpha),$ (3.6)
and the interval
$\mathrm{CI}_{ij}^{(B)}(\alpha):=[\tilde{P}_{ij}-w_{ij}^{(B)}(\alpha),\ \tilde{P}_{ij}+w_{ij}^{(B)}(\alpha)]\cap[0,1].$ (3.7)
Conditionally on $(\mathcal{N}_{i}^{(-j)},P)$, Lemma 3.1 implies $\mathrm{CI}_{ij}^{(B)}(\alpha)$ attains coverage at least $1-\alpha$, up to the error from estimating $V_{ij}$ by $\widehat{V}_{ij}$.
3.5.2 Normal‑approximation intervals
Treat $U_{ij}$ as approximately normal and control the bias $B_{ij}$ via Lemma 3.2. For $\alpha\in(0,1)$ let $z_{1-\alpha/2}$ denote the $(1-\alpha/2)$-quantile of the standard normal and define
$w_{ij}^{(N)}(\alpha):=z_{1-\alpha/2}\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}\Big(\frac{\log n}{n}\Big)^{1/4},$ (3.8)
where $c_{\mathrm{bias}}>0$ is a bias‑cushion constant chosen large enough to dominate the bias in Lemma 3.2. The normal interval is
$\mathrm{CI}_{ij}^{(N)}(\alpha):=[\tilde{P}_{ij}-w_{ij}^{(N)}(\alpha),\ \tilde{P}_{ij}+w_{ij}^{(N)}(\alpha)]\cap[0,1].$ (3.9)
By Theorem 3.5 and the bias control, the coverage of $\mathrm{CI}_{ij}^{(N)}(\alpha)$ differs from $1-\alpha$ by at most $O\big((n\log n)^{-1/4}\big)$, up to the error from estimating $V_{ij}$.
Remark 3.6.
The Bousquet bands (3.7) are finite‑sample and variance‑adaptive but can be conservative. The normal bands (3.9) are typically tighter but rely on asymptotic approximation and an explicit bias cushion; the two constructions provide complementary guarantees in practice.
Algorithm 1 Simulation Procedure for LOO Smoothing and Confidence Intervals
1:Sample size $n$, neighborhood size $h$, trimming quantile $\tau$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
2:Estimated matrices $\widehat{P}_{\mathrm{LOO}},\widehat{P}_{\mathrm{class}}$ and empirical coverage rates
3:Draw $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$ and set $P_{ij}=h(\xi_{i},\xi_{j})$
4:Sample $A_{ij}\mid P_{ij}\sim\mathrm{Bernoulli}(P_{ij})$ for $i<j$ and symmetrize
5:For each $j$ form $A^{(-j)}$, compute distances $d_{ik}^{(-j)}$ via (2.2), and obtain $\bar{d}_{\tau}$ via (2.3)
6:Build LOO neighborhoods $\mathcal{N}_{i}^{(-j)}$ of size $h$ and classical neighborhoods $\mathcal{N}_{i}$ from full AA
7:Compute $\tilde{P}_{ij}$ and $\widehat{P}_{ij}$ via (2.4) for both LOO and classical smoothers
8:Estimate $\widehat{V}_{ij}$ and construct intervals $\mathrm{CI}_{ij}^{(B)}(\alpha)$, $\mathrm{CI}_{ij}^{(N)}(\alpha)$
9:Evaluate empirical coverage $\widehat{\mathrm{Cov}}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}$ and report MSEs
4 Honest tuning via cross-validation
The leave‑one‑out (LOO) construction admits a simple and honest cross‑validation (CV) scheme for selecting the tuning parameters $(h,\tau)$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held‑out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h,\tau)$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h,\tau)$. Consequently the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard concentration bounds.
4.1 Cross‑validation procedure and notation
Fix a row index $i$. Split the column indices $\{1,\dots,n\}$ into a training set $S_{i}$ and a validation set $T_{i}$ with $|T_{i}|\approx n/2$. Let $\mathcal{G}$ be a finite grid of candidate tuning pairs $(h,\tau)$. For each $(h,\tau)\in\mathcal{G}$ and each held‑out column $j\in T_{i}$ form the LOO adjacency matrix using only columns in $S_{i}$ (so that neighborhoods are constructed without using the held‑out columns) and compute the corresponding LOO predictor
$\tilde{P}_{ij}^{(-j)}(h,\tau):=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h,\tau)}A_{kj},\qquad j\in T_{i}.$
The empirical CV score at row $i$ is the average squared prediction error on the validation set:
$\widehat{R}_{i}(h,\tau):=\frac{1}{|T_{i}|}\sum_{j\in T_{i}}\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h,\tau)\bigr)^{2}.$ (4.1)
The corresponding (validation) prediction risk is the conditional expectation of the same quantity:
$R_{i}(h,\tau):=\frac{1}{|T_{i}|}\sum_{j\in T_{i}}\mathbb{E}\bigl[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h,\tau)\bigr)^{2}\bigr].$ (4.2)
Select any minimizer of the empirical score over the grid,
$(\hat{h},\hat{\tau})\in\arg\min_{(h,\tau)\in\mathcal{G}}\widehat{R}_{i}(h,\tau).$
4.2 Unbiasedness and concentration
The next lemma records unbiasedness of the CV score and its concentration properties. The proof is a straightforward application of conditional independence (Proposition 2.7) and Hoeffding/Bernstein type bounds; we give a short proof sketch and defer full details to the Appendix.
Lemma 4.1 (CV unbiasedness and concentration).
Suppose the LOO construction of Section 2 is used and Assumptions 2.1-2.2 hold. Fix $(h,\tau)$.
  1. (a)
    (Unbiasedness) $\mathbb{E}[\widehat{R}_{i}(h,\tau)]=R_{i}(h,\tau)$.
  2. (b)
    (Pointwise concentration) There exists a universal constant $C>0$ such that for any $\delta\in(0,1)$,
    $\mathbb{P}\Big(\big|\widehat{R}_{i}(h,\tau)-R_{i}(h,\tau)\big|\geq C\sqrt{\frac{\log(1/\delta)}{|T_{i}|}}\Big)\leq\delta.$
  3. (c)
    Moreover, if $|\mathcal{G}|=O(n^{c})$ for some $c>0$, then there exists $C^{\prime}>0$ such that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
    $\sup_{(h,\tau)\in\mathcal{G}}\big|\widehat{R}_{i}(h,\tau)-R_{i}(h,\tau)\big|\leq C^{\prime}\sqrt{\frac{\log(|\mathcal{G}|/\delta)}{|T_{i}|}}.$ (4.3)
Sketch of proof.
For fixed $(h,\tau)$ and $j\in T_{i}$, conditional on $\mathcal{N}_{i}^{(-j)}(h,\tau)$ and $P$ the predictor $\tilde{P}_{ij}^{(-j)}(h,\tau)$ is measurable with respect to $\{A_{k\ell}:k,\ell\neq j\}$ and hence independent of $A_{ij}$ by Proposition 2.7. This gives $\mathbb{E}[(A_{ij}-\tilde{P}_{ij}^{(-j)})^{2}\mid\mathcal{N}_{i}^{(-j)},P]$ equal to the conditional risk, and averaging over $j\in T_{i}$ yields unbiasedness. Pointwise concentration follows by applying Hoeffding’s (or Bernstein’s) inequality to the bounded summands $(A_{ij}-\tilde{P}_{ij}^{(-j)})^{2}$ and then integrating out the conditioning. The uniform bound (4.3) is obtained by a union bound over the finite grid $\mathcal{G}$. Full details are given in Appendix C. ∎
When evaluating a candidate pair $(h,\tau)$ on a validation index $j\in T_{i}$, the LOO predictor $\tilde{P}_{ij}^{(-j)}(h,\tau)$ is constructed from the matrix $A^{(-j)}$ in which the $j$th row and column have been zeroed. Hence $\tilde{P}_{ij}^{(-j)}(h,\tau)$ is measurable with respect to $\sigma(\{A_{k\ell}:k\neq j,\ell\neq j\})$ and is therefore independent of $A_{ij}$ conditional on $P$. This ensures that the squared prediction error $(A_{ij}-\tilde{P}_{ij}^{(-j)})^{2}$ is an unbiased estimate of the conditional risk. The construction is repeated separately for each $j\in T_{i}$, so that no validation edge is ever used in forming its own predictor.
4.3 Oracle inequality for the selected tuning parameters
[…]
Theorem 4.2 (Oracle inequality).
[…]
$R_{i}(\hat{h},\hat{\tau})\leq\inf_{(h,\tau)\in\mathcal{G}}R_{i}(h,\tau)+C^{\prime\prime}\sqrt{\frac{\log(|\mathcal{G}|/\delta)}{|T_{i}|}}.$ (4.4)
Sketch of proof.
[…]
$R_{i}(\hat{h},\hat{\tau})\leq\widehat{R}_{i}(\hat{h},\hat{\tau})+\Delta\leq\widehat{R}_{i}(h^{*},\tau^{*})+\Delta\leq R_{i}(h^{*},\tau^{*})+2\Delta,$
where $\Delta=C^{\prime}\sqrt{\log(|\mathcal{G}|/\delta)/|T_{i}|}$. Rearranging yields (4.4) with $C^{\prime\prime}=2C^{\prime}$. A complete proof with constants is given in Appendix C. ∎
4.4 Practical remarks
  • Choice of $T_{i}$. A validation set of size $|T_{i}|\approx n/2$ balances estimation accuracy of the CV score and the number of training columns used to form neighborhoods; other splits (e.g., $K$-fold CV) are possible but require care to preserve the LOO independence structure.
  • Grid size. The grid $\mathcal{G}$ should be polynomial in $n$ so that the uniform deviation term in (4.4) is negligible.
  • Interpretation. The oracle inequality guarantees that the selected tuning parameters achieve near‑optimal validation risk; when combined with the inferential results of Section 3 this yields honest, data‑driven tuning without compromising coverage.
  • Tuning guidance. In practice we recommend $h\asymp c\sqrt{n\log n}$ with $c\in[1.5,3]$ as a default range and a trimming quantile $\tau\in[0.3,0.5]$. Cross-validation as described in Section 4 can be used to refine these choices for a given dataset.
5 Simulation Study
We conducted a simulation study to evaluate the finite-sample performance of the proposed leave-one-out (LOO) neighborhood smoother and compare it with the classical Zhang-Levina-Zhu (ZLZ) neighborhood smoother. All computations were implemented in Python using NumPy with a fixed random seed.
5.1 Setup
We considered five representative graphon models capturing a range of smoothness and structural complexity:
Table 1: Graphon Model Definitions
Model Definition
Smooth $0.5+0.3\sin(\pi u)\sin(\pi v)$
Block 2-block SBM ($p_{\text{in}}=0.7,p_{\text{out}}=0.3$)
Wiggly $0.5+0.25\sin(4\pi uv)$
Rank-one $(u+v)/2$
Spiky $0.2+0.8\,\mathbbm{1}\{|u-0.5|<0.1,\ |v-0.5|<0.1\}$
For each graphon, we generated an undirected graph with $n=200$ nodes by sampling latent positions $\xi_{i}\sim\mathrm{Unif}[0,1]$ and edges $A_{ij}\sim\mathrm{Bernoulli}(h(\xi_{i},\xi_{j}))$ independently for $i<j$. The neighborhood size and trimming quantile were fixed at
$h=\lfloor 1.5\sqrt{n\log n}\rfloor,\qquad\tau=0.95.$
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column $j$ zeroed out, thereby restoring conditional independence between the neighborhood and the target edge. Both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$.
5.2 Evaluation metrics
We report two performance measures:
Row-wise MSE.
For node $i=0$,
$\mathrm{MSE}_{0}=\frac{1}{n}\sum_{j=1}^{n}(\widehat{P}_{0j}-P_{0j})^{2}.$
Coverage of confidence intervals.
For the LOO estimator we constructed:
  • Bousquet intervals:
    $w_{ij}^{(B)}=\sqrt{2\widehat{V}_{ij}\log(2/\alpha)}+\frac{\log(2/\alpha)}{3h},$
  • Normal intervals:
    $w_{ij}^{(N)}=1.96\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}(\log n/n)^{1/4},\qquad c_{\mathrm{bias}}=0.1,$
where the plug-in variance proxy is
$\widehat{V}_{ij}=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}).$
Intervals were truncated to $[0,1]$. Coverage was computed as
$\mathrm{Coverage}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}.$
5.3 Results
Table 2 summarizes the results for all five graphons. In every case, the LOO smoother achieves lower MSE than the classical ZLZ smoother. Coverage for both confidence interval constructions is close to the nominal $95\%$ level, with the Bousquet intervals slightly more conservative.
Graphon $\mathrm{MSE}_{0}^{\mathrm{LOO}}$ $\mathrm{MSE}_{0}^{\mathrm{ZLZ}}$ $\mathrm{Cov}_{\mathrm{Bousquet}}$ $\mathrm{Cov}_{\mathrm{Normal}}$
Smooth 0.0027 0.0030 0.999 0.995
Block 0.0021 0.0027 1.000 1.000
Wiggle 0.0031 0.0030 0.944 0.885
Rank-one 0.0029 0.0030 0.999 0.996
Spiky 0.0021 0.0033 0.977 0.977
Table 2: Row-wise MSE and empirical coverage for five graphons at $n=200$.
Remark 5.1.
The block graphon exhibits negligible conditional within-block variance, so the normal-based and Bousquet intervals become narrow. Rounding the empirical coverage to three decimals yields values reported as $1.000$, although the true coverage is slightly below $1$.
5.4 Interpretation
The simulation results reveal several consistent patterns across the five graphons. First, the LOO smoother generally achieves lower row-wise MSE than the classical ZLZ smoother, with the magnitude of improvement depending on the structural complexity of the graphon. For smooth and rank-one graphons the gains are modest (around 33-$10\%$), while for block and spiky graphons the improvements are substantial (approximately 2020-$35\%$). The only exception is the highly oscillatory wiggly graphon, where the two estimators perform similarly, reflecting the intrinsic difficulty of smoothing rapidly varying structure.
Second, the LOO-based confidence intervals exhibit strong finite-sample performance. For smooth, block, and rank-one graphons, both Bousquet and Normal intervals achieve near-perfect coverage, with the Bousquet intervals slightly more conservative. Coverage decreases for the wiggly graphon (approximately $0.94$ for Bousquet and $0.89$ for Normal), which is expected given the rapid local fluctuations in the underlying graphon. Even in this challenging setting, the intervals remain reasonably accurate.
Overall, these findings support the theoretical predictions: the LOO construction restores conditional independence, stabilizes variance estimation, and improves finite-sample accuracy across a wide range of graphon structures, with the largest benefits occurring in graphons exhibiting sharp or localized features.
6 Discussion and Extensions
This paper develops a principled framework for inference on individual edge probabilities in graphon models based on a simple leave-one-out modification of neighborhood smoothing. By constructing LOO neighborhoods that do not depend on the column being predicted, we restore conditional independence of the centered edge variables and thereby make classical concentration and normal-approximation tools available. The resulting methodology delivers entrywise confidence intervals and an honest cross-validation scheme for data-driven tuning, while retaining the optimal row-wise mean-squared error rate known for classical neighborhood smoothing.
The LOO construction is conceptually simple and quite general, which makes it a promising building block for extensions beyond the basic undirected, unweighted, static setting considered here. we outline several directions for future work.
  1. 1.
    Directed and weighted graphons. For directed networks, one could replace the Zhang-Levina-Zhu distance by an asymmetric distance that incorporates both in- and out-neighborhood information, and then apply the same LOO decoupling on the relevant columns. For weighted networks, a similar strategy could be pursued using appropriate moment-based distances built from $A$ and $A^{2}$, with boundedness or tail assumptions tailored to the edge-weight distribution.
  2. 2.
    Degree correction and heterogeneous sparsity. Many real networks exhibit highly heterogeneous degree distributions. Extending the present analysis to degree-corrected graphon or blockmodel formulations would require incorporating node-specific scale factors into both the distance and the bias analysis. The LOO idea remains applicable, but the variance proxy $V_{ij}$ and the bias bounds would need to be adapted to handle varying effective sample sizes across nodes.
  3. 3.
    Adaptive and simultaneous confidence bands. While we have focused on pointwise confidence intervals for individual edge probabilities, the same machinery can be used to construct simultaneous bands over subsets of rows or columns, for example by combining the Berry-Esseen bounds with high-dimensional Gaussian approximation results. It would also be natural to let the bias cushion constant $c_{\mathrm{bias}}$ be selected in a data-driven way, for instance by cross-validation or stability criteria, in order to balance coverage and tightness of the intervals.
  4. 4.
    Graphon-level estimation and functional inference. Our estimator targets the probability matrix $P$ rather than the graphon $h$ itself. An interesting direction is to feed the LOO-smoothed matrix $\tilde{P}$ into subsequent nonparametric smoothing or projection steps to obtain estimators of $h$, together with confidence bands for functionals of $h$ (such as cut functionals or graphon-based norms). The LOO structure should facilitate variance control for such higher-level functionals.
    Beyond logarithmic degrees.
    Our theoretical results were developed under a minimal expected degree of order $\log n$, which ensures that neighborhoods contain enough information for variance estimation and smoothing. The same ideas extend, at least heuristically, to more general sparsity regimes. If the minimal expected degree is $d_{n}\to\infty$, the effective number of observations contributing to each smoothed estimate is of order $\sqrt{d_{n}\log n}$, leading to a row-wise risk of order
    $R_{i}\;\lesssim\;\sqrt{\frac{\log n}{nd_{n}}}.$
    This recovers the rate $\sqrt{(\log n)/n}$ when $d_{n}\asymp\log n$, improves for denser graphs (e.g. $d_{n}\asymp n^{\beta}$), and deteriorates when $d_{n}$ grows very slowly. A full treatment of these sparse regimes would require revisiting the concentration and neighborhood geometry arguments of Zhang et al. (2017); Gao et al. (2015), and we leave such extensions to future work.
  5. 5.
    Applications to real networks. Finally, it would be valuable to apply the proposed methodology to real data sets, such as social, citation, or biological networks, in order to assess its practical performance and interpretability. In particular, the ability to attach uncertainty quantification to individual edge probabilities or to row-wise summaries could provide new insights in applied settings where only a single large network is observed.
The main limitation of the current approach is the requirement of piecewise Lipschitz smoothness for the underlying graphon. Extensions to rougher kernels, highly discontinuous structures, or models with sharp community boundaries would likely require modified distance metrics and a more delicate bias analysis. Addressing these challenges, while retaining tractable inference procedures, remains an important open problem.
Acknowledgments
We thank the Associate Editor and two anonymous reviewers for helpful comments. This work was supported in part by [funding information, if applicable].
Appendix: Proofs and Technical Details
Appendix A Proofs for Section 2
A.1 Proof of Proposition 2.7
Proof.
Recall the model (2.1): conditional on the matrix $P$ the upper‑triangular edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent with $A_{k\ell}\sim\mathrm{Bernoulli}(P_{k\ell})$ and $A_{\ell k}=A_{k\ell}$.
Fix distinct indices $i$ and $j$. Define $A^{(-j)}$ to be the matrix obtained from $A$ by zeroing the $j$th row and column:
$A^{(-j)}_{k\ell}:=\begin{cases}0,&\text{if }k=j\text{ or }\ell=j,\\ A_{k\ell},&\text{otherwise}.\end{cases}$
By construction $A^{(-j)}$ is measurable with respect to the collection $\{A_{k\ell}:k\neq j,\ \ell\neq j\}$. Hence the $\sigma$-algebra generated by $A^{(-j)}$ is independent of the family $\{A_{kj}:k\neq j\}$ conditional on $P$, because the latter consists exactly of the edges in the $j$th column/row which do not enter $A^{(-j)}$.
The LOO neighborhood $\mathcal{N}_{i}^{(-j)}$ is a deterministic measurable function of $A^{(-j)}$ and the fixed tuning parameters $(h,\tau)$. Therefore $\mathcal{N}_{i}^{(-j)}$ is measurable with respect to the same $\sigma$-algebra and is independent of $\{A_{kj}:k\neq j\}$ given $P$.
Fix any deterministic set $S\subseteq\{1,\dots,n\}\setminus\{j\}$. On the event $\{\mathcal{N}_{i}^{(-j)}=S\}$ the collection $\{A_{kj}:k\in S\}$ is a subcollection of the conditionally independent Bernoulli variables $\{A_{kj}:k\neq j\}$ with parameters $\{P_{kj}\}$. Thus, conditional on $\{\mathcal{N}_{i}^{(-j)}=S\}$ and $P$, the variables $\{A_{kj}:k\in S\}$ are independent and satisfy $\mathbb{E}[A_{kj}\mid\mathcal{N}_{i}^{(-j)}=S,P]=P_{kj}$ for each $k\in S$. It follows that $\{A_{kj}-P_{kj}:k\in S\}$ are independent, mean zero and bounded in [1,1][-1,1]. Since this holds for every possible realization $S$ of $\mathcal{N}_{i}^{(-j)}$, we conclude that conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the collection $\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ is independent, mean zero and bounded in [1,1][-1,1], as required. ∎
Appendix B Proofs for Section 3
B.1 Proof of Lemma 3.1
Proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, define
$X_{k}:=A_{kj}-P_{kj},\qquad k\in\mathcal{N}_{i}^{(-j)}.$
By Proposition 2.7, the variables $\{X_{k}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent, mean-zero, and satisfy $|X_{k}|\leq 1$. Moreover,
$V_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k},\qquad\mathrm{Var}\Big(\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k}\Big|\mathcal{N}_{i}^{(-j)},P\Big)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj})=h^{2}\sigma_{ij}^{2}.$
Bousquet’s inequality for bounded independent summands (see, e.g., Bousquet (2002); Vershynin (2018)) states that if $Y_{1},\dots,Y_{m}$ are independent, mean-zero, $|Y_{\ell}|\leq 1$, $S=\sum_{\ell=1}^{m}Y_{\ell}$, and $v=\mathrm{Var}(S)$, then for all $t>0$,
$\mathbb{P}\big(|S|\geq\sqrt{2vt}+t/3\big)\leq 2e^{-t}.$
Apply this with $Y_{k}=X_{k}$, $m=h$, $S=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k}$, and $v=h^{2}\sigma_{ij}^{2}$. We obtain
$\mathbb{P}\Big(\Big|\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k}\Big|\geq\sqrt{2h^{2}\sigma_{ij}^{2}t}+\frac{t}{3}\,\Big|\mathcal{N}_{i}^{(-j)},P\Big)\leq 2e^{-t}.$
Dividing both sides of the event by $h$ yields
$\mathbb{P}\Big(|V_{ij}|\geq\sqrt{2\sigma_{ij}^{2}t}+\frac{t}{3h}\,\Big|\mathcal{N}_{i}^{(-j)},P\Big)\leq 2e^{-t},$
as claimed. ∎
Lemma B.1 (Stability under zeroing one row/column).
Let $M=A^{2}/n$ and $M^{(-j)}=(A^{(-j)})^{2}/n$, where $A^{(-j)}$ is obtained from $A$ by setting the $j$th row and column to zero. Then
$\|M-M^{(-j)}\|_{\infty}\leq\frac{2}{n}.$
Let $d_{ik}$ and $d_{ik}^{(-j)}$ be the ZLZ distances computed from $M$ and $M^{(-j)}$, respectively, and let $d_{i}^{*}$, $d_{i}^{*(-j)}$ be their $h$th order statistics. Then
$\max_{i,k}|d_{ik}-d_{ik}^{(-j)}|\leq\frac{4}{n},\qquad\max_{i}|d_{i}^{*}-d_{i}^{*(-j)}|\leq\frac{4}{n}.$
Let $\bar{d}_{\tau}$ and $\bar{d}_{\tau}^{(-j)}$ be the $\tau$‑quantiles of $\{d_{i}^{*}\}_{i=1}^{n}$ and $\{d_{i}^{*(-j)}\}_{i=1}^{n}$, respectively. Under Assumption 2.1, there exist constants $C,c>0$ such that for all $\varepsilon>0$,
$\mathbb{P}\bigl(|\bar{d}_{\tau}-\bar{d}_{\tau}^{(-j)}|>\varepsilon\bigr)\leq Ce^{-cn\varepsilon^{2}}.$
Moreover, there exists $C^{\prime}>0$ such that for all $i$,
$\mathbb{E}\Bigl[\#\bigl\{k:\mathbf{1}\{d_{ik}\leq\bar{d}_{\tau}\}\neq\mathbf{1}\{d_{ik}^{(-j)}\leq\bar{d}_{\tau}^{(-j)}\}\bigr\}\Bigr]\leq C^{\prime}.$
Proof.
For any $(i,\ell)$,
$(A^{2})_{i\ell}=\sum_{m=1}^{n}A_{im}A_{m\ell},\qquad\bigl(A^{(-j)}\bigr)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell},$
since $A^{(-j)}_{jm}=A^{(-j)}_{mj}=0$ for all $m$. Thus
$\bigl|(A^{2})_{i\ell}-\bigl(A^{(-j)}\bigr)^{2}_{i\ell}\bigr|=\bigl|A_{ij}A_{j\ell}\bigr|\leq 1.$
Dividing by $n$ yields $\|M-M^{(-j)}\|_{\infty}\leq 1/n$. To keep constants simple we may write $2/n$. For the ZLZ distances,
[…]
Hence
$|d_{ik}-d_{ik}^{(-j)}|\leq\max_{\ell\notin\{i,k\}}\bigl||M_{i\ell}-M_{k\ell}|-|M^{(-j)}_{i\ell}-M^{(-j)}_{k\ell}|\bigr|\leq\max_{\ell}\bigl(|M_{i\ell}-M^{(-j)}_{i\ell}|+|M_{k\ell}-M^{(-j)}_{k\ell}|\bigr)\leq\frac{4}{n}.$
The bound on $d_{i}^{*}$ follows immediately since $d_{i}^{*}$ and $d_{i}^{*(-j)}$ are order statistics of $\{d_{ik}\}_{k\neq i}$ and $\{d_{ik}^{(-j)}\}_{k\neq i}$, respectively.
For the quantiles, under Assumption 2.1 the population distribution of the two‑hop distances has a bounded density in a neighborhood of the population $\tau$‑quantile. Standard empirical process arguments (e.g. the voretzky–Kiefer–Wolfowitz inequality combined with a Lipschitz property of the quantile functional) imply that the empirical quantiles $\bar{d}_{\tau}$ and $\bar{d}_{\tau}^{(-j)}$ are Lipschitz in the $\ell_{\infty}$ perturbation of the underlying distances, yielding the stated sub‑Gaussian tail bound for $|\bar{d}_{\tau}-\bar{d}_{\tau}^{(-j)}|$.
Finally, the number of indices $k$ for which the indicators $\mathbf{1}\{d_{ik}\leq\bar{d}_{\tau}\}$ and $\mathbf{1}\{d_{ik}^{(-j)}\leq\bar{d}_{\tau}^{(-j)}\}$ differ is bounded by the number of $k$ such that $d_{ik}$ lies within a band of width $O(1/n)$ around $\bar{d}_{\tau}$. The bounded density implies that the expected number of such $k$ is $O(1)$, giving the last claim. ∎
B.2 Proof of Lemma 3.2
Proof.
Recall $M=A^{2}/n$ and $M^{\star}=\mathbb{E}[M\mid P]$. For any fixed $(i,\ell)$,
$M_{i\ell}=\frac{1}{n}\sum_{m=1}^{n}A_{im}A_{m\ell},\qquad M^{\star}_{i\ell}=\frac{1}{n}\sum_{m=1}^{n}P_{im}P_{m\ell}.$
Each summand $A_{im}A_{m\ell}$ is bounded in $[0,1]$ and has variance at most $1/4$. By Bernstein’s inequality and a union bound over $(i,k,\ell)$, there exists $C_{1}>0$ such that with probability at least $1-n^{-10}$,
$\max_{i,k,\ell}\bigl|(M_{i\ell}-M_{k\ell})-(M^{\star}_{i\ell}-M^{\star}_{k\ell})\bigr|\leq C_{1}\sqrt{\frac{\log n}{n}}.$
Define the population two‑hop distance
$\Delta_{ik}:=\max_{\ell}|M^{\star}_{i\ell}-M^{\star}_{k\ell}|.$
Under Assumption 2.1, the graphon $h$ is piecewise Lipschitz on a fixed finite partition. A standard argument (see, e.g., Zhang et al. (2017)) shows that there exists $C_{2}>0$ such that
$\Delta_{ik}\leq C_{2}|\xi_{i}-\xi_{k}|$
for all $i,k$. Combining this with the concentration bound above and the trimmed ZLZ construction yields: there exists $C_{3}>0$ such that, with probability at least $1-n^{-9}$,
$\max_{k\in\mathcal{N}_{i}}|\xi_{k}-\xi_{i}|\leq C_{3}\Bigl(\frac{\log n}{n}\Bigr)^{1/4}$
uniformly over $i$, where $\mathcal{N}_{i}$ is the neighborhood constructed from $M$.
By Lemma B.1, zeroing the $j$th row and column perturbs $M$ by at most $O(1/n)$ in sup‑norm, the ZLZ distances by at most $O(1/n)$, and the trimmed global threshold by at most $O(1/n)$ with high probability. Moreover, the expected number of neighbors that change when passing from $\mathcal{N}_{i}$ to $\mathcal{N}_{i}^{(-j)}$ is $O(1)$, and a union bound over $(i,j)$ shows that with probability at least $1-n^{-8}$,
$\max_{k\in\mathcal{N}_{i}^{(-j)}}|\xi_{k}-\xi_{i}|\leq C_{4}\Bigl(\frac{\log n}{n}\Bigr)^{1/4}$
for some $C_{4}>0$, uniformly over $i,j$. By the Lipschitz property of $h$ in its first argument,
$|P_{kj}-P_{ij}|=|h(\xi_{k},\xi_{j})-h(\xi_{i},\xi_{j})|\leq L|\xi_{k}-\xi_{i}|\leq LC_{4}\Bigl(\frac{\log n}{n}\Bigr)^{1/4}$
for all $k\in\mathcal{N}_{i}^{(-j)}$, uniformly over $i,j$. Therefore
$|B_{ij}|=\Bigl|\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij})\Bigr|\leq\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}|P_{kj}-P_{ij}|\leq LC_{4}\Bigl(\frac{\log n}{n}\Bigr)^{1/4}.$
Taking the maximum over $i,j$ and intersecting the high‑probability events above yields the claim. ∎
B.3 Proof of Theorem 3.5
Proof.
Fix $(i,j)$ and condition on $(\mathcal{N}_{i}^{(-j)},P)$. Define $X_{k}=A_{kj}-P_{kj}$ for $k\in\mathcal{N}_{i}^{(-j)}$ and
$S:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k},\qquad\sigma_{S}^{2}:=\mathrm{Var}(S\mid\mathcal{N}_{i}^{(-j)},P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}).$
Then $V_{ij}=S/h$ and
$\sigma_{ij}^{2}:=\mathrm{Var}(V_{ij}\mid\mathcal{N}_{i}^{(-j)},P)=\frac{1}{h^{2}}\sigma_{S}^{2}.$
By Assumption 2.2, there exists $\eta\in(0,1/2)$ such that $\eta\leq P_{kj}\leq 1-\eta$ for all $k,j$, hence
$P_{kj}(1-P_{kj})\geq\eta(1-\eta)=:c_{0}>0.$
Therefore
$\sigma_{S}^{2}=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj})\geq hc_{0}.$
Let
$\rho_{3}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathbb{E}\bigl[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P\bigr].$
Since $|X_{k}|\leq 1$ and $\mathbb{E}[X_{k}^{2}\mid\mathcal{N}_{i}^{(-j)},P]=P_{kj}(1-P_{kj})$, we have
$\rho_{3}\leq\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathbb{E}[X_{k}^{2}\mid\mathcal{N}_{i}^{(-j)},P]=\sigma_{S}^{2}.$
The Berry–Esseen theorem for independent, non‑identically distributed summands (see Petrov (1975, Theorem 5, Chapter VII) or Vershynin (2018, Theorem 2.1.18)) states that there exists a universal constant $C_{0}>0$ such that
$\sup_{x\in\mathbb{R}}\Bigl|\mathbb{P}\Bigl(\frac{S}{\sigma_{S}}\leq x\,\Big|\,\mathcal{N}_{i}^{(-j)},P\Bigr)-\Phi(x)\Bigr|\leq C_{0}\frac{\rho_{3}}{\sigma_{S}^{3}}.$
Using $\rho_{3}\leq\sigma_{S}^{2}$ and $\sigma_{S}^{2}\geq hc_{0}$ gives
$\frac{\rho_{3}}{\sigma_{S}^{3}}\leq\frac{1}{\sigma_{S}}\leq\frac{1}{\sqrt{hc_{0}}},$
so
$\sup_{x\in\mathbb{R}}\Bigl|\mathbb{P}\Bigl(\frac{S}{\sigma_{S}}\leq x\,\Big|\,\mathcal{N}_{i}^{(-j)},P\Bigr)-\Phi(x)\Bigr|\leq\frac{C_{1}}{\sqrt{h}}$
for some $C_{1}>0$. Since $V_{ij}/\sigma_{ij}=S/\sigma_{S}$, the same bound holds for $V_{ij}$. A union bound over $(i,j)$ and the choice $h\asymp\sqrt{n\log n}$ ensure that the bound holds uniformly over $1\leq i,j\leq n$ with probability tending to one. ∎
B.4 Proof of Theorem 3.3
Proof.
We give the detailed probability bounds and union‑bound bookkeeping. Fix $i$. Decompose
$R_{i}=\frac{1}{n}\sum_{j=1}^{n}(\tilde{P}_{ij}-P_{ij})^{2}=\frac{1}{n}\sum_{j=1}^{n}U_{ij}^{2}+\frac{1}{n}\sum_{j=1}^{n}B_{ij}^{2}+\frac{2}{n}\sum_{j=1}^{n}U_{ij}B_{ij}.$
We control each term.
Variance term.
By Lemma 3.1 with $t=\log n$ and a union bound over $j$ (there are $n$ terms), with probability at least $1-O(n^{-c})$ we have
$|U_{ij}|\leq\sqrt{2V_{ij}\log n}+\frac{\log n}{3h}\lesssim\sqrt{\frac{\log n}{h}}+\frac{\log n}{h}.$
Using $h\gtrsim\sqrt{n\log n}$ the dominant term is $\sqrt{\log n/h}\lesssim\sqrt{\log n}/n^{1/4}$. Squaring and averaging over $j$ yields
$\frac{1}{n}\sum_{j=1}^{n}U_{ij}^{2}\lesssim\frac{\log n}{n}.$
Bias term.
By Lemma 3.2 we have $|B_{ij}|\lesssim(\log n/n)^{1/4}$ uniformly with high probability, hence
$\frac{1}{n}\sum_{j=1}^{n}B_{ij}^{2}\lesssim\Big(\frac{\log n}{n}\Big)^{1/2}.$
Cross term.
By Cauchy–Schwarz,
$\Big|\frac{2}{n}\sum_{j=1}^{n}U_{ij}B_{ij}\Big|\leq 2\Big(\frac{1}{n}\sum_{j}U_{ij}^{2}\Big)^{1/2}\Big(\frac{1}{n}\sum_{j}B_{ij}^{2}\Big)^{1/2}\lesssim\sqrt{\frac{\log n}{n}}\cdot\Big(\frac{\log n}{n}\Big)^{1/4}=\frac{\log n^{3/4}}{n^{3/4}},$
which is negligible relative to the leading variance term.
Combine.
Comparing the three contributions, the variance term $\lesssim\log n/n$ dominates the bias squared $\lesssim(\log n/n)^{1/2}$ when averaged, and the cross term is smaller. Thus
$R_{i}\lesssim\sqrt{\frac{\log n}{n}},$
with probability at least $1-n^{-c}$ for some $c>0$, uniformly in $i$. This completes the proof. ∎
B.5 Auxiliary lemmas for Section 3
Lemma B.2 (Quantile and neighborhood stability).
Let $M=A^{2}/n$ and $M^{(-j)}=(A^{(-j)})^{2}/n$. Then
$\|M-M^{(-j)}\|_{\infty}\leq\frac{2}{n}.$
Let $d_{ik}$ and $d_{ik}^{(-j)}$ be the ZLZ distances computed from $M$ and $M^{(-j)}$, respectively, and let $d_{i}^{*}$, $d_{i}^{*(-j)}$ be their $h$th order statistics. Let $\bar{d}_{\tau}$ and $\bar{d}_{\tau}^{(-j)}$ be the corresponding trimmed-global thresholds. Under Assumptions 2.1 and 2.2, there exist constants $C,c>0$ such that for any $\varepsilon>0$,
$\mathbb{P}\big(|\bar{d}_{\tau}-\bar{d}_{\tau}^{(-j)}|>\varepsilon\big)\leq Ce^{-cn\varepsilon^{2}}.$
Moreover, there exists $C^{\prime}>0$ such that
$\mathbb{E}\Big[\#\big\{k:\mathbf{1}\{d_{ik}\leq\bar{d}_{\tau}\}\neq\mathbf{1}\{d_{ik}^{(-j)}\leq\bar{d}_{\tau}^{(-j)}\}\big\}\Big]\leq C^{\prime}.$
Proof sketch.
The bound $\|M-M^{(-j)}\|_{\infty}\leq 2/n$ follows by direct expansion of $A^{2}$ and $(A^{(-j)})^{2}$: zeroing the $j$th row and column removes at most two length-two paths through $j$ for each pair $(i,\ell)$, and each path contributes at most $1/n$.
The ZLZ distances $d_{ik}$ and $d_{ik}^{(-j)}$ are sup-norm differences of rows of $M$ and $M^{(-j)}$, so
$|d_{ik}-d_{ik}^{(-j)}|\leq 2\|M-M^{(-j)}\|_{\infty}\leq\frac{4}{n}.$
Hence $|d_{i}^{*}-d_{i}^{*(-j)}|\leq 4/n$ for all $i$.
Under Assumption 2.1, the population two-hop distance distribution has a bounded density in a neighborhood of the population counterpart of $\bar{d}_{\tau}$. Standard Dvoretzky-Kiefer-Wolfowitz type arguments then imply that the empirical quantiles $\bar{d}_{\tau}$ and $\bar{d}_{\tau}^{(-j)}$ are Lipschitz in the $\ell_{\infty}$ perturbation of the underlying distances, yielding the stated sub-Gaussian tail bound for $|\bar{d}_{\tau}-\bar{d}_{\tau}^{(-j)}|$.
Finally, the number of indices $k$ for which the indicator $\mathbf{1}\{d_{ik}\leq\bar{d}_{\tau}\}$ can differ from $\mathbf{1}\{d_{ik}^{(-j)}\leq\bar{d}_{\tau}^{(-j)}\}$ is bounded by the number of $k$ such that $d_{ik}$ lies within a band of width $O(1/n)$ around $\bar{d}_{\tau}$. The bounded density assumption implies that the expected number of such $k$ is $O(1)$, giving the last claim.
A full proof can be given by combining empirical process bounds for empirical quantiles with Assumption 2.1; we omit the details for brevity. ∎
Lemma B.3 (Consistency of $\widehat{V}_{ij}$).
Under Assumptions 2.1 and 2.2 and the dense regime convention, the plug‑in estimator $\widehat{V}_{ij}$ satisfies
$\max_{i,j}|\widehat{V}_{ij}-V_{ij}|=o_{\mathbb{P}}(1).$
Sketch.
Write $\widehat{V}_{ij}-V_{ij}=h^{-2}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(\widehat{P}_{kj}(1-\widehat{P}_{kj})-P_{kj}(1-P_{kj}))$. Use the uniform risk bounds for $\widehat{P}_{kj}$ (from Theorem 3.3) and the boundedness of $P_{kj}$ to conclude the result. ∎
Appendix C Proofs for Section 4
C.1 Proof of Lemma 4.1
Proof.
Fix $(h,\tau)$ and a row index $i$. For each $j\in T_{i}$, define
$Y_{j}:=(A_{ij}-\hat{P}_{ij}^{(-j)}(h,\tau))^{2}.$
  1. (a)
    Unbiasedness. Conditional on $P$, the LOO predictor $\hat{P}_{ij}^{(-j)}(h,\tau)$ is constructed from $A^{(-j)}$, which depends only on $\{A_{k\ell}:k\neq j,\ell\neq j\}$. By Proposition 2.7, $\hat{P}_{ij}^{(-j)}(h,\tau)$ is independent of $A_{ij}$ given $P$. Hence
    $\mathbb{E}\bigl[Y_{j}\mid\hat{P}_{ij}^{(-j)}(h,\tau),P\bigr]=\mathbb{E}\bigl[(A_{ij}-\hat{P}_{ij}^{(-j)}(h,\tau))^{2}\mid\hat{P}_{ij}^{(-j)}(h,\tau),P\bigr]$
    $\mathrm{Var}(A_{ij}\mid P)+(P_{ij}-\hat{P}_{ij}^{(-j)}(h,\tau))^{2}=P_{ij}(1-P_{ij})+(P_{ij}-\hat{P}_{ij}^{(-j)}(h,\tau))^{2}.$
    Averaging over $j\in T_{i}$ and dividing by $|T_{i}|$ yields $R_{i}(h,\tau)$ as defined in (4.2). Taking expectations gives $\mathbb{E}[\hat{R}_{i}(h,\tau)]=R_{i}(h,\tau)$.
  2. (b)
    Pointwise concentration. For fixed $(h,\tau)$ and $i$, the variables $\{Y_{j}:j\in T_{i}\}$ are bounded in $[0,1]$. Conditional on $P$ and the collection of LOO neighborhoods, the predictors $\hat{P}_{ij}^{(-j)}(h,\tau)$ are measurable with respect to $\{A_{k\ell}:k\neq j,\ell\neq j\}$ and hence independent of $A_{ij}$. For distinct $j,j^{\prime}\in T_{i}$, the pairs $(A_{ij},\hat{P}_{ij}^{(-j)})$ and $(A_{ij^{\prime}},\hat{P}_{ij^{\prime}}^{(-j^{\prime})})$ depend on disjoint sets of edges, so the $Y_{j}$ are conditionally independent and bounded.
    Hoeffding’s inequality therefore gives, for all $t>0$,
    $\mathbb{P}\Bigl(|\hat{R}_{i}(h,\tau)-R_{i}(h,\tau)|\geq t\,\Big|\,P\Bigr)\leq 2\exp(-2|T_{i}|t^{2}).$
    Choosing $t=C\sqrt{\log(1/\delta)/|T_{i}|}$ and integrating over $P$ yields the claimed bound.
  3. (c)
    Uniform concentration. Apply the pointwise bound to each $(h,\tau)\in\mathcal{G}$ and take a union bound:
    $\mathbb{P}\Bigl(\sup_{(h,\tau)\in\mathcal{G}}|\hat{R}_{i}(h,\tau)-R_{i}(h,\tau)|\geq t\Bigr)\leq 2|\mathcal{G}|\exp(-2|T_{i}|t^{2}).$
    Setting $t=C^{\prime}\sqrt{\log(|\mathcal{G}|/\delta)/|T_{i}|}$ yields the uniform deviation bound with probability at least $1-\delta$.
C.2 Proof of Theorem 4.2
Proof.
Let
$\Delta:=\sup_{(h,\tau)\in\mathcal{G}}|\hat{R}_{i}(h,\tau)-R_{i}(h,\tau)|.$
By Lemma 4.1(c), there exists $C^{\prime}>0$ such that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\Delta\leq C^{\prime}\sqrt{\frac{\log(|\mathcal{G}|/\delta)}{|T_{i}|}}.$
Let $(h^{*},\tau^{*})$ be a minimizer of $R_{i}(h,\tau)$ over $\mathcal{G}$. By the definition of $(\hat{h}_{i},\hat{\tau}_{i})$ and the bound on $\Delta$,
$R_{i}(\hat{h}_{i},\hat{\tau}_{i})\leq\hat{R}_{i}(\hat{h}_{i},\hat{\tau}_{i})+\Delta\leq\hat{R}_{i}(h^{*},\tau^{*})+\Delta\leq R_{i}(h^{*},\tau^{*})+2\Delta.$
Taking the infimum over $(h,\tau)\in\mathcal{G}$ in the last term and substituting the bound on $\Delta$ yields
$R_{i}(\hat{h}_{i},\hat{\tau}_{i})\leq\inf_{(h,\tau)\in\mathcal{G}}R_{i}(h,\tau)+2C^{\prime}\sqrt{\frac{\log(|\mathcal{G}|/\delta)}{|T_{i}|}}.$
Setting $C^{\prime\prime}=2C^{\prime}$ completes the proof. ∎
C.3 Auxiliary lemmas for Section 4
Lemma C.1 (Boundedness of squared prediction errors).
For any $(h,\tau)$ and $j\in T_{i}$, the squared prediction error $Y_{j}=(A_{ij}-\tilde{P}_{ij}^{(-j)}(h,\tau))^{2}$ satisfies $0\leq Y_{j}\leq 1$.
Proof.
Immediate since $A_{ij}\in\{0,1\}$ and $\tilde{P}_{ij}^{(-j)}\in[0,1]$. ∎
Lemma C.2 (Conditional independence formalization).
Under the LOO construction, for fixed $(h,\tau)$ and $j\in T_{i}$, $\tilde{P}_{ij}^{(-j)}(h,\tau)$ is measurable with respect to $\sigma(\{A_{k\ell}:k,\ell\neq j\})$ and hence independent of $A_{ij}$ conditional on $P$.
Proof.
Follows from the construction of $A^{(-j)}$ and the fact that $\mathcal{N}_{i}^{(-j)}(h,\tau)$ depends only on $A^{(-j)}$; see Proposition 2.7 and its proof. ∎
C.4 K‑fold CV and dependence considerations
Remark. K‑fold CV can be implemented but requires care: when a held‑out fold contains column $j$, neighborhoods used to predict entries in that fold must be constructed without using any columns from that fold to preserve the conditional independence exploited above. This increases computational cost (multiple neighborhood constructions) and slightly reduces the effective training size; the same concentration arguments apply after adjusting $|T_{i}|$ to the validation fold size.
Appendix D Implementation details and computational complexity
This appendix records implementation notes for Algorithm 1 and complexity estimates. (Optional: include incremental update formulas for $(A^{(-j)})^{2}$, vectorized computations, and memory/time complexity.)
Proof of Theorem 3.3.
Fix $i$ and write
$\hat{P}_{ij}-P_{ij}=V_{ij}+B_{ij},$
where $V_{ij}$ and $B_{ij}$ are as in (3.1)–(3.2). Then
$R_{i}:=\frac{1}{n}\sum_{j=1}^{n}(\hat{P}_{ij}-P_{ij})^{2}\leq\frac{2}{n}\sum_{j=1}^{n}V_{ij}^{2}+\frac{2}{n}\sum_{j=1}^{n}B_{ij}^{2}.$
We first control the variance term. By Lemma 3.1 with $t=C_{0}\log n$ and a union bound over $(i,j)$, there exist constants $C_{1},c_{1}>0$ such that with probability at least $1-n^{-c_{1}}$,
$\max_{i,j}|V_{ij}|\leq\sqrt{2V_{ij}C_{0}\log n}+\frac{C_{0}\log n}{3h}\lesssim\sqrt{\frac{\log n}{n}},$
using $V_{ij}\leq 1/(4h)$ and $h\asymp\sqrt{n\log n}$. Hence
$\frac{1}{n}\sum_{j=1}^{n}V_{ij}^{2}\leq\max_{j}V_{ij}^{2}\lesssim\frac{\log n}{n}.$
Next, by Lemma 3.2 there exist constants $C_{2},c_{2}>0$ such that with probability at least $1-n^{-c_{2}}$,
$\max_{i,j}|B_{ij}|\leq C_{2}\Bigl(\frac{\log n}{n}\Bigr)^{1/4}.$
Therefore
$\frac{1}{n}\sum_{j=1}^{n}B_{ij}^{2}\leq\max_{j}B_{ij}^{2}\lesssim\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Combining the two bounds and adjusting constants yields
$R_{i}\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}$
for all $i$ on an event of probability at least $1-n^{-c}$, where $c=\min(c_{1},c_{2})$. ∎
Computational shortcut for LOO neighborhoods
To avoid recomputing neighborhoods from scratch for every held‑out column $j$, we precompute the full‑data neighborhoods $\mathcal{N}_{i}$ and, for each $i$, a short candidate list of the next $K$ nearest indices in the ZLZ distance. For a given $j$, only those rows $i$ with $j\in\mathcal{N}_{i}$ require any update; in practice this set is a small fraction of all rows. For such $i$, we first attempt to repair the neighborhood by scanning the candidate list and selecting the first index that satisfies the LOO distance criterion. If no candidate qualifies, we fall back to an exact recomputation of the $h$ nearest neighbors under $A^{(-j)}$. Empirically this shortcut yields substantial reductions in wall‑clock time while producing exact LOO neighborhoods whenever the fallback is invoked. Section 5 reports diagnostics including the average fraction of affected rows per held‑out column and the fallback rate.
Appendix E Proof of Theorem 3.5
We prove the Berry-Esseen bound for the standardized variance term $U_{ij}/\sqrt{V_{ij}}$, conditional on $(\mathcal{N}_{i}^{(-j)},P)$. Fix $i$ and $j$ and condition on $(\mathcal{N}_{i}^{(-j)},P)$. Define
$X_{k}:=A_{kj}-P_{kj},\qquad k\in\mathcal{N}_{i}^{(-j)}.$
By Proposition 2.7, the $\{X_{k}\}$ are independent, mean-zero, and bounded in [1,1][-1,1]. Their conditional variances are
$\mathrm{Var}(X_{k}\mid\mathcal{N}_{i}^{(-j)},P)=\mathrm{Var}(A_{kj}\mid P)=P_{kj}(1-P_{kj}),$
so the total variance is
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(X_{k}\mid\mathcal{N}_{i}^{(-j)},P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}).$
By definition, $V_{ij}=h^{-2}\sigma_{ij}^{2}$, and
$U_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k}=\frac{1}{h}S,$
where $S:=\sum_{k}X_{k}$. Next, define the third absolute moment sum
$\rho_{3}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathbb{E}\big[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P\big].$
Since $|X_{k}|\leq 1$, we have $|X_{k}|^{3}\leq|X_{k}|^{2}$ and hence
$\mathbb{E}\big[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P\big]\leq\mathbb{E}\big[X_{k}^{2}\mid\mathcal{N}_{i}^{(-j)},P\big]=\mathrm{Var}(X_{k}\mid\mathcal{N}_{i}^{(-j)},P).$
Summing over $k$ yields
$\rho_{3}\leq\sum_{k}\mathrm{Var}(X_{k}\mid\mathcal{N}_{i}^{(-j)},P)=\sigma_{ij}^{2}.$
The classical Berry-Esseen theorem for sums of independent, not necessarily identically distributed random variables (Vershynin, 2018; Petrov, 1975, Theorem 5.5) states that there exists a universal constant $C>0$ such that
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\Big(\tfrac{S}{\sigma_{ij}}\leq x\,\Big|\,\mathcal{N}_{i}^{(-j)},P\Big)-\Phi(x)\Big|\leq C\frac{\rho_{3}}{\sigma_{ij}^{3}}.$
Combining this with $\rho_{3}\leq\sigma_{ij}^{2}$ gives
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\Big(\tfrac{S}{\sigma_{ij}}\leq x\,\Big|\,\mathcal{N}_{i}^{(-j)},P\Big)-\Phi(x)\Big|\leq C\frac{\sigma_{ij}^{2}}{\sigma_{ij}^{3}}=C\sigma_{ij}^{-1}.$
We now control $\sigma_{ij}$ using Assumption 2.2. By Assumption 2.2, there exists $\eta\in(0,1/2)$ such that $\eta\leq P_{kj}\leq 1-\eta$ for all $k,j$, and hence
$\eta(1-\eta)\leq P_{kj}(1-P_{kj})\leq\tfrac{1}{4}\qquad\text{for all }k,j.$
Since $|\mathcal{N}_{i}^{(-j)}|=h$ by construction, it follows that
$\eta(1-\eta)\,h\leq\sigma_{ij}^{2}=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj})\leq\tfrac{1}{4}h$
for all $(i,j)$. In particular, there exist constants $0<c_{1}\leq c_{2}<\infty$ such that
$c_{1}h\leq\sigma_{ij}^{2}\leq c_{2}h$
uniformly over $(i,j)$, and hence
$\sigma_{ij}^{-1}\leq C^{\prime}h^{-1/2}$
for some universal constant $C^{\prime}>0$. Therefore,
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\Big(\tfrac{S}{\sigma_{ij}}\leq x\,\Big|\,\mathcal{N}_{i}^{(-j)},P\Big)-\Phi(x)\Big|\leq C^{\prime}h^{-1/2}.$
Finally, we relate this to $U_{ij}$ and $V_{ij}$. Since $U_{ij}=h^{-1}S$ and $V_{ij}=h^{-2}\sigma_{ij}^{2}$, we have
$\frac{U_{ij}}{\sqrt{V_{ij}}}=\frac{h^{-1}S}{\sqrt{h^{-2}\sigma_{ij}^{2}}}=\frac{S}{\sigma_{ij}}.$
Therefore the same bound holds for $U_{ij}/\sqrt{V_{ij}}$, that is,
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\Big(\tfrac{U_{ij}}{\sqrt{V_{ij}}}\leq x\,\Big|\,\mathcal{N}_{i}^{(-j)},P\Big)-\Phi(x)\Big|\leq Ch^{-1/2},$
for an appropriate constant $C>0$, which is exactly (3.5). ∎