### v3

# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

Abstract
Estimating latent edge probabilities in network models often relies on aggregating information from structurally similar nodes. However, using the same adjacency matrix to both determine node similarities and estimate probabilities induces complex statistical dependencies. This coupling invalidates standard inferential tools, such as concentration inequalities and central limit theorems, leaving a critical gap in entrywise uncertainty quantification for networks. To resolve this, we introduce a leave-one-out (LOO) decoupling strategy. When predicting a target edge, the local neighborhood is constructed from a reduced adjacency matrix with the target row and column entirely removed. This structural modification-rather than the specific smoothing mechanism-is the core innovation: it restores exact conditional independence between the data-driven neighborhood selection and the observed edges being aggregated. Leveraging this restored independence, we derive variance-adaptive empirical Bernstein bounds and establish uniform Berry-Esseen normal approximations with explicit rates. These results enable the construction of valid finite-sample and asymptotic confidence intervals for individual edge probabilities. Furthermore, the LOO architecture naturally admits an honest cross-validation scheme for tuning parameter selection, providing a complete and rigorous inferential framework while preserving strong mean-squared error performance.
1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2007; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $h$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, modern estimators-including universal singular value decomposition (Chatterjee, 2015), sorting-based smoothing (Chan and Airoldi, 2014), and block-approximations (Gao et al., 2015)-have been proven to reliably achieve minimax-optimal rates for point estimation.
Among these, the neighborhood smoothing method introduced by Zhang et al. (2017) stands out. By adaptively aggregating structurally similar nodes based on empirical two-hop distances, neighborhood smoothing achieves optimal mean-squared error (MSE) without requiring the strict structural assumptions of block models or the selection of a global low-rank embedding.
However, a fundamental gap remains: the transition from optimal point estimation to rigorous statistical inference. Given an entrywise estimate $\widehat{P}_{ij}$, it is critically important for downstream scientific tasks to attach a valid measure of uncertainty to this prediction. Yet, existing inferential theory for networks has focused almost exclusively on low-dimensional or global functionals. Significant progress has been made in proving limit theorems for spectral embeddings (Sussman et al., 2012; Athreya et al., 2016), asymptotic normality for community parameters in SBMs (Bickel et al., 2013), and global goodness-of-fit testing (Lei, 2016). Conversely, entrywise uncertainty quantification-specifically, constructing valid confidence intervals and proving central limit theorems for individual edge probabilities under continuous graphons-has remained theoretically elusive. Recent attempts to bridge this gap have relied on computationally intensive resampling methods, such as network bootstrapping (Levin and Levina, 2025; Green and Shalizi, 2022), or generalized fiducial inference (Su et al., 2022), but exact distribution-free analytical bounds for neighborhood estimators are lacking.
The obstacle preventing entrywise inference is not merely technical, but structural. To achieve optimal MSE, neighborhood smoothing estimators (Zhang et al., 2017) rely on a data-driven aggregation of similar nodes. This creates a severe “double-dipping” problem: the procedure reuses the adjacency matrix twice. The matrix $A$ is first used to compute pairwise topological distances to select a neighborhood $\mathcal{N}_{i}$, and then the very same matrix $A$ is used to average the edges within that neighborhood. Consequently, the random neighborhood weights depend directly on the edges being averaged. Even after conditioning on the latent probability matrix $P$, the summands in the estimator are strongly coupled. This intractable dependence invalidates the foundational assumptions of classical concentration inequalities and Berry-Esseen normal approximations.
In this paper, we demonstrate that this probabilistic bottleneck can be cleanly bypassed via a simple but highly principled structural modification: leave-one-out (LOO) neighborhood smoothing.
When estimating a single entry $P_{ij}$, we perform a topological surgery on the graph by constructing the neighborhood of node $i$ using a reduced adjacency matrix $A^{(-j)}$ in which the $j$-th row and column have been entirely deleted. All pairwise distances and nearest-neighbor selections are computed on this reduced matrix. Only after the neighborhood $\mathcal{N}_{i}^{(-j)}$ is fixed do we average the original target edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$.
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling bridges the gap between network estimation and classical empirical process theory. Specifically, under standard piecewise Lipschitz assumptions on the graphon, we leverage this restored independence to establish three main contributions:
  1. 1.
    Variance-Adaptive Concentration: We derive tight, finite-sample concentration inequalities based on empirical Bernstein bounds that automatically adapt to the local variance of the edge probabilities.
  2. 2.
    Asymptotic Normality: We establish uniform Berry-Esseen bounds with explicit convergence rates for the normalized estimation error, enabling the construction of valid normal-approximation confidence intervals for individual edge probabilities.
  3. 3.
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$.inequality.
Crucially, these inferential guarantees do not come at the cost of statistical efficiency. We prove that the proposed LOO estimator preserves the optimal row-wise mean-squared error rates of classical neighborhood smoothing.
The remainder of the paper is organized as follows. Section 2 formalizes the graphon model, the Zhang-Levina-Zhu (ZLZ), Zhang et al. (2017) empirical distance, and our leave-one-out neighborhood construction. Section 3.2 establishes the foundational conditional independence, proving the variance-adaptive concentration bounds, Berry-Esseen normal approximations, and the validity of the resulting confidence intervals. Section 4 details the honest cross-validation scheme for tuning parameter selection. Section 5 reports comprehensive simulation results illustrating the finite-sample coverage and MSE performance, and Section 6 concludes with a discussion of extensions to directed, weighted, and degree-corrected settings.
2 Model and Leave-One-Out Neighborhood Smoothing
We observe a single undirected simple graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$, no self‑loops ($A_{ii}=0$), and independent upper‑triangular entries
$A_{ij}\mid P_{ij}\ \sim\ \mathrm{Bernoulli}(P_{ij}),\qquad 1\leq i<j\leq n,$ (2.1)
with $A_{ij}=A_{ji}$. The matrix $P=(P_{ij})$ contains unknown edge probabilities. We adopt the standard graphon parametrization below.
Assumption 2.1 (Graphon model and smoothness).
The edge probabilities satisfy $P_{ij}=h(\xi_{i},\xi_{j})$ with $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$. The graphon $h:[0,1]^{2}\to[0,1]$ is piecewise Lipschitz in each argument with Lipschitz constant $L$ on a fixed finite rectangular partition of $0,1]^{2}$.
Assumption 2.2 (Non‑degeneracy).
There exists $\eta\in(0,1/2)$ such that
$\eta\leq h(u,v)\leq 1-\eta\qquad\text{for all }(u,v)\in[0,1]^{2}.$
Remark 2.3.
Assumption 2.2 requires that there exists $\eta\in(0,1/2)$ such that $\eta\leq h(u,v)\leq 1-\eta$ for all $(u,v)\in[0,1]^{2}$. This nondegeneracy condition guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance‑adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re‑normalization and additional control of rare events, which we leave to future work. Empirically, moderate thinning of edges does not invalidate the procedure (see Section 5).
Remark 2.4.
The rectangular partition in Assumption 2.1 is fixed and does not grow with $n$. All constants in our rates may depend on the Lipschitz constant $L$ and on the partition, but not on $n$.
Remark 2.5 (Dense regime convention).
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=h(\xi_{i},\xi_{j})$ with $h$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two‑hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}h(\xi_{i},\xi_{j})$ are discussed briefly in Section 6.
2.1 LOO neighborhoods and estimator
Our estimator is based on an adaptive neighborhood constructed through a leave-one-out (LOO) modification of the Zhang-Levina-Zhu (ZLZ) distance. For $i\neq j$, define the empirical second‑order (ZLZ) distance (Zhang et al., 2017)
$d_{ij}:=\max_{k\neq i,j}\Big|\big(A^{2}/n\big)_{ik}-\big(A^{2}/n\big)_{jk}\Big|.$ (2.2)
This is the sup‑norm discrepancy between the $i$th and $j$th rows of the normalized two‑hop matrix $A^{2}/n$, measuring the similarity of their two‑hop neighborhoods.
For estimation of a fixed entry $(i,j)$, remove node $j$ and form the reduced adjacency matrix $A^{(-j)}$ obtained by deleting the $j$th row and column. Define the two-hop matrix
$M^{(-j)}=\frac{1}{n-1}A^{(-j)}A^{(-j)}.$ (2.3)
For node $i$ and node $k\neq i,j$, define the ZLZ distance as follows:
[…]
Let
$d_{i,(1)}^{(-j)}\leq\cdots\leq d_{i,(n-2)}^{(-j)},$
denote the order statistics of all nodes except $j$ to node $i$, namely $\{d_{ik}^{(-j)}:k\neq i,j\}$. Define the leave-one-out neighborhood of node $i$ by
$\mathcal{N}_{i}^{(-j)}=\left\{k\neq i,j:d_{ik}^{(-j)}\leq d_{i,(h_{n})}^{(-j)}\right\},$ (2.5)
that is, the $h_{n}$ nearest neighbors of $i$ according to the ZLZ two-hop distance computed on the reduced graph.
This construction preserves the original ZLZ neighborhood rule, with the sole modification that node $j$ is removed prior to distance computation.
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry-Esseen theory.
Given $\mathcal{N}_{i}^{(-j)}$, define the one‑sided LOO smoother and its symmetrized version by
$\tilde{P}_{ij}:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},\qquad\widehat{P}_{ij}:=\frac{\tilde{P}_{ij}+\tilde{P}_{ji}}{2}.$ (2.6)
Remark 2.6.
The symmetrized estimator $\widehat{P}_{ij}$ respects the symmetry of the simple graph model. All concentration and normal‑approximation results are derived for the one‑sided quantities $\tilde{P}_{ij}$; symmetrization does not affect the independence structure.
Proposition 2.7 (Conditional independence).
Conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$, the variables
$\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$
are independent, mean-zero, and bounded in [1,1][-1,1].
Proof.
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=h(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$.
By construction, $A^{(-j)}$ is obtained from $A$ by removing the $j$th row and column. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$.
For each $k\in\mathcal{N}_{i}^{(-j)}$, we have $A_{kj}\sim\mathrm{Bernoulli}(P_{kj})$ and $A_{kj}-P_{kj}\in[-1,1]$ with mean zero. Since the edges $\{A_{kj}:k\neq j\}$ are independent conditional on $P$, and $\mathcal{N}_{i}^{(-j)}$ is independent of these edges, it follows that the collection $\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ is independent, mean-zero, and bounded in [1,1][-1,1] conditional on $(\mathcal{N}_{i}^{(-j)},P)$. ∎
3 Consistency
In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write For $i\neq j$, write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}+B_{ij},$
where
$U_{ij}^{(-j)}$ $:=\tilde{P}_{ij}-\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.1)
$B_{ij}^{(-j)}$ $:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}-P_{ij}.$ (3.2)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}^{(-j)}$ are independent, mean zero and bounded, so $U_{ij}^{(-j)}$ behaves like an average of $h$ independent observations. The term $B_{ij}^{(-j)}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2.
3.1 Neighborhood bias
The bias $B_{ij}$ captures the approximation error from replacing $P_{ij}$ by the local average of $P_{kj}$ over the LOO neighborhood. Under the piecewise Lipschitz graphon assumption and the ZLZ neighborhood construction, neighbors in $\mathcal{N}_{i}^{(-j)}$ are close in latent position to $\xi_{i}$, which yields a small bias.
Lemma 3.1 (Uniform two-hop stability under deletion of one vertex).
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of an undirected simple graph with $A_{ii}=0$. For each $j\in[n]$, let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting the $j$th row and column (i.e., removing vertex $j$). Define the normalized two-hop matrices
$M:=\frac{1}{n}A^{2},\qquad M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}.$
For each $j$, embed $M^{(-j)}$ into an $n\times n$ matrix by placing it on the index set $n]\setminus\{j\}$ and setting entries involving index $j$ to zero. Then, deterministically,
$\sup_{1\leq j\leq n}\ \|M-M^{(-j)}\|_{\max}\ \leq\ \frac{3}{n},$
where $\|B\|_{\max}:=\max_{1\leq i,\ell\leq n}|B_{i\ell}|$ denotes the entrywise max norm. In particular, for every $j\in[n]$ and all $i,\ell\in[n]\setminus\{j\}$,
$\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{3}{n}.$
Proof.
Fix $j\in[n]$ and $i,\ell\in[n]\setminus\{j\}$. Expanding the $(i,\ell)$ entry gives
$(A^{2})_{i\ell}=\sum_{m=1}^{n}A_{im}A_{m\ell},\qquad\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell},$
hence
$(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}=A_{ij}A_{j\ell},$
so $\big|(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}\big|\leq 1$. Therefore,
[…]
$\leq\left|\frac{1}{n}(A^{2})_{i\ell}-\frac{1}{n}\big(A^{(-j)}\big)^{2}_{i\ell}\right|+\left|\frac{1}{n}-\frac{1}{n-1}\right|\big(A^{(-j)}\big)^{2}_{i\ell}$
$\leq\frac{1}{n}+\frac{1}{n(n-1)}\big(A^{(-j)}\big)^{2}_{i\ell}.$
Since $\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell}\leq n-1$, we get
$\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{1}{n}+\frac{n-1}{n(n-1)}=\frac{2}{n}.$
Thus the bound holds for every $j$ (in fact with constant $2$), and taking the supremum over $j$ yields $\|M-M^{(-j)}\|_{\max,[n]\setminus\{j\}}:=\max_{i,\ell\in[n]\setminus\{j\}}|M_{i\ell}-M^{(-j)}_{i\ell}|\leq 2/n$. ∎
Lemma 3.2 (Two-hop concentration is preserved under leave-one-out deletion).
Let $A\in\{0,1\}^{n\times n}$ be generated from the graphon model (2.1) with probability matrix $P$. Define
$M:=\frac{1}{n}A^{2},\qquad M^{\star}:=\mathbb{E}\!\left[M\,\middle|\,P\right],$
and for each $j\in[n]$ let $A^{(-j)}$ be the adjacency matrix obtained by deleting the $j$th row and column. Define the leave-one-out two-hop matrices
$M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2},\qquad M^{(-j)\star}:=\mathbb{E}\!\left[M^{(-j)}\,\middle|\,P\right].$
Embed $M^{(-j)}$ and $M^{(-j)\star}$ into $n\times n$ matrices on the index set $n]\setminus\{j\}$ (and set entries involving $j$ to zero).
Then, for every $n$,
$\sup_{1\leq j\leq n}\ \big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}\ \leq\ \|M-M^{\star}\|_{\max}\ +\ \frac{C}{n},$ (3.3)
for a universal constant $C>0$. In particular, if
$\|M-M^{\star}\|_{\max}=\mathcal{O}_{\mathbb{P}}\!\Big(\sqrt{\tfrac{\log n}{n}}\Big),$
(as established for the ZLZ two-hop matrix in Zhang et al. (2017) and its supplementary material), then
$\sup_{1\leq j\leq n}\ \big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}=\mathcal{O}_{\mathbb{P}}\!\Big(\sqrt{\tfrac{\log n}{n}}\Big).$
Proof.
Fix $j\in[n]$. By the triangle inequality,
$\big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}$ $\leq\big\|M^{(-j)}-M\big\|_{\max}+\big\|M-M^{\star}\big\|_{\max}+\big\|M^{\star}-M^{(-j)\star}\big\|_{\max}.$
The first term is deterministic and satisfies
$\sup_{1\leq j\leq n}\|M^{(-j)}-M\|_{\max}\leq\frac{C_{1}}{n}$
by Lemma 3.1. For the last term, note that
$M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)\star}=\mathbb{E}[M^{(-j)}\mid P],$
so Jensen’s inequality and Lemma 3.1 give
$\sup_{1\leq j\leq n}\|M^{\star}-M^{(-j)\star}\|_{\max}=\sup_{1\leq j\leq n}\big\|\mathbb{E}[M-M^{(-j)}\mid P]\big\|_{\max}\leq\sup_{1\leq j\leq n}\mathbb{E}\big[\|M-M^{(-j)}\|_{\max}\mid P\big]\leq\frac{C_{1}}{n}.$
Combining the bounds yields (3.3) with $C=2C_{1}$. Finally, since $n^{-1}=o\!\big(\sqrt{(\log n)/n}\big)$, the stated rate transfer follows immediately. ∎
Proposition 3.3 (ZLZ neighborhood smoothing: row-risk (black box)).
Work under the assumptions of Zhang, Levina, and Zhu (2017, Biometrika) and choose the bandwidth $h\asymp\sqrt{n\log n}$ as in their Theorem 1. Let $N_{i}$ be the ZLZ neighborhood of node $i$ constructed from the two-hop matrix $M=A^{2}/n$ via their distance and thresholding rule, and define the ZLZ estimator
$\tilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|N_{i}|}\sum_{k\in N_{i}}A_{kj}.$
Then there exist constants $C,c>0$ such that with probability at least $1-n^{-c}$, uniformly over $i$,
$\frac{1}{n}\sum_{j\neq i}\big(\tilde{P}^{\mathrm{ZLZ}}_{ij}-P_{ij}\big)^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$
Proof.
This is the row-wise conclusion proved for $\tilde{P}_{ij}$ in Zhang et al. (2017), using their supplementary decomposition into $J_{1}(i,j)$ and $J_{2}(i,j)$ (equation 1 in the supplementary material of Zhang et al. (2017)) and the two-hop concentration input. ∎
Corollary 3.4 (Leave-one-out ZLZ row-wise risk + $(U_{ij},B_{ij})$ decomposition).
Fix $j\in[n]$. Let $A^{(-j)}$ be the adjacency matrix with row/column $j$ deleted, and let $\widetilde{P}^{(-j)}$ denote the ZLZ neighborhood-smoothing estimator computed from $A^{(-j)}$ (with neighborhoods $N_{i}^{(-j)}$ defined exactly as in ZLZ but using the two-hop matrix $(A^{(-j)})^{2}/(n-1)$). For $i\neq j$ and $k\in N_{i}^{(-j)}$, define
$U^{(-j)}_{ij}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(A_{kj}-P_{kj}),\qquad B^{(-j)}_{ij}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(P_{kj}-P_{ij}),$
so that
$\widetilde{P}^{(-j)}_{ij}-P_{ij}=U^{(-j)}_{ij}+B^{(-j)}_{ij}$.
Then for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\big)^{2}\;\leq\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;+\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2},$
and, with probability at least $1-c_{1}n^{-c_{2}}$ (for the same type of constants as ZLZ), uniformly for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\big)^{2}\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
Moreover, identifying ZLZ’s terms by
$J_{1}^{(-j)}(i,\ell):=\big(U^{(-j)}_{i\ell}\big)^{2},\qquad J_{2}^{(-j)}(i,\ell):=\big(B^{(-j)}_{i\ell}\big)^{2},$
one has the ZLZ-style bounds (with (n1)(n-1) in place of $n$):
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2},\qquad\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
Proof.
First, (ZLZ algebra; defines $J_{1},J_{2}$ and matches $U,B$). Exactly as in ZLZ, we begin from the decomposition of the squared error into two contributions:
$\Big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\Big)^{2}=\Bigg[\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\Big\{(A_{k\ell}-P_{k\ell})+(P_{k\ell}-P_{i\ell})\Big\}\Bigg]^{2},$
and the inequality
$(\cdot)^{2}\leq 2J_{1}^{(-j)}(i,\ell)+2J_{2}^{(-j)}(i,\ell)$
is the same inequality as ZLZ equation (1), with the identifications $J_{1}^{(-j)}(i,\ell)=(U^{(-j)}_{i\ell})^{2}$ and $J_{2}^{(-j)}(i,\ell)=(B^{(-j)}_{i\ell})^{2}$. (See equation (1), supplementary material of Zhang et al. (2017) .)
Second, two-hop concentration under deletion preserves ZLZ inputs. ZLZ’s neighborhood construction and the control of $J_{2}$ rely on uniform control of two-hop quantities of the form $(A^{2}/n)_{ab}-(P^{2}/n)_{ab}$, established in ZLZ equations, (4)-(6) in their supplementary material. By Lemma 3.2, the same bound holds when replacing $A$ by $A^{(-j)}$ (and $n$ by n1n-1). Therefore, every subsequent ZLZ step that uses (4)-(6) remains valid for the graph with the deleted vertex.
Third, Bounding the $J_{1}$ row-average ZLZ does not bound each $J_{1}(i,\ell)$ separately; instead it controls the row-average $(n^{-1}\sum_{\ell}J_{1}(i,\ell))$ via the decomposition in equation (11), together with the bounds (12), (14), and the conclusion (15) in the ZLZ supplementary material. The same argument applies verbatim to $A^{(-j)}$ and $N_{i}^{(-j)}$, using equation 3.2 to justify the same concentration inputs. Hence
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2},$
by the ZLZ chain (11)-(15).
Fourth, ZLZ bounds $(n^{-1}\sum_{\ell}J_{2}(i,\ell))$ using Cauchy-Schwarz and the “claim 2 of their Lemma 2” input, concluding in equation (16). The same reasoning applies here (again by Lemma 3.2 to justify the same concentration inputs), yielding the same order bound:
$\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
This is exactly ZLZ equation (16) with the leave-one-out objects.
Finally, we have the same claimed row-wise risk bound, matching ZLZ’s stated goal (their inequality right after introducing $\widetilde{P}$, i.e., the display in the supplementary around the start of the proof of Theorem 1). ∎
Assumption 3.5 (Piecewise bi-Lipschitz smoothness).
The graphon $W:[0,1]^{2}\to[0,1]$ admits a finite vertical partition $0=b_{0}<b_{1}<\cdots<b_{K}=1$ with blocks $I_{m}:=(b_{m},b_{m+1})$ such that:
  1. 1.
    (Within-block Lipschitz in the first argument.) For all $m$ and all $u,u^{\prime}\in I_{m}$, $x\in[0,1]$,
    $|W(u,x)-W(u^{\prime},x)|\leq L|u-u^{\prime}|.$
  2. 2.
    (Across-block $L^{2}$ separation.) There exist $\eta,\delta>0$ such that for any $u\in I_{m}$ and $u^{\prime}\in I_{m^{\prime}}$ with $m\neq m^{\prime}$,
    $\lambda\big(\{x\in[0,1]:|W(u,x)-W(u^{\prime},x)|\geq\eta\}\big)\geq\delta.$
  3. 3.
    (Within-block $L^{2}$ identifiability.) For any $\varepsilon>0$ with $b_{m}+\varepsilon<b_{m+1}-\varepsilon$, there exists $\kappa(\varepsilon)>0$ such that for all $u,u^{\prime}\in[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$,
    $\int_{0}^{1}(W(u,x)-W(u^{\prime},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|u-u^{\prime}|^{2}.$
Lemma 3.6 (Rowwise $L^{2}$ closeness implies latent closeness).
Assume Assumption 3.5. Fix $\varepsilon>0$ and define
$\mathcal{I}_{\varepsilon}=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,b_{m+1}-\varepsilon].$
Let $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$ and set $P_{ab}=W(U_{a},U_{b})$. Then there exist constants $c_{0},c_{1}>0$ (depending only on the constants in Assumption 3.5 and on $\varepsilon$) such that, for all sufficiently large $n$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;c_{0}\Big(\frac{\log n}{n}\Big)^{1/2}\quad\Longrightarrow\quad\begin{cases}U_{i},U_{k}\text{ lie in the same block }I_{m},\\ |U_{i}-U_{k}|\leq c_{1}(\log n/n)^{1/4}.\end{cases}$
Proof.
Write $f_{i}(x)=W(U_{i},x)$ and $f_{k}(x)=W(U_{k},x)$.
We can easily see that if $U_{i}$ and $U_{k}$ lie in different blocks, then by Assumption 3.5(ii),
$\int_{0}^{1}(f_{i}-f_{k})^{2}\;\geq\;\eta^{2}\delta.$
For $c_{0}>0$ small enough and $n$ large, $c_{0}(\log n/n)^{1/2}<\eta^{2}\delta/2$, so the assumed upper bound on $\frac{1}{n}\sum_{\ell}(P_{i\ell}-P_{k\ell})^{2}$ (which concentrates around $\int_{0}^{1}(f_{i}-f_{k})^{2}$ conditional on $(U_{\ell})$) is incompatible with this. Hence $U_{i},U_{k}$ must lie in the same block $I_{m}$.
On the other hand, since $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$, we have $U_{i},U_{k}\in[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$ for some $m$, and Assumption 3.5(iii) gives
$\int_{0}^{1}(f_{i}-f_{k})^{2}=\int_{0}^{1}(W(U_{i},x)-W(U_{k},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|U_{i}-U_{k}|^{2}.$
Combining this with the assumed upper bound yields
$\kappa(\varepsilon)\,|U_{i}-U_{k}|^{2}\;\leq\;c_{0}\Big(\frac{\log n}{n}\Big)^{1/2},$
so
$|U_{i}-U_{k}|\;\leq\;\sqrt{\frac{c_{0}}{\kappa(\varepsilon)}}\,\Big(\frac{\log n}{n}\Big)^{1/4}.$
Set $c_{1}:=\sqrt{c_{0}/\kappa(\varepsilon)}$. ∎
Lemma 3.7 (Bias bound for the LOO neighborhood smoother).
Assume Assumption 3.5. Let $h_{n}\asymp\sqrt{n\log n}$ and define $r_{n}=(\log n/n)^{1/4}$. For $i\neq j$, set
$B^{(-j)}_{ij}=\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}),\qquad P_{ab}=W(U_{a},U_{b}).$
Suppose there exist constants $C_{2},c>0$ such that, with probability at least $1-n^{-c}$, the LOO neighborhoods satisfy
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C_{2}\Big(\frac{\log n}{n}\Big)^{1/2}\qquad\text{for all }i\neq j\text{ and all }k\in\mathcal{N}_{i}^{(-j)}.$ (3.4)
Then for all $i\neq j$ with $U_{i},U_{j}\in\mathcal{I}_{\varepsilon}$,
$|B^{(-j)}_{ij}|\;\leq\;C\,r_{n},$
for some constant $C>0$ depending only on $(L,\eta,\delta,\kappa(\varepsilon),C_{2})$.
Proof.
Fix $i\neq j$ with $U_{i},U_{j}\in\mathcal{I}_{\varepsilon}$ and work on the event where (3.4) holds. Define
$\mathcal{B}_{i}=\Big\{k\in\mathcal{N}_{i}^{(-j)}:\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}>M\Big(\frac{\log n}{n}\Big)^{1/2}\Big\},$
for some constant $M>C_{2}$.
Averaging $(\star)$ over $k\in\mathcal{N}_{i}^{(-j)}$ gives
$\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C_{2}\Big(\frac{\log n}{n}\Big)^{1/2}.$
If $|\mathcal{B}_{i}|\geq 1$, then the left-hand side is at least
$\frac{|\mathcal{B}_{i}|}{|\mathcal{N}_{i}^{(-j)}|}M\Big(\frac{\log n}{n}\Big)^{1/2}\;\geq\;\frac{M}{|\mathcal{N}_{i}^{(-j)}|}\Big(\frac{\log n}{n}\Big)^{1/2},$
which is strictly larger than $C_{2}(\log n/n)^{1/2}$ for $M>C_{2}$ and $|\mathcal{N}_{i}^{(-j)}|\geq 1$, a contradiction. Hence $|\mathcal{B}_{i}|=0$, and for every $k\in\mathcal{N}_{i}^{(-j)}$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;M\Big(\frac{\log n}{n}\Big)^{1/2}.$
Applying Lemma 3.6 to each such $k$ yields that $U_{k}$ lies in the same block as $U_{i}$ and
$|U_{k}-U_{i}|\leq C^{\prime}r_{n}$
for some constant $C^{\prime}>0$. Since $U_{j}\in\mathcal{I}_{\varepsilon}$, the points $(U_{i},U_{j})$ and $(U_{k},U_{j})$ lie in the same Lipschitz region of $W$ in the first argument, so by Assumption 3.5(i),
$|P_{kj}-P_{ij}|=|W(U_{k},U_{j})-W(U_{i},U_{j})|\leq L|U_{k}-U_{i}|\leq LC^{\prime}r_{n}.$
Averaging over $k\in\mathcal{N}_{i}^{(-j)}$ gives
$|B^{(-j)}_{ij}|\;\leq\;Cr_{n}$
for some $C>0$, as claimed. ∎
Lemma 3.7 establishes a bias bound for the leave-one-out neighborhood smoother, showing that under the piecewise bi-Lipschitz smoothness assumption, the bias term $B_{ij}^{(-j)}$ is of order $r_{n}=(\log n/n)^{1/4}$ for nodes $i$ and $j$ whose latent positions lie in the interior of the same block. This result relies on the control of the average two-hop distance between node $i$ and its neighbors in the leave-one-out neighborhood, as well as on the Lipschitz continuity of the graphon within blocks.
3.2 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
We introduce the conditional variance for $U_{ij}^{(-j)}$:
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.5)
so that $V_{ij}\leq 1/(4h)$. While $V_{ij}$ is analytically useful and measurable with respect to the $\sigma$-field generated by $(\mathcal{N}_{i}^{(-j)},P)$, it depends on the unknown latent probabilities $\{P_{kj}\}$.
To obtain tight, variance-adaptive control of $U_{ij}^{(-j)}$ using only observable quantities, we define the unbiased sample variance of the edges in the leave-one-out neighborhood:
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h(h-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h}{h-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}),$ (3.6)
where $\tilde{P}_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical mean of the observed edges. Instead of relying on classical Bernstein bounds that require the true variance $V_{ij}$, we utilize the Empirical Bernstein inequality (Theorem 4 of Maurer and Pontil (2009)), which directly incorporates the sample variance $\hat{\sigma}_{ij}^{2}$. The key advantage of this approach is that it completely bypasses the need for asymptotic plug-iin justifications, providing a strictly finite-sample, variance-adaptive control of $U_{ij}^{(-j)}$ based entirely on the observable quantity $\hat{\sigma}_{ij}^{2}$.
Lemma 3.8 (Empirical Bernstein inequality).
For any $t>0$,
$\mathbb{P}\Bigg(\big|U_{ij}^{(-j)}\big|\geq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\,t}{h}}+\frac{7t}{3(h-1)}\ \Bigg|\ \mathcal{N}_{i}^{(-j)},P\Bigg)\leq 2e^{-t}.$
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Their empirical mean is
$\tilde{P}_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},$
and its conditional expectation is
$\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}.$
Thus $U_{ij}$ is exactly the deviation of the empirical mean from its conditional expectation. Applying Theorem 4 of Maurer and Pontil (2009) to the variables $X_{k}:=A_{kj}$, $k\in\mathcal{N}_{i}^{(-j)}$, and taking a union bound over both tails yields the stated inequality. ∎
Lemma 3.8 provides a strictly finite-sample, variance-adaptive control of $U_{ij}$ directly in terms of the empirical variance $\hat{\sigma}_{ij}^{2}$, thereby completely bypassing the need for asymptotic plug-in justifications.
3.3 Berry-Esseen normal approximation
We now establish a normal approximation for the variance component $U_{ij}^{(-j)}$ defined in (3.1). Recall
$U_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$
and define the conditional variance
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}\mid P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
so that
$V_{ij}=\mathrm{Var}(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P)=\frac{1}{h_{n}^{2}}\sigma_{ij}^{2}.$
Under Assumption 2.2, there exists $\eta\in(0,1/2)$ such that
$\eta\leq P_{kj}\leq 1-\eta\qquad\text{for all }k,j.$
Hence
$\eta(1-\eta)\leq P_{kj}(1-P_{kj})\leq\frac{1}{4}.$
Since $|\mathcal{N}_{i}^{(-j)}|=h_{n}$, it follows that
$\eta(1-\eta)\,h_{n}\leq\sigma_{ij}^{2}\leq\frac{1}{4}\,h_{n}.$ (3.7)
In particular,
$\sigma_{ij}\geq\sqrt{\eta(1-\eta)}\,h_{n}^{1/2},\qquad\sigma_{ij}^{-1}\leq C\,h_{n}^{-1/2}$ (3.8)
for a constant $C>0$ independent of $(i,j,n)$. Consequently,
$V_{ij}=\frac{\sigma_{ij}^{2}}{h_{n}^{2}}\asymp h_{n}^{-1},$
and the conditional variance of $U_{ij}^{(-j)}$ does not degenerate.
Theorem 3.9 (Uniform Berry-Esseen bound).
Suppose Assumptions 2.1 and 2.2 hold and $h_{n}\asymp\sqrt{n\log n}$. Then there exists a universal constant $C>0$, independent of $i,j,n$, such that, uniformly over $1\leq i,j\leq n$,
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2}.$ (3.9)
Proof.
Fix $(i,j)$ and condition on $(\mathcal{N}_{i}^{(-j)},P)$. Define
$X_{k}:=A_{kj}-P_{kj},\qquad k\in\mathcal{N}_{i}^{(-j)}.$
By Proposition 2.7, the variables $\{X_{k}\}$ are independent, mean zero and bounded in [1,1][-1,1], but not necessarily identically distributed.
Let
$S:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k},\qquad\sigma_{ij}^{2}=\mathrm{Var}(S\mid\mathcal{N}_{i}^{(-j)},P).$
Define the third absolute moment sum
$\rho_{3}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathbb{E}\!\left[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P\right].$
Since $|X_{k}|\leq 1$, we have
$\rho_{3}\leq\sum_{k}\mathbb{E}[X_{k}^{2}\mid\mathcal{N}_{i}^{(-j)},P]=\sigma_{ij}^{2}.$
The Berry-Esseen inequality for independent, not necessarily identically distributed summands (see, e.g., Feller (1971, Chapter XVI)) states that there exists a universal constant $C_{0}>0$ such that
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\frac{\rho_{3}}{\sigma_{ij}^{3}}.$
Using $\rho_{3}\leq\sigma_{ij}^{2}$ yields
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\sigma_{ij}^{-1}.$
By (3.8), $\sigma_{ij}^{-1}\leq Ch_{n}^{-1/2}$ and therefore it does not depend on $i$ and $j$. Finally, since
$U_{ij}^{(-j)}=\frac{1}{h_{n}}S,\qquad V_{ij}=\frac{1}{h_{n}^{2}}\sigma_{ij}^{2},$
we have the identity
$\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}=\frac{S}{\sigma_{ij}},$
and the bound (3.9) follows. ∎
Corollary 3.10 (Asymptotic normality).
Under the conditions of Theorem 3.9,
$\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,1)\qquad\text{uniformly over }1\leq i,j\leq n.$
Proof.
Immediate from Theorem 3.9, since $h_{n}^{-1/2}\to 0$. ∎
3.4 Confidence intervals
Undder the canonical choice $h_{n}\asymp\sqrt{n\log n},$ we have
$h_{n}^{-1/2}=\left(\sqrt{n\log n}\right)^{-1/2}=(n\log n)^{-1/4}.$
Therefore the Berry-Esseen remainder in (3.9) satisfies
$\sup_{i,j}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|=O\!\big((n\log n)^{-1/4}\big).$
By Lemma 3.7, the bias component satisfies
$\sup_{i,j}|B_{ij}^{(-j)}|=O\!\big((\log n/n)^{1/4}\big).$
Hence the stochastic fluctuation scale and the bias scale are of the same order, up to logarithmic factors. In particular, to obtain asymptotically valid normal-approximation confidence intervals for $P_{ij}$, one must either undersmooth (so that the bias is $o(h_{n}^{-1/2})$) or incorporate an explicit bias correction of order $(\log n/n)^{1/4}$.
We construct two families of confidence intervals for $P_{ij}$: (i) finite-sample, variance-adaptive bands based on Lemma 3.8, and (ii) asymptotic normal-approximation bands based on Theorem 3.9 together with a bias cushion from Lemma 3.7.
3.4.1 Finite-sample Confidence Intervals via Empirical Bernstein Bounds
To construct finite-sample confidence intervals for the variance component $U_{ij}^{(-j)}$, we require a concentration inequality that depends only on observable quantities. Classical Bernstein inequalities involve the true variance $V_{ij}$, which is not directly accessible. In contrast, the Empirical Bernstein bound of Maurer and Pontil (2009) provides a deviation inequality that replaces the unknown variance with the sample variance, yielding a fully data-driven and nonasymptotic confidence interval.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Their empirical mean is the leave-one-out estimator
$\tilde{P}_{ij}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}.$ (3.10)
The unbiased sample variance of the observed edges in the neighborhood is
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}.$ (3.11)
Since the edges are binary, this expression simplifies to
$\hat{\sigma}_{ij}^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}).$ (3.12)
We apply Theorem 4 of Maurer and Pontil (2009) to the independent bounded variables $X_{k}:=A_{kj}$, $k\in\mathcal{N}_{i}^{(-j)}$, with sample size $h_{n}$. For any $\alpha\in(0,1)$, a union bound over the upper and lower tails yields the conditional deviation inequality
$\mathbb{P}\!\left(\big|U_{ij}^{(-j)}\big|\leq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}\;\middle|\;\mathcal{N}_{i}^{(-j)},P\right)\geq 1-\alpha.$ (3.13)
We therefore define the variance-adaptive half-width
$w_{ij}^{(B)}(\alpha):=\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}.$ (3.14)
The resulting finite-sample confidence interval for the localized average is
$CI_{ij}^{(B)}(\alpha):=\Big[\tilde{P}_{ij}-w_{ij}^{(B)}(\alpha),\;\tilde{P}_{ij}+w_{ij}^{(B)}(\alpha)\Big]\cap[0,1].$ (3.15)
By construction, conditionally on $(\mathcal{N}_{i}^{(-j)},P)$, this interval captures the neighborhood average
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$
with probability at least $1-\alpha$ for any finite $n$.
Remark 3.11 (Advantage of the Empirical Bound).
Unlike standard concentration inequalities that require substituting an estimated variance $\hat{V}_{ij}$ for the true variance $V_{ij}$, a step that typically degrades finite-sample guarantees and forces reliance on Slutsky’s Theorem and asymptotic normality, the Empirical Bernstein interval completely avoids the cost of studentization. The interval $CI_{ij}^{(B)}(\alpha)$ automatically adapts to the empirical variance of the neighborhood while strictly preserving its $1-\alpha$ coverage probability for any finite sample size.
3.4.2 Normal‑approximation intervals
Treat $U_{ij}$ as approximately normal and control the bias $B_{ij}$ via Lemma 3.7. For $\alpha\in(0,1)$ let $z_{1-\alpha/2}$ denote the $(1-\alpha/2)$-quantile of the standard normal and define
$w_{ij}^{(N)}(\alpha):=z_{1-\alpha/2}\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}\Big(\frac{\log n}{n}\Big)^{1/4},$ (3.16)
where $c_{\mathrm{bias}}>0$ is a bias‑cushion constant chosen large enough to dominate the bias in Lemma 3.7. The normal interval is
$\mathrm{CI}_{ij}^{(N)}(\alpha):=[\tilde{P}_{ij}-w_{ij}^{(N)}(\alpha),\ \tilde{P}_{ij}+w_{ij}^{(N)}(\alpha)]\cap[0,1].$ (3.17)
By Theorem 3.9 and the bias control, the coverage of $\mathrm{CI}_{ij}^{(N)}(\alpha)$ differs from $1-\alpha$ by at most $O\big((n\log n)^{-1/4}\big)$, up to the error from estimating $V_{ij}$.
Remark 3.12.
The Empirical Bernstein bands (3.15) are finite-sample and variance-adaptive but can be conservative. The normal bands (3.17) are typically tighter but rely on asymptotic approximation and an explicit bias cushion; the two constructions provide complementary guarantees in practice.
Algorithm 1 Simulation Procedure for LOO Smoothing and Confidence Intervals
1:Sample size $n$, neighborhood size $h$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
2:Estimated matrices $\widehat{P}_{\mathrm{LOO}},\widehat{P}_{\mathrm{class}}$ and empirical coverage rates
3:Draw $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$ and set $P_{ij}=h(\xi_{i},\xi_{j})$
4:Sample $A_{ij}\mid P_{ij}\sim\mathrm{Bernoulli}(P_{ij})$ for $i<j$ and symmetrize
5:For each $j$ form $A^{(-j)}$, compute distances $d_{ik}^{(-j)}$ via (2.4)
6:Build LOO neighborhoods $\mathcal{N}_{i}^{(-j)}$ of size $h_{n}$ and classical neighborhoods $\mathcal{N}_{i}$ from full AA
7:Compute $\tilde{P}_{ij}$ and $\widehat{P}_{ij}$ via (2.6) for both LOO and classical smoothers
8:Estimate $\widehat{V}_{ij}$ and construct intervals $\mathrm{CI}_{ij}^{(B)}(\alpha)$, $\mathrm{CI}_{ij}^{(N)}(\alpha)$
9:Evaluate empirical coverage $\widehat{\mathrm{Cov}}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}$ and report MSEs
4 Honest tuning via cross-validation
The leave-one-out (LOO) construction admits a simple and honest cross-validation (CV) scheme for selecting the neighborhood size $h$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held-out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h)$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h)$. Consequently, the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard
4.1 Cross-validation procedure and notation
Fix a row index $i\in[n]$. Let $\mathcal{G}$ be a finite grid of candidate neighborhood sizes,
$\mathcal{G}\subset\{1,2,\dots,n-2\},$
typically chosen around the theoretical scale $h_{n}\asymp\sqrt{n\log n}$. For each $h\in\mathcal{G}$ and each held-out column $j\neq i$, form the reduced adjacency matrix $A^{(-j)}$ by deleting the $j$-th row and column. This ensures that the neighborhood $\mathcal{N}_{i}^{(-j)}(h)$ is constructed without using any information from the held-out target column $j$. Compute the corresponding leave-one-out predictor
$\tilde{P}_{ij}^{(-j)}(h):=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h)}A_{kj},\qquad j\neq i.$
The empirical cross-validation score at row $i$ is the average squared prediction error across all available held-out columns:
$\widehat{R}_{i}(h):=\frac{1}{n-1}\sum_{j\neq i}\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h)\bigr)^{2}.$ (4.1)
The corresponding prediction risk is the conditional expectation of the same quantity. Because $A_{ij}$ is conditionally independent of $\mathcal{N}_{i}^{(-j)}(h)$ given $P$, this evaluates to
$R_{i}(h):=\frac{1}{n-1}\sum_{j\neq i}\mathbb{E}\!\left[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h)\bigr)^{2}\,\middle|\,\mathcal{N}_{i}^{(-j)}(h),P\right].$ (4.2)
We then select any minimizer of the empirical score over the grid,
$\hat{h}\in\arg\min_{h\in\mathcal{G}}\widehat{R}_{i}(h).$
5 Simulation Study
We conducted a simulation study to evaluate the finite-sample performance of the proposed leave-one-out (LOO) neighborhood smoother and compare it with the classical (ZLZ) neighborhood smoother. All computations were implemented in Python using NumPy with a fixed random seed.
5.1 Setup
We considered five representative graphon models capturing a range of smoothness and structural complexity:
Table 1: Graphon Model Definitions
Model Definition
Smooth $0.5+0.3\sin(\pi u)\sin(\pi v)$
Block 2-block SBM ($p_{\text{in}}=0.7,p_{\text{out}}=0.3$)
Wiggly $0.5+0.25\sin(4\pi uv)$
Rank-one $(u+v)/2$
Spiky $0.2+0.8\,\mathbbm{1}\{|u-0.5|<0.1,\ |v-0.5|<0.1\}$
For each graphon, we generated an undirected graph with $n=200$ nodes by sampling latent positions $\xi_{i}\sim\mathrm{Unif}[0,1]$ and edges $A_{ij}\sim\mathrm{Bernoulli}(h(\xi_{i},\xi_{j}))$ independently for $i<j$. The neighborhood size and trimming quantile were fixed at
$h=\lfloor 1.5\sqrt{n\log n}\rfloor.$
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column and row $j$ removed, thereby restoring conditional independence between the neighborhood and the target edge. Both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$.
5.2 Evaluation metrics
We report two performance measures.
Row-wise MSE.
For node $i=0$,
$\mathrm{MSE}_{0}=\frac{1}{n}\sum_{j=1}^{n}(\widehat{P}_{0j}-P_{0j})^{2}.$
Coverage of confidence intervals.
For the LOO estimator, we consider:
  • Empirical Bernstein intervals:
    $w_{ij}^{(B)}=\sqrt{2\widehat{V}_{ij}\log(2/\alpha)}+\frac{\log(2/\alpha)}{3h},$
  • Normal intervals:
    $w_{ij}^{(N)}=1.96\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}(\log n/n)^{1/4},\qquad c_{\mathrm{bias}}=0.1,$
where the plug-in variance proxy is
$\widehat{V}_{ij}=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}).$
Intervals are truncated to $[0,1]$. Coverage is computed as
$\mathrm{Coverage}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}.$
5.3 Results
Table 2 summarizes the results for all five graphons. The LOO smoother achieves a comparable MSE to the classical ZLZ smoother. Coverage for both confidence interval constructions is close to the nominal $95\%$ level, with the Empirical Bernstein intervals slightly more conservative.
Graphon $\mathrm{MSE}_{0}^{\mathrm{LOO}}$ $\mathrm{MSE}_{0}^{\mathrm{ZLZ}}$ $\mathrm{Cov}_{\mathrm{Bousquet}}$ $\mathrm{Cov}_{\mathrm{Normal}}$
Smooth 0.0027 0.0030 0.999 0.995
Block 0.0021 0.0027 1.000 1.000
Wiggle 0.0031 0.0030 0.944 0.885
Rank-one 0.0029 0.0030 0.999 0.996
Spiky 0.0021 0.0033 0.977 0.977
Table 2: Row-wise MSE and empirical coverage for five graphons at $n=200$.
6 Discussion and Extensions
This paper develops a principled framework for inference on individual edge probabilities in graphon models based on a simple leave-one-out modification of neighborhood smoothing. By constructing LOO neighborhoods that do not depend on the column being predicted, we restore conditional independence of the centered edge variables and thereby make classical concentration and normal-approximation tools available. The resulting methodology delivers entrywise confidence intervals and an honest cross-validation scheme for data-driven tuning, while retaining the optimal row-wise mean-squared error rate known for classical neighborhood smoothing.
The LOO construction is conceptually simple and quite general, which makes it a promising building block for extensions beyond the basic undirected, unweighted, static setting considered here. we outline several directions for future work.
  1. 1.
    Directed and weighted graphons. For directed networks, one could replace the Zhang et al. (2017) distance by an asymmetric distance that incorporates both in and out neighborhood information, and then apply the same LOO decoupling on the relevant columns. For weighted networks, a similar strategy could be pursued using appropriate moment-based distances built from $A$ and $A^{2}$, with boundedness or tail assumptions tailored to the edge-weight distribution.
  2. 2.
    Degree correction and heterogeneous sparsity. Many real networks exhibit highly heterogeneous degree distributions. Extending the present analysis to degree corrected graphon or blockmodel formulations would require incorporating node-specific scale factors into both the distance and the bias analysis. The LOO idea remains applicable, but the variance proxy $V_{ij}$ and the bias bounds would need to be adapted to handle varying effective sample sizes across nodes.
  3. 3.
    Adaptive and simultaneous confidence bands. While we have focused on pointwise confidence intervals for individual edge probabilities, the same machinery can be used to construct simultaneous bands over subsets of rows or columns, for example by combining the Berry-Esseen bounds with high-dimensional Gaussian approximation results. It would also be natural to let the bias cushion constant $c_{\mathrm{bias}}$ be selected in a data-driven way, for instance by cross-validation or stability criteria, in order to balance coverage and tightness of the intervals.
  4. 4.
    Applications to real networks. Finally, it would be valuable to apply the proposed methodology to real data sets, such as social, citation, or biological networks, in order to assess its practical performance and interpretability. In particular, the ability to attach uncertainty quantification to individual edge probabilities or to row-wise summaries could provide new insights in applied settings where only a single large network is observed.
The main limitation of the current approach is the requirement of piecewise Lipschitz smoothness for the underlying graphon. Extensions to rougher kernels, highly discontinuous structures, or models with sharp community boundaries would likely require modified distance metrics and a more delicate bias analysis. Addressing these challenges, while retaining tractable inference procedures, remains an important open problem.