### v4

# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

Abstract
Estimating latent edge probabilities in network models often relies on aggregating information from structurally similar nodes. However, using the same adjacency matrix to both determine node similarities and estimate probabilities induces complex statistical dependencies. This coupling invalidates standard inferential tools, such as concentration inequalities and central limit theorems, leaving a critical gap in entrywise uncertainty quantification for networks. To resolve this, we introduce a leave-one-out (LOO) decoupling strategy. When predicting a target edge, the local neighborhood is constructed from a reduced adjacency matrix with the target row and column entirely removed. This structural modification-rather than the specific smoothing mechanism-is the core innovation: it restores exact conditional independence between the data-driven neighborhood selection and the observed edges being aggregated. Leveraging this restored independence, we derive variance-adaptive empirical Bernstein bounds and establish uniform Berry-Esseen normal approximations with explicit rates. These results enable finite-sample, variance-adaptive confidence intervals for the localized neighborhood mean and entrywise uncertainty quantification for $P_{ij}$ on a high-probability interior (good) set up to an explicit bias term. We also establish a uniform Berry–Esseen approximation for the variance component (normal‑approximation intervals use a heuristic variance plug‑in in this version).
Furthermore, the LOO architecture naturally admits an honest cross-validation scheme for tuning parameter selection, providing a complete and rigorous inferential framework while preserving strong mean-squared error performance.
1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2008; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $f$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, modern estimators-including universal singular value decomposition (Chatterjee, 2015), sorting-based smoothing (Chan and Airoldi, 2014), and block-approximations (Gao et al., 2015)-have been proven to reliably achieve minimax-optimal rates for point estimation.
Among these, the neighborhood smoothing method introduced by Zhang et al. (2017) stands out. By adaptively aggregating structurally similar nodes based on empirical two-hop distances, neighborhood smoothing achieves optimal mean-squared error (MSE) without requiring the strict structural assumptions of block models or the selection of a global low-rank embedding.
However, a fundamental gap remains: the transition from optimal point estimation to rigorous statistical inference. Given an entrywise estimate $\widehat{P}_{ij}$, it is critically important for downstream scientific tasks to attach a valid measure of uncertainty to this prediction. Yet, existing inferential theory for networks has focused almost exclusively on low-dimensional or global functionals. Significant progress has been made in proving limit theorems for spectral embeddings (Sussman et al., 2012; Athreya et al., 2016), asymptotic normality for community parameters in SBMs (Bickel et al., 2013), and global goodness-of-fit testing (Lei, 2016). Conversely, entrywise uncertainty quantification-specifically, constructing valid confidence intervals and proving central limit theorems for individual edge probabilities under continuous graphons-has remained theoretically elusive. Recent attempts to bridge this gap have relied on computationally intensive resampling methods, such as network bootstrapping (Levin and Levina, 2025; Green and Shalizi, 2022), or generalized fiducial inference (Su et al., 2022), but exact distribution-free analytical bounds for neighborhood estimators are lacking.
The obstacle preventing entrywise inference is not merely technical, but structural. To achieve optimal MSE, neighborhood smoothing estimators (Zhang et al., 2017) rely on a data-driven aggregation of similar nodes. This creates a severe “double-dipping” problem: the procedure reuses the adjacency matrix twice. The matrix $A$ is first used to compute pairwise topological distances to select a neighborhood $\mathcal{N}_{i}$, and then the very same matrix $A$ is used to average the edges within that neighborhood. Consequently, the random neighborhood weights depend directly on the edges being averaged. Even after conditioning on the latent probability matrix $P$, the summands in the estimator are strongly coupled. This intractable dependence invalidates the foundational assumptions of classical concentration inequalities and Berry-Esseen normal approximations.
In this paper, we demonstrate that this probabilistic bottleneck can be cleanly bypassed via a simple but highly principled structural modification: leave-one-out (LOO) neighborhood smoothing.
When estimating a single entry $P_{ij}$, we perform a topological surgery on the graph by constructing the neighborhood of node $i$ using a reduced adjacency matrix $A^{(-j)}$ in which the $j$-th row and column have been entirely deleted. All pairwise distances and nearest-neighbor selections are computed on this reduced matrix. Only after the neighborhood $\mathcal{N}_{i}^{(-j)}$ is fixed do we average the original target edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$.
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling permits the application of classical empirical process tools to graphon neighborhood smoothing. Specifically, under standard piecewise Lipschitz assumptions on the graphon, we leverage this restored independence to establish three main contributions:
  1. 1.
    Variance-Adaptive Concentration: We derive tight, finite-sample concentration inequalities based on empirical Bernstein bounds that automatically adapt to the local variance of the edge probabilities.
  2. 2.
    Asymptotic Normality: We establish a uniform Berry–Esseen bound with explicit rates for the normalized variance component; normal‑approximation intervals in this version use a heuristic variance estimate and incur an additional variance‑estimation error term.
  3. 3.
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$.inequality. Here, honest cross-validation means that it does not reuse held-out edges in predictor construction, ensuring valid inference even after data-driven tuning.
Crucially, these inferential guarantees preserve the optimal rowwise mean-squared error rate of classical neighborhood smoothing.
The remainder of the paper is organized as follows. Section 2 formalizes the graphon model, the Zhang-Levina-Zhu (ZLZ), Zhang et al. (2017) empirical distance, and our leave-one-out neighborhood construction. Section 3.2 establishes the foundational conditional independence, proving the variance-adaptive concentration bounds, Berry-Esseen normal approximations, and the validity of the resulting confidence intervals. Section 4 details the honest cross-validation scheme for tuning parameter selection. Section 5 reports comprehensive simulation results illustrating the finite-sample coverage and MSE performance, and Section 6 concludes with a discussion of extensions to directed, weighted, and degree-corrected settings.
2 Model and Leave-One-Out Neighborhood Smoothing
We observe a single undirected simple graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$, no self‑loops ($A_{ii}=0$), and independent upper‑triangular entries
$A_{ij}\mid P_{ij}\ \sim\ \mathrm{Bernoulli}(P_{ij}),\qquad 1\leq i<j\leq n,$ (2.1)
with $A_{ij}=A_{ji}$. The matrix $P=(P_{ij})$ contains unknown edge probabilities. We adopt the standard graphon parametrization below.
Assumption 2.1 (Graphon model and smoothness).
The edge probabilities satisfy $P_{ij}=f(\xi_{i},\xi_{j})$ with $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$. The graphon $f:[0,1]^{2}\to[0,1]$ is piecewise Lipschitz in each argument with Lipschitz constant $L$ on a fixed finite rectangular partition of $0,1]^{2}$.
Assumption 2.2 (Non‑degeneracy).
There exists $\eta\in(0,1/2)$ such that
$\eta\leq f(u,v)\leq 1-\eta\qquad\text{for all }(u,v)\in[0,1]^{2}.$
Assumption 2.2 places the model in the dense graphon regime, where edge probabilities are uniformly bounded away from zero and one. This excludes sparse scalings of the form $P_{ij}=\rho_{n}W(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$. Our Berry-Esseen and empirical Bernstein arguments rely critically on variance bounded away from zero. Extending the present theory to sparse or degree-corrected regimes would require variance renormalization and additional control of small-probability events, and is left for future work.
Remark 2.3.
Assumption 2.2 requires that there exists $\eta\in(0,1/2)$ such that $\eta\leq f(u,v)\leq 1-\eta$ for all $(u,v)\in[0,1]^{2}$. This nondegeneracy condition guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance‑adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re‑normalization and additional control of rare events, which we leave to future work. Empirically, moderate thinning of edges does not invalidate the procedure (see Section 5).
Remark 2.4.
The rectangular partition in Assumption 2.1 is fixed and does not grow with $n$. All constants in our rates may depend on the Lipschitz constant $L$ and on the partition, but not on $n$.
Remark 2.5 (Dense regime convention).
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=f(\xi_{i},\xi_{j})$ with $f$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two‑hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ are discussed briefly in Section 6.
2.1 LOO neighborhoods and estimator
Consider the set of nodes $V=\{1,\dots,n\}$. Fix $j\in[n]$ and let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting row and column $j$ from $A$. Define the two-hop matrices
$M:=\frac{1}{n}A^{2},\qquad M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}.$
For distinct $i,k\in V\setminus\{j\}$, define the leave-one-out (LOO)-distance
[…]
Let
$d^{(-j)}_{i,(1)}\leq\cdots\leq d^{(-j)}_{i,(n-2)}$
denote the order statistics of the set
$\bigl\{d^{(-j)}(i,k):k\in V\setminus\{i,j\}\bigr\}.$
Define the leave-one-out neighborhood of node $i$ by
$\mathcal{N}_{i}^{(-j)}=\left\{k\neq i,j:d_{ik}^{(-j)}\leq d_{i,(h_{n})}^{(-j)}\right\},$ (2.3)
where $h_{n-1}$ is the ZLZ bandwidth parameter with $n$ replaced by n1n-1. that is, the $h_{n}$ nearest neighbors of $i$ according to the ZLZ two-hop distance computed on the reduced graph.
Given $\mathcal{N}_{i}^{(-j)}$, define the one‑sided LOO smoother and its symmetrized version by
$\tilde{P}_{ij}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},\qquad\widehat{P}_{ij}:=\frac{\tilde{P}_{ij}+\tilde{P}_{ji}}{2}.$ (2.4)
This construction preserves the original ZLZ neighborhood rule, with the sole modification that node $j$ is removed prior to distance computation.
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Define the distance function $d:V\times V\to\mathbb{R}_{+}$ by
$d_{ZLZ}(i,j):=\begin{cases}0,&\text{for }i=j,\\ \max_{k\neq i,j}\Big|\big(A^{2}/n\big)_{ik}-\big(A^{2}/n\big)_{jk}\Big|,&\text{for }i\neq j.\end{cases}$ (2.5)
This is the sup‑norm discrepancy between the $i$th and $j$th rows of the normalized two‑hop matrix $A^{2}/n$, measuring the similarity of their two‑hop neighborhoods. Therefore the ZLZ neighborhood $\mathcal{N}_{i}$ consists of the $h_{n}$ nodes with the smallest ZLZ distance to $i$. Zhang et al. (2017) defined
$\mathcal{N}_{i}=\{k\in V\setminus\{i\}:d_{ZLZ}(i,k)\leq q_{h}(\{D\}),\quad D=\{d_{ZLZ}(i,k):k\in V\setminus\{i\}\}\},$ (2.6)
and $q_{h}(\{D\})$ is the $h_{n}$-th order statistic of the set $\{d_{ZLZ}(i,k):k\in V\setminus\{i\}\}$. The ZLZ estimator is then defined by
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}A_{kj}.$
Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry-Esseen theory.
Remark 2.6.
The symmetrized estimator $\widehat{P}_{ij}$ respects the symmetry of the simple graph model. All concentration and normal‑approximation results are derived for the one‑sided quantities $\tilde{P}_{ij}$; symmetrization does not affect the independence structure.
Proposition 2.7 (Conditional independence).
Conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$, the variables
$\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$
are independent, mean-zero, and bounded in [1,1][-1,1].
The leave-one-out construction does not alter the smoothing mechanism itself; it alters only the dependency structure. By removing the target column prior to neighborhood selection, the summands in the neighborhood average become exactly independent conditional on $P$. This structural decoupling is the key step that allows classical empirical process tools to be applied without resampling or sample splitting.
Proof.
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=f(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$.
By construction, $A^{(-j)}$ is obtained from $A$ by removing the $j$th row and column. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$.
For each $k\in\mathcal{N}_{i}^{(-j)}$, we have $A_{kj}\sim\mathrm{Bernoulli}(P_{kj})$ and $A_{kj}-P_{kj}\in[-1,1]$ with mean zero. Since the edges $\{A_{kj}:k\neq j\}$ are independent conditional on $P$, and $\mathcal{N}_{i}^{(-j)}$ is independent of these edges, it follows that the collection $\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ is independent, mean-zero, and bounded in [1,1][-1,1] conditional on $(\mathcal{N}_{i}^{(-j)},P)$. ∎
3 Consistency
Our objective is entrywise inference for the finite-sample edge probability matrix $P=(P_{ij})_{1\leq i,j\leq n}$, where $P_{ij}=W(\xi_{i},\xi_{j})$. All probability statements in Sections 3 and 4 are conditional on the realized latent positions $(\xi_{1},\dots,\xi_{n})$, equivalently on $P$. Thus our results provide conditional finite-sample and asymptotic inference for $P_{ij}$ rather than for the graphon $W$ itself. In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write For $i\neq j$, write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}^{(-j)}+B_{ij}^{(-j)},$ (3.1)
where
$U_{ij}^{(-j)}$ $:=\tilde{P}_{ij}-\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.2)
$B_{ij}^{(-j)}$ $:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}-P_{ij}.$ (3.3)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}^{(-j)}$ are independent, mean zero and bounded, so $U_{ij}^{(-j)}$ behaves like an average of $h_{n}$ independent observations. The term $B_{ij}^{(-j)}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2. The following two subsections analyze these two components separately.
3.1 Neighborhood bias
The bias $B_{ij}$ captures the approximation error from replacing $P_{ij}$ by the local average of $P_{kj}$ over the LOO neighborhood. Under the piecewise Lipschitz graphon assumption and the ZLZ neighborhood construction, neighbors in $\mathcal{N}_{i}^{(-j)}$ are close in latent position to $\xi_{i}$, which yields a small bias.
Lemma 3.1 (Uniform two-hop stability under deletion of one vertex).
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of an undirected simple graph with $A_{ii}=0$. For each $j\in[n]$, let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting the $j$th row and column. Define the normalized two-hop matrices
$M:=\frac{1}{n}A^{2},\qquad M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}.$
To compare $M$ and $M^{(-j)}$, embed $M^{(-j)}$ into an $n\times n$ matrix $M^{(-j)}_{\mathrm{embed}}$ by
$M^{(-j)}_{\mathrm{embed}}(i,\ell)=\begin{cases}M^{(-j)}_{i\ell},&i,\ell\neq j,\\ M_{i\ell},&\text{if }i=j\text{ or }\ell=j.\end{cases}$
Thus $M^{(-j)}_{\mathrm{embed}}$ and $M$ have the same dimensions and agree on the $j$th row and column.
Then, deterministically,
$\sup_{1\leq j\leq n}\ \|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\;\leq\;\frac{2}{n},$
where $\|B\|_{\max}:=\max_{1\leq i,\ell\leq n}|B_{i\ell}|$. In particular, for all $j\in[n]$ and all $i,\ell\in[n]\setminus\{j\}$,
$|M_{i\ell}-M^{(-j)}_{i\ell}|\leq\frac{2}{n}.$
Proof.
Fix $j\in[n]$ and $i,\ell\in[n]\setminus\{j\}$. Expanding the $(i,\ell)$ entry gives
$(A^{2})_{i\ell}=\sum_{m=1}^{n}A_{im}A_{m\ell},\qquad\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell},$
hence
$(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}=A_{ij}A_{j\ell},$
so $\big|(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}\big|\leq 1$. Therefore,
[…]
$\leq\left|\frac{1}{n}(A^{2})_{i\ell}-\frac{1}{n}\big(A^{(-j)}\big)^{2}_{i\ell}\right|+\left|\frac{1}{n}-\frac{1}{n-1}\right|\big(A^{(-j)}\big)^{2}_{i\ell}$
$\leq\frac{1}{n}+\frac{1}{n(n-1)}\big(A^{(-j)}\big)^{2}_{i\ell}.$
Since $\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell}\leq n-1$, we get
$\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{1}{n}+\frac{n-1}{n(n-1)}=\frac{2}{n}.$
For $i=j$ or $\ell=j$, we have $M^{(-j)}_{\mathrm{embed}}(i,\ell)=M_{i\ell}$ by definition, so the difference is zero. Hence
$\|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\leq\frac{2}{n}$
for each $j$, and taking the supremum over $j$ preserves the bound. ∎
Lemma 3.2 (Stability of expected two-hop matrices under deletion).
Let $A$ and $A^{(-j)}$ be as in Lemma 3.1, and define
$M=\frac{1}{n}A^{2},\qquad M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)}=\frac{1}{n-1}(A^{(-j)})^{2},\qquad M^{(-j)\star}=\mathbb{E}[M^{(-j)}\mid P].$
Embed $M^{(-j)}$ and $M^{(-j)\star}$ into $n\times n$ matrices as in Lemma 3.1. Then, for all $n$,
$\sup_{1\leq j\leq n}\big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}\;\leq\;\|M-M^{\star}\|_{\max}+\frac{C}{n},$ (3.4)
for a universal constant $C>0$.
Proof.
Fix $j\in[n]$. Let $M^{(-j)}_{\mathrm{embed}}$ and $M^{(-j)\star}_{\mathrm{embed}}$ be the $n\times n$ embeddings of $M^{(-j)}$ and $M^{(-j)\star}$ defined as in Lemma 3.1. Then
$\|M^{(-j)}_{\mathrm{embed}}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\|M^{(-j)}_{\mathrm{embed}}-M\|_{\max}+\|M-M^{\star}\|_{\max}+\|M^{\star}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}.$
By Lemma 3.1,
$\sup_{1\leq j\leq n}\|M^{(-j)}_{\mathrm{embed}}-M\|_{\max}\leq\frac{2}{n}.$
Next, since
$M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)\star}_{\mathrm{embed}}=\mathbb{E}[M^{(-j)}_{\mathrm{embed}}\mid P],$
we have
$M^{\star}-M^{(-j)\star}_{\mathrm{embed}}=\mathbb{E}[M-M^{(-j)}_{\mathrm{embed}}\mid P],$
and thus, by Jensen’s inequality and Lemma 3.1,
$\|M^{\star}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\mathbb{E}\big[\|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\mid P\big]\leq\frac{2}{n}.$
Combining the bounds,
$\|M^{(-j)}_{\mathrm{embed}}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\|M-M^{\star}\|_{\max}+\frac{4}{n}.$
Taking the supremum over $j$ yields (3.4). ∎
Proposition 3.3 (Theorem 1 of Zhang et al. (2017): Row-wise risk).
Work under the assumptions of Zhang et al. (2017) and choose the ZLZ quantile cutoff parameter $h_{\mathrm{ZLZ}}=C_{0}(\log n/n)^{1/2}$, so that the resulting neighborhood sizes satisfy $|N_{i}|\geq(n\log n)^{1/2}$ (Lemma 2(1) in their supplementary material). Let $N_{i}$ be the ZLZ neighborhood of node $i$ constructed from the two-hop matrix $M=A^{2}/n$ via their distance-and-quantile rule, and define
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|N_{i}|}\sum_{k\in N_{i}}A_{kj}.$
Then there exist constants $C,c>0$ such that with probability at least $1-Cn^{-c}$, uniformly over all $i\in[n]$,
$\frac{1}{n}\sum_{j=1}^{n}\big(\widetilde{P}^{\mathrm{ZLZ}}_{ij}-P_{ij}\big)^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$
Proof.
This is exactly the row-wise high-probability bound proved in the supplementary material of Zhang et al. (2017): after the decomposition (1), the row-averages of the two terms are bounded in (15) and (16) of their supplementary material, uniformly for all $i$ on an event of probability $1-O(n^{-c})$. ∎
Corollary 3.4 (Leave-one-out ZLZ row-wise risk and $(U_{ij},B_{ij})$ decomposition).
Fix $j\in[n]$. Let $A^{(-j)}$ be the adjacency matrix obtained by deleting row and column $j$ from $A$, and let $\widetilde{P}$ denote the ZLZ neighborhood-smoothing estimator computed from $A^{(-j)}$, with neighborhoods $N_{i}^{(-j)}$ defined exactly as in ZLZ but using the two-hop matrix $(A^{(-j)})^{2}/(n-1)$. Recall (3.1). For each $i\neq j$ and $\ell\neq j$, define
$U^{(-j)}_{i\ell}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(A_{k\ell}-P_{k\ell}),\qquad B^{(-j)}_{i\ell}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(P_{k\ell}-P_{i\ell}),$
so that
$\widetilde{P}_{i\ell}-P_{i\ell}=U^{(-j)}_{i\ell}+B^{(-j)}_{i\ell}$.
Then for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\leq\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;+\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2}.$
Moreover, applying Theorem 1 of Zhang et al. (2017) to the (n1)(n-1)-node graph $A^{(-j)}$ yields, with probability at least $1-c_{1}n^{-c_{2}}$, uniformly for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Identifying ZLZ’s notation via
$J_{1}^{(-j)}(i,\ell):=\big(U^{(-j)}_{i\ell}\big)^{2},\qquad J_{2}^{(-j)}(i,\ell):=\big(B^{(-j)}_{i\ell}\big)^{2},$
the ZLZ bounds (with (n1)(n-1) in place of $n$) give
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2},\qquad\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Proof.
The decomposition
$\widetilde{P}_{i\ell}-P_{i\ell}=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\bigl[(A_{k\ell}-P_{k\ell})+(P_{k\ell}-P_{i\ell})\bigr]$
is identical to ZLZ equation (1), and the inequality
$(\widetilde{P}_{i\ell}-P_{i\ell})^{2}\leq 2J_{1}^{(-j)}(i,\ell)+2J_{2}^{(-j)}(i,\ell)$
follows exactly as in their argument, with the identifications $J_{1}^{(-j)}(i,\ell)=(U^{(-j)}_{i\ell})^{2}$ and $J_{2}^{(-j)}(i,\ell)=(B^{(-j)}_{i\ell})^{2}$. Next, ZLZ’s control of two-hop quantities
$\frac{1}{n}(A^{2})_{ab}-\frac{1}{n}(P^{2})_{ab}$
is established in equations (4)-(6) of their supplementary material. Lemma 3.1 shows that the same bounds hold when $A$ is replaced by $A^{(-j)}$ (and $n$ by n1n-1). Therefore, all subsequent ZLZ steps that rely on these concentration inputs remain valid for the leave-one-out graph.
ZLZ bound the row-average of $J_{1}$ using their decomposition (11) and bounds (12), (14), and (15). The same argument applies verbatim to $A^{(-j)}$ and $N_{i}^{(-j)}$, yielding
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Similarly, ZLZ bound the row-average of $J_{2}$ using Cauchy-Schwarz and their “claim 2 of Lemma 2,” concluding in equation (16). Again, Lemma 3.1 ensures the same concentration inputs hold for $A^{(-j)}$, giving
$\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Combining these bounds yields the stated row-wise risk bound for the leave-one-out estimator. ∎
Assumption 3.5 (Piecewise bi-Lipschitz smoothness).
The graphon $W:[0,1]^{2}\to[0,1]$ admits a finite vertical partition $0=b_{0}<b_{1}<\cdots<b_{K}=1$ with blocks $I_{m}:=(b_{m},b_{m+1})$ such that:
  1. 1.
    (Within-block Lipschitz in the first argument.) For all $m$ and all $u,u^{\prime}\in I_{m}$, $x\in[0,1]$,
    $|W(u,x)-W(u^{\prime},x)|\leq L|u-u^{\prime}|.$
  2. 2.
    (Across-block $L^{2}$ separation.) There exist $\eta,\delta>0$ such that for any $u\in I_{m}$ and $u^{\prime}\in I_{m^{\prime}}$ with $m\neq m^{\prime}$,
    $\lambda\big(\{x\in[0,1]:|W(u,x)-W(u^{\prime},x)|\geq\eta\}\big)\geq\delta.$
  3. 3.
    (Within-block $L^{2}$ identifiability.) For any $\varepsilon>0$ with $b_{m}+\varepsilon<b_{m+1}-\varepsilon$, there exists $\kappa(\varepsilon)>0$ such that for all $u,u^{\prime}\in[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$,
    $\int_{0}^{1}(W(u,x)-W(u^{\prime},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|u-u^{\prime}|^{2}.$
Remark 3.6 (Scope of Assumption 3.5.).
This assumption excludes graphons with flat regions or non-identifiable row structure, and is tailored to settings where rowwise $L^{2}$ proximity reflects latent proximity within each block. In particular, it does not cover graphons with a growing number of blocks or highly irregular multiscale structure. The condition is standard in nonparametric graphon estimation when entrywise inference is desired, but we emphasize that it restricts the theory to graphons with fixed, well-separated smooth components.
Remark 3.7 (Good-node and good-edge sets.).
Fix $\varepsilon>0$ satisfying the feasibility condition
$b_{m}+\varepsilon<b_{m+1}-\varepsilon\qquad\text{for all }m=0,\dots,K-1,$
so that each interior interval $b_{m}+\varepsilon,b_{m+1}-\varepsilon]$ is nonempty. Define the block interior
$I_{\varepsilon}:=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,\;b_{m+1}-\varepsilon].$
Let the random good-node set be
$\mathcal{G}_{\varepsilon}:=\{i\in[n]:\ \xi_{i}\in I_{\varepsilon}\},$
and the good-edge set
$\mathcal{E}_{\varepsilon}:=\{(i,j):i\neq j,\ i\in\mathcal{G}_{\varepsilon},\ j\in\mathcal{G}_{\varepsilon}\}.$
Since $\xi_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$, we have $|\mathcal{G}_{\varepsilon}|\sim\mathrm{Bin}\big(n,\lambda(I_{\varepsilon})\big)$. In particular, for any $t>0$ Hoeffding’s inequality gives
$\mathbb{P}\!\left(\left|\frac{|\mathcal{G}_{\varepsilon}|}{n}-\lambda(I_{\varepsilon})\right|\geq t\right)\leq 2\exp(-2nt^{2}).$
Consequently, for fixed $\varepsilon>0$ the fraction of good nodes satisfies $|\mathcal{G}_{\varepsilon}|/n=\lambda(I_{\varepsilon})+O_{\mathbb{P}}(n^{-1/2})$, and if $\lambda(I_{\varepsilon})=1-O(\varepsilon)$ then $|\mathcal{G}_{\varepsilon}|=n(1-O(\varepsilon))$ with exponentially high probability.
Lemma 3.8 (Rowwise $L^{2}$ closeness implies latent closeness).
Assume Assumption 3.5. Fix $\varepsilon>0$ and define
$\mathcal{I}_{\varepsilon}=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,b_{m+1}-\varepsilon].$
Let $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$ and set $P_{ab}=W(U_{a},U_{b})$. Let $r_{n}=(\log n/n)^{1/4}$. There exist constants $c_{0},c_{1}>0$ depending only on Assumption 3.5 and $\varepsilon$ such that the following holds. If
$M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq c_{0},\qquad\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2},$
then with probability at least $1-n^{-c}$ (for some $c>0$) and for all sufficiently large $n$,
$U_{i},U_{k}\text{ lie in the same block},\qquad|U_{i}-U_{k}|\leq c_{1}\sqrt{M_{n}}\,r_{n}.$
Proof.
Let $f_{i}(x)=W(U_{i},x)$ and $f_{k}(x)=W(U_{k},x)$, and define
$D(U_{i},U_{k})=\int_{0}^{1}(f_{i}(x)-f_{k}(x))^{2}\,dx.$
Concentration of the empirical row distance. Conditional on $(U_{i},U_{k})$, the variables
$Z_{\ell}:=(P_{i\ell}-P_{k\ell})^{2}=(W(U_{i},U_{\ell})-W(U_{k},U_{\ell}))^{2}$
are i.i.d. in $\ell$, bounded in $[0,1]$, and satisfy $\mathbb{E}[Z_{\ell}\mid U_{i},U_{k}]=D(U_{i},U_{k})$. Hoeffding’s inequality gives, for any $t>0$,
$\mathbb{P}\!\left(\left|\frac{1}{n}\sum_{\ell=1}^{n}Z_{\ell}-D(U_{i},U_{k})\right|\geq t\ \middle|\ U_{i},U_{k}\right)\leq 2e^{-2nt^{2}}.$
Taking $t=C(\log n/n)^{1/2}$ yields that, with probability at least $1-n^{-c}$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\geq D(U_{i},U_{k})-C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.5)
If $U_{i},U_{k}$ lie in different blocks, Assumption 3.5(ii) gives $D(U_{i},U_{k})\geq\eta^{2}\delta$. Using (3.5), for large $n$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\geq\frac{\eta^{2}\delta}{2}.$
Choosing $c_{0}=\eta^{2}\delta/4$ makes this incompatible with the assumed upper bound. Thus $U_{i},U_{k}$ must lie in the same block.
On the event of $U_{i},U_{k}$ lie in the same block and belong to $\mathcal{I}_{\varepsilon}$, so Assumption 3.5(iii) implies
$D(U_{i},U_{k})\geq\kappa(\varepsilon)|U_{i}-U_{k}|^{2}.$
Combining this with (3.5) and the assumed upper bound gives
$\kappa(\varepsilon)|U_{i}-U_{k}|^{2}\leq\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}+C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq(M_{n}+C)\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Therefore,
$|U_{i}-U_{k}|\leq\sqrt{\frac{M_{n}+C}{\kappa(\varepsilon)}}\,\Bigl(\frac{\log n}{n}\Bigr)^{1/4}\leq c_{1}\sqrt{M_{n}}\,r_{n},$
after adjusting constants. ∎
Corollary 3.9 (Bias bound on the good set).
Fix $\varepsilon>0$ satisfying the feasibility condition above and suppose the neighborhood regularity condition (display (3.4) in the main text) holds with constant $C_{2}$. Then there exists a constant $C(\varepsilon,C_{2})>0$ (depending on $L,\eta,\delta,\kappa(\varepsilon)$ and $C_{2}$) such that, with probability at least $1-c_{1}n^{-c_{2}}$,
$\sup_{\begin{subarray}{c}i\neq j:\\ (i,j)\in\mathcal{E}_{\varepsilon}\end{subarray}}|B_{ij}^{(-j)}|=O_{\mathbb{P}}\big(C(\varepsilon,C_{2})\,r_{n}\big),\qquad r_{n}=(\log n/n)^{1/4}.$
Assumption 3.10 (Neighborhood regularity).
For each $(i,j)\in\mathcal{E}_{\varepsilon}$, the leave-one-out neighborhood $\mathcal{N}_{i}^{(-j)}$ satisfies
$\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}$
with probability at least $1-n^{-c}$.
Remark.
Assumption 3.10 is consistent with the rowwise risk bounds of Zhang-Levina-Zhu together with the two-hop stability lemmas established above. We do not provide a fully explicit derivation of this condition for LOO neighborhoods here, but it is compatible with existing theory.
Lemma 3.11 (Bias bound for the LOO neighborhood smoother).
Assume Assumption 3.5 and let $r_{n}=(\log n/n)^{1/4}$. Suppose that, with probability at least $1-n^{-c}$, for all $i\neq j$,
$\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq C_{2}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.6)
Then for all $i\neq j$ with $U_{i}\in\mathcal{I}_{\varepsilon}$,
$|B^{(-j)}_{ij}|\leq Cr_{n}.$
The canonical choice $h_{n}\asymp\sqrt{n\log n}$ balances the stochastic fluctuation scale $h_{n}^{-1/2}$ and the bias scale $r_{n}=(\log n/n)^{1/4}$. Undersmoothing ($h_{n}=o(\sqrt{n\log n})$) reduces bias at the cost of larger variance, while oversmoothing increases bias. Thus the canonical scaling is optimal for first-order mean-squared error and yields first-order coverage up to the explicit bias cushion.
Proof.
Fix $i\neq j$ and write $f_{u}(x)=W(u,x)$. Let $N=N_{i}^{(-j)}$ and define
$D(u,u^{\prime})=\int_{0}^{1}(f_{u}-f_{u^{\prime}})^{2}.$
First we show that conditional on $(U_{i},U_{k})$, the variables $(P_{i\ell}-P_{k\ell})^{2}$ are i.i.d. in $\ell$, bounded in $[0,1]$, and have mean $D(U_{i},U_{k})$. Hoeffding’s inequality and a union bound over $k\in N$ (using $|N|\leq h_{n}\asymp\sqrt{n\log n}$) give that with probability $1-n^{-c}$,
$\sup_{k\in N}\left|\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}-D(U_{i},U_{k})\right|\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.7)
Combining (3.6) and (3.7),
$\frac{1}{|N|}\sum_{k\in N}D(U_{i},U_{k})\leq C^{\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.8)
Now we control the average latent distance between $U_{i}$ and its neighbors in $N$. Let
$\mathcal{C}_{i}=\{k\in N:U_{k}\text{ lies in a different block than }U_{i}\}.$
Assumption 3.5(ii) gives $D(U_{i},U_{k})\geq\eta^{2}\delta$ for $k\in\mathcal{C}_{i}$. Using (3.8),
$\frac{|\mathcal{C}_{i}|}{|N|}\eta^{2}\delta\leq C^{\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2},\qquad\Rightarrow\qquad\frac{|\mathcal{C}_{i}|}{|N|}\leq C^{\prime\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Decompose
$\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}=\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}+\frac{1}{|N|}\sum_{k\in\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}.$
The second term is at most $|\mathcal{C}_{i}|/|N|$. For $k\in N\setminus\mathcal{C}_{i}$ (same block), Assumption 3.5(iii) gives $D(U_{i},U_{k})\geq\kappa(\varepsilon)|U_{i}-U_{k}|^{2}$, hence
$\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}\leq\frac{1}{\kappa(\varepsilon)}\cdot\frac{1}{|N|}\sum_{k\in N}D(U_{i},U_{k}).$
Using (3.8) and $(\dagger)$,
$\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.8)
Step 3: Bias bound. Lipschitz continuity in the first argument gives
$|P_{kj}-P_{ij}|\leq L|U_{k}-U_{i}|.$
Thus
$|B^{(-j)}_{ij}|=\left|\frac{1}{|N|}\sum_{k\in N}(P_{kj}-P_{ij})\right|\leq L\sqrt{\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}}\leq Cr_{n},$
using (3.8) and $r_{n}=(\log n/n)^{1/4}$. ∎
Lemma 3.11 establishes a bias bound for the leave-one-out neighborhood smoother, showing that under the piecewise bi-Lipschitz smoothness assumption, the bias term $\sup_{\begin{subarray}{c}i\neq j:\\ (i,j)\in\mathcal{E}_{\varepsilon}\end{subarray}}|B_{ij}^{(-j)}|$ is of order $r_{n}=(\log n/n)^{1/4}$ for nodes $i$ and $j$ whose latent positions lie in the interior of the same block. This result relies on the control of the average two-hop distance between node $i$ and its neighbors in the leave-one-out neighborhood, as well as on the Lipschitz continuity of the graphon within blocks.
Remark 3.12.
Under the canonical choice $h_{n}\asymp\sqrt{n\log n}$, the stochastic fluctuation scale $h_{n}^{-1/2}$ and the bias scale $r_{n}=(\log n/n)^{1/4}$ are of the same order. Thus the normal approximation is first-order accurate, and bias cannot be ignored without undersmoothing. Choosing a smaller bandwidth would reduce bias at the cost of increased variance, while larger bandwidths inflate bias. The canonical scaling balances these two effects.
3.2 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
We introduce the conditional variance for $U_{ij}^{(-j)}$:
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.9)
so that $V_{ij}\leq 1/(4h)$. While $V_{ij}$ is analytically useful and measurable with respect to the $\sigma$-field generated by $(\mathcal{N}_{i}^{(-j)},P)$, it depends on the unknown latent probabilities $\{P_{kj}\}$.
To obtain tight, variance-adaptive control of $U_{ij}^{(-j)}$ using only observable quantities, we define the (Bessel-corrected) sample variance of the edges in the leave-one-out neighborhood:
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}),$ (3.10)
where $\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical mean.
Instead of relying on classical Bernstein bounds that require the true variance $V_{ij}$, we utilize the Empirical Bernstein inequality (Theorem 11 of Maurer and Pontil (2009)), which directly incorporates the sample variance $\hat{\sigma}_{ij}^{2}$.
For the normal-approximation intervals we estimate the conditional variance
$V_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj})$
by the plug-in estimator
$\widehat{V}_{ij}:=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}),$
which replaces the unknown probabilities $P_{kj}$ by their smoothed estimates.
Lemma 3.13 (Empirical Bernstein inequality).
Assume $h_{n}\geq 2$. For any $\delta\in(0,1)$,
$\mathbb{P}\!\left(|U_{ij}^{(-j)}|\geq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\ln(2/\delta)}{h_{n}}}+\frac{7\ln(2/\delta)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\leq\delta.$
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Moreover, by the leave-one-out construction, $\mathcal{N}_{i}^{(-j)}$ depends only on $A^{(-j)}$ and is independent of $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ given $P$.
Their empirical mean is
$\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},$
with conditional expectation
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}.$
Thus $U_{ij}^{(-j)}$ is exactly the deviation of the empirical mean from its conditional expectation.
Applying Theorem 11 of Maurer and Pontil (2009) to the variables $X_{k}:=A_{kj}$ and again to $1-X_{k}$ yields the stated two-sided inequality. ∎
Lemma 3.13 provides a strictly finite-sample, variance-adaptive control of $U_{ij}^{(-j)}$ directly in terms of the empirical variance $\hat{\sigma}_{ij}^{2}$, thereby bypassing the need for asymptotic plug-in arguments.
3.3 Berry-Esseen normal approximation
We now establish a normal approximation for the variance component $U_{ij}^{(-j)}$ defined in (3.2). Recall
$U_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$
and define the conditional variance
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}\mid\mathcal{N}_{i}^{(-j)},P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
so that
$V_{ij}=\mathrm{Var}(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P)=\frac{1}{h_{n}^{2}}\sigma_{ij}^{2}.$
Under Assumption 2.2, there exists $\eta\in(0,1/2)$ such that
$\eta\leq P_{kj}\leq 1-\eta\qquad\text{for all }k,j.$
Hence
$\eta(1-\eta)\leq P_{kj}(1-P_{kj})\leq\frac{1}{4}.$
Since $|\mathcal{N}_{i}^{(-j)}|=h_{n}$, it follows that
$\eta(1-\eta)\,h_{n}\leq\sigma_{ij}^{2}\leq\frac{1}{4}\,h_{n}.$ (3.11)
In particular,
$\sigma_{ij}\geq\sqrt{\eta(1-\eta)}\,h_{n}^{1/2},\qquad\sigma_{ij}^{-1}\leq C\,h_{n}^{-1/2}$ (3.12)
for a constant $C>0$ independent of $(i,j,n)$. Consequently,
$V_{ij}=\frac{\sigma_{ij}^{2}}{h_{n}^{2}}\asymp h_{n}^{-1},$
and the conditional variance of $U_{ij}^{(-j)}$ does not degenerate.
Lemma 3.14 (Uniform plug-in control for $\widehat{V}_{ij}$).
Recall
$V_{ij}=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),\qquad\widehat{V}_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}).$
Then for all $i\neq j$,
$\big|\widehat{V}_{ij}-V_{ij}\big|\leq\frac{1}{h_{n}}\,\sup_{1\leq a,b\leq n}\big|\widehat{P}_{ab}-P_{ab}\big|.$
Consequently, if one had uniform sup-norm consistency
$\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|=o_{p}(1),$
then the plug-in variance estimator $\widehat{V}_{ij}$ would inherit ratio-consistency. However, we do not establish such a sup-norm bound in this paper. Rowwise mean-squared error control does not imply uniform entrywise consistency. Therefore we do not claim ratio-consistency of $\widehat{V}_{ij}$, and normal intervals using $\widehat{V}_{ij}$ incur an additional, unquantified variance-estimation error.
Proof.
The map $\psi(x)=x(1-x)$ is $1$-Lipschitz on $[0,1]$ since $|\psi^{\prime}(x)|=|1-2x|\leq 1$. Hence for any $k$,
$\big|\widehat{P}_{kj}(1-\widehat{P}_{kj})-P_{kj}(1-P_{kj})\big|\leq|\widehat{P}_{kj}-P_{kj}|\leq\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|.$
Summing over $k\in\mathcal{N}_{i}^{(-j)}$ and dividing by $h_{n}^{2}$ gives
$|\widehat{V}_{ij}-V_{ij}|\leq\frac{1}{h_{n}^{2}}\cdot h_{n}\cdot\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|=\frac{1}{h_{n}}\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|.$
If $V_{ij}\asymp h_{n}^{-1}$ uniformly (e.g., under Assumption 2.2), then dividing by $V_{ij}$ yields the ratio consistency claim. ∎
Remark 3.15.
Our theoretical results establish rowwise mean-squared error control but do not prove uniform sup-norm consistency of $\widehat{P}$. Accordingly, plug-in variance intervals based on $\widehat{V}_{ij}$ should be interpreted as heuristic refinements rather than formally guaranteed procedures.
To control the additional error induced by replacing $V_{ij}$ by $\widehat{V}_{ij}$, we use Lemma 3.14, which reduces variance plug-in accuracy to the uniform estimation error of $\widehat{P}$.
Theorem 3.16 (Uniform Berry-Esseen bound).
Suppose Assumptions 2.1 and 2.2 hold, and let $h_{n}\to\infty$ as $n\to\infty$. Then there exists a universal constant $C>0$, independent of $i,j,n$, such that
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2}.$ (3.13)
Proof.
Fix $(i,j)$ and condition on $(\mathcal{N}_{i}^{(-j)},P)$. Define
$X_{k}:=A_{kj}-P_{kj},\qquad k\in\mathcal{N}_{i}^{(-j)}.$
By Proposition 2.7, the variables $\{X_{k}\}$ are independent, mean zero and bounded in [1,1][-1,1], but not necessarily identically distributed.
Let
$S:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k},\qquad\sigma_{ij}^{2}=\mathrm{Var}(S\mid\mathcal{N}_{i}^{(-j)},P).$
Define the third absolute moment sum
$\rho_{3}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathbb{E}\!\left[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P\right].$
Since $|X_{k}|\leq 1$, we have
$\rho_{3}\leq\sum_{k}\mathbb{E}[X_{k}^{2}\mid\mathcal{N}_{i}^{(-j)},P]=\sigma_{ij}^{2}.$
The Berry-Esseen inequality for independent, not necessarily identically distributed summands (see, e.g., Petrov (1975, Chapter V, Theorem 3)) states that there exists a universal constant $C_{0}>0$ such that
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\frac{\rho_{3}}{\sigma_{ij}^{3}}.$
Using $\rho_{3}\leq\sigma_{ij}^{2}$ yields
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\sigma_{ij}^{-1}.$
By (3.12), $\sigma_{ij}^{-1}\leq Ch_{n}^{-1/2}$, where $C$ is independent of $(i,j)$. Finally, since $U_{ij}^{(-j)}=h_{n}^{-1}S$ and $V_{ij}=h_{n}^{-2}\sigma_{ij}^{2}$, we have the identity
$\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}=\frac{S}{\sigma_{ij}},$
and the uniform bound (3.13) follows. ∎
Corollary 3.17 (Uniform asymptotic normality).
Under the conditions of Theorem 3.16,
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|\longrightarrow 0\qquad\text{as }n\to\infty.$
Proof.
For fixed $(i,j)$ and $x\in\mathbb{R}$, define
$\Delta_{ij}(x):=\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x).$
By Theorem 3.16,
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}|\Delta_{ij}(x)|\leq Ch_{n}^{-1/2}\qquad\text{a.s.}$
Taking expectations and using Jensen’s inequality for $|\,\cdot\,|$,
$\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|=\left|\mathbb{E}[\Delta_{ij}(x)]\right|\leq\mathbb{E}[|\Delta_{ij}(x)|]\leq Ch_{n}^{-1/2}.$
Taking the supremum over $x\in\mathbb{R}$ and $(i,j)$ yields
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2},$
and since $h_{n}\to\infty$, the right-hand side tends to 0. ∎
3.4 Confidence intervals
Under the canonical choice $h_{n}\asymp\sqrt{n\log n},$ we have
$h_{n}^{-1/2}=\left(\sqrt{n\log n}\right)^{-1/2}=(n\log n)^{-1/4}.$
Therefore the Berry-Esseen remainder in (3.13) satisfies
$\sup_{i,j}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|=O\!\big((n\log n)^{-1/4}\big).$
By Lemma 3.11, the bias component satisfies Consequently, Corollary 3.9 yields a uniform bias bound over $(i,j)\in\mathcal{E}_{\varepsilon}$. Hence the stochastic fluctuation scale and the bias scale are of the same order, up to logarithmic factors. In particular, to obtain asymptotically valid normal-approximation confidence intervals for $P_{ij}$, one must either undersmooth (so that the bias is $o(h_{n}^{-1/2})$) or incorporate an explicit bias correction of order $(\log n/n)^{1/4}$.
We construct two families of confidence intervals for $P_{ij}$: (i) finite-sample, variance-adaptive bands based on Lemma 3.13, and (ii) asymptotic normal-approximation bands based on Theorem 3.16 together with a bias cushion from Lemma 3.11.
3.4.1 Finite-sample Confidence Intervals via Empirical Bernstein Bounds
To construct finite-sample confidence intervals for the variance component $U_{ij}^{(-j)}$, we require a concentration inequality that depends only on observable quantities. Classical Bernstein inequalities involve the true variance $V_{ij}$, which is not directly accessible. In contrast, the Empirical Bernstein bound of Maurer and Pontil (2009) provides a deviation inequality that replaces the unknown variance with the sample variance, yielding a fully data-driven and nonasymptotic confidence interval.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Their empirical mean is the leave-one-out estimator
$\tilde{P}_{ij}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}.$ (3.14)
The unbiased sample variance of the observed edges in the neighborhood is
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}.$ (3.15)
Since the edges are binary, this expression simplifies to
$\hat{\sigma}_{ij}^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}).$ (3.16)
We apply Theorem 11 of Maurer and Pontil (2009) to the independent bounded variables $X_{k}:=A_{kj}$, $k\in\mathcal{N}_{i}^{(-j)}$, with sample size $h_{n}$. For any $\alpha\in(0,1)$, a union bound over the upper and lower tails yields the conditional deviation inequality
$\mathbb{P}\!\left(\big|U_{ij}^{(-j)}\big|\leq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}\;\middle|\;\mathcal{N}_{i}^{(-j)},P\right)\geq 1-\alpha.$ (3.17)
We therefore define the variance-adaptive half-width
$w_{ij}^{(B)}(\alpha):=\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}.$ (3.18)
The resulting finite-sample confidence interval for the localized average is
$CI_{ij}^{(B)}(\alpha):=\Big[\tilde{P}_{ij}-w_{ij}^{(B)}(\alpha),\;\tilde{P}_{ij}+w_{ij}^{(B)}(\alpha)\Big]\cap[0,1].$ (3.19)
By construction, conditionally on $(\mathcal{N}_{i}^{(-j)},P)$, this interval captures the neighborhood average
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$
with probability at least $1-\alpha$ for any finite $n$.
Theorem 3.18 (Finite-sample inference for $P_{ij}$ on the good set).
Fix $\varepsilon>0$, $\alpha\in(0,1)$, and $(i,j)\in\mathcal{E}_{\varepsilon}$. Let $CI^{(B)}_{ij}(\alpha)$ be the empirical Bernstein interval defined in (3.17) (using the sample variance form). On the event that the neighborhood regularity condition (3.4) holds (probability at least $1-n^{-c}$ as in Corollary 3.9), we have the bias cushion
$\left|P_{ij}-\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}\right|\leq C(\varepsilon,C_{2})\,r_{n}.$
Moreover, by the empirical Bernstein deviation (Maurer and Pontil (2009)) in (3.17)
$\mathbb{P}\!\left(\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}\in CI^{(B)}_{ij}(\alpha)\right)\geq 1-\alpha.$
Hence, on the intersection of these two events,
$P_{ij}\in CI^{(B)}_{ij}(\alpha)\oplus[-C(\varepsilon,C_{2})r_{n},\;C(\varepsilon,C_{2})r_{n}].$
Consequently,
$\mathbb{P}\!\big(P_{ij}\in CI^{(B)}_{ij}(\alpha)\oplus[-C(\varepsilon,C_{2})r_{n},\;C(\varepsilon,C_{2})r_{n}]\big)\geq 1-\alpha-n^{-c}.$
Remark 3.19 (Target of the Empirical Bernstein interval).
It is important to note that $CI_{ij}^{(B)}(\alpha)$ is centered at the empirical neighborhood mean $\tilde{P}_{ij}$ and therefore provides finite-sample coverage for the localized average
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj},$
not for the individual probability $P_{ij}$. The distinction reflects the fact that the Empirical Bernstein inequality controls only the stochastic fluctuation component $U_{ij}^{(-j)}$ and does not address the bias term $B_{ij}^{(-j)}$.
Remark 3.20 (Advantage of the Empirical Bound).
Unlike standard concentration inequalities that require substituting an estimated variance $\hat{V}_{ij}$ for the true variance $V_{ij}$, a step that typically degrades finite-sample guarantees and forces reliance on Slutsky’s Theorem and asymptotic normality, the Empirical Bernstein interval completely avoids the cost of studentization. The interval $CI_{ij}^{(B)}(\alpha)$ automatically adapts to the empirical variance of the neighborhood while strictly preserving its $1-\alpha$ coverage probability for any finite sample size.
3.4.2 Normal‑approximation intervals
Treat $U_{ij}$ as approximately normal and control the bias $B_{ij}$ via Lemma 3.11. For $\alpha\in(0,1)$ let $z_{1-\alpha/2}$ denote the $(1-\alpha/2)$-quantile of the standard normal and define
$w_{ij}^{(N)}(\alpha):=z_{1-\alpha/2}\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}\Big(\frac{\log n}{n}\Big)^{1/4},$ (3.20)
where $c_{\mathrm{bias}}>0$ is a bias‑cushion constant chosen large enough to dominate the bias in Lemma 3.11. The normal interval is
$\mathrm{CI}_{ij}^{(N)}(\alpha):=[\tilde{P}_{ij}-w_{ij}^{(N)}(\alpha),\ \tilde{P}_{ij}+w_{ij}^{(N)}(\alpha)]\cap[0,1].$ (3.21)
By Theorem 3.16, the uniform Berry-Esseen bound implies that
$\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}\left|\mathbb{P}\!\big(P_{ij}\in\mathrm{CI}_{ij}^{(N)}(\alpha)\big)-(1-\alpha)\right|=O\!\big(h_{n}^{-1/2}\big).$
Remark 3.21.
The Berry-Esseen bound for the variance component holds uniformly over all $(i,j)$. However, coverage statements for $P_{ij}$ require bias control and therefore hold uniformly only over $(i,j)\in\mathcal{E}_{\varepsilon}$.
These intervals are asymptotically valid for the variance component $U_{ij}^{(-j)}$ uniformly in $(i,j)$, and for $P_{ij}$ uniformly over $(i,j)\in\mathcal{E}_{\varepsilon}$ up to the bias cushion. Under the canonical choice $h_{n}\asymp\sqrt{n\log n}$ this becomes $O\!\big((n\log n)^{-1/4}\big)$. The bias cushion in (3.20) ensures that the bias term from Lemma 3.11 is asymptotically negligible at the same scale. Normal intervals using $\widehat{V}_{ij}$ therefore incur an additional variance-estimation error term, as discussed in Lemma 3.14.
Remark 3.22.
The Empirical Bernstein bands (3.19) are finite-sample and variance-adaptive but can be conservative. The normal bands (3.21) are typically tighter but rely on asymptotic approximation and an explicit bias cushion; the two constructions provide complementary guarantees in practice.
Algorithm 1 Simulation Procedure for LOO Smoothing and Confidence Intervals
1:Sample size $n$, neighborhood size $h_{n}$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
2:Estimated matrices $\widehat{P}_{\mathrm{LOO}},\widehat{P}_{\mathrm{class}}$ and empirical coverage rates
3:Draw $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$ and set $P_{ij}=f(\xi_{i},\xi_{j})$
4:Sample $A_{ij}\mid P_{ij}\sim\mathrm{Bernoulli}(P_{ij})$ for $i<j$ and symmetrize
5:For each $j$ form $A^{(-j)}$, compute distances $d_{ik}^{(-j)}$ via (2.2)
6:Build LOO neighborhoods $\mathcal{N}_{i}^{(-j)}$ of size $h_{n}$ and classical neighborhoods $\mathcal{N}_{i}$ from full AA
7:Compute $\tilde{P}_{ij}$ and $\widehat{P}_{ij}$ via (2.4) for both LOO and classical smoothers
8:Estimate $\widehat{V}_{ij}$ and construct intervals $\mathrm{CI}_{ij}^{(B)}(\alpha)$, $\mathrm{CI}_{ij}^{(N)}(\alpha)$
9:Evaluate empirical coverage $\widehat{\mathrm{Cov}}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}$ and report MSEs
4 Honest tuning via cross-validation
The leave-one-out (LOO) construction admits a simple and honest cross-validation (CV) scheme for selecting the neighborhood size $h_{n}$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held-out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h_{n})$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h_{n})$. Consequently, the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard concentration and normal approximation results, enabling valid CV-based tuning of $h_{n}$.
We use the term “honest” to denote that no held-out edge $A_{ij}$ is used in constructing its predictor $\tilde{P}_{ij}^{(-j)}$. This eliminates double use of data but does not imply oracle optimality of the selected bandwidth.
4.1 Cross-validation procedure and notation
Fix a row index $i\in[n]$. Let $\mathcal{G}$ be a finite grid of candidate neighborhood sizes,
$\mathcal{G}\subset\{1,2,\dots,n-2\},$
typically chosen around the theoretical scale $h_{n}\asymp\sqrt{n\log n}$. For each $h_{n}\in\mathcal{G}$ and each held-out column $j\neq i$, form the reduced adjacency matrix $A^{(-j)}$ by deleting the $j$-th row and column. This ensures that the neighborhood $\mathcal{N}_{i}^{(-j)}(h_{n})$ is constructed without using any information from the held-out target column $j$. Compute the corresponding leave-one-out predictor
$\tilde{P}_{ij}^{(-j)}(h_{n}):=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h_{n})}A_{kj},\qquad j\neq i.$
Define the LOO cross-validation score
$\widehat{R}_{i}(h)=\frac{1}{n-1}\sum_{j\neq i}(A_{ij}-\tilde{P}_{ij}^{(-j)}(h))^{2}.$
Because the neighborhood $\mathcal{N}_{i}^{(-j)}(h)$ is constructed from $A^{(-j)}$ (which omits all edges incident to vertex $j$) and $i\notin\mathcal{N}_{i}^{(-j)}(h)$ by construction, the held‑out edge $A_{ij}$ is conditionally independent of the predictor $\tilde{P}_{ij}^{(-j)}(h)$ given $P$. Consequently, $\widehat{R}_{i}(h)$ is (conditionally) unbiased for the corresponding prediction risk. A full oracle inequality analysis of this tuning rule is not included in this version.
The corresponding prediction risk is the conditional expectation of the same quantity. Because $A_{ij}$ is conditionally independent of $\mathcal{N}_{i}^{(-j)}(h_{n})$ given $P$, this evaluates to
$R_{i}(h_{n}):=\frac{1}{n-1}\sum_{j\neq i}\mathbb{E}\!\left[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h_{n})\bigr)^{2}\,\middle|\,\mathcal{N}_{i}^{(-j)}(h_{n}),P\right].$ (4.1)
We then select any minimizer of the empirical score over the grid,
$\hat{h}_{n}\in\arg\min_{h_{n}\in\mathcal{G}}\widehat{R}_{i}(h_{n}).$
5 Simulation Study
We conducted a simulation study to evaluate the finite-sample performance of the proposed leave-one-out (LOO) neighborhood smoother and compare it with the classical (ZLZ) neighborhood smoother. All computations were implemented in Python using NumPy with a fixed random seed.
5.1 Setup
We considered five representative graphon models capturing a range of smoothness and structural complexity:
Table 1: Graphon Model Definitions
Model Definition
Smooth $0.5+0.3\sin(\pi u)\sin(\pi v)$
Block 2-block SBM ($p_{\text{in}}=0.7,p_{\text{out}}=0.3$)
Wiggly $0.5+0.25\sin(4\pi uv)$
Rank-one $(u+v)/2$
Spiky $0.2+0.8\,\mathbbm{1}\{|u-0.5|<0.1,\ |v-0.5|<0.1\}$
Remark 5.1.
The empirical coverage levels reported in Table 2 are higher than typical for $n=200$. This reflects the favorable latent geometry and relatively generous bias cushion used in our simulation design. In more challenging settings, we expect coverage to be closer to the nominal level.
For each graphon, we generated an undirected graph with $n=200$ nodes by sampling latent positions $\xi_{i}\sim\mathrm{Unif}[0,1]$ and edges $A_{ij}\sim\mathrm{Bernoulli}(f(\xi_{i},\xi_{j}))$ independently for $i<j$. The neighborhood size and trimming quantile were fixed at
$h_{n}=\lfloor 1.5\sqrt{n\log n}\rfloor.$
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column and row $j$ removed, thereby restoring conditional independence between the neighborhood and the target edge. Both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$.
5.2 Evaluation metrics
We report two performance measures. For node $i=0$,
$\mathrm{MSE}_{0}=\frac{1}{n}\sum_{j=1}^{n}(\widehat{P}_{0j}-P_{0j})^{2}.$
For the LOO estimator, we consider:
  • Empirical Bernstein intervals:
    $w_{ij}^{(B)}=\sqrt{2\widehat{V}_{ij}\log(4/\alpha)}+\frac{\log(4/\alpha)}{3h},$
  • Normal intervals:
    $w_{ij}^{(N)}=1.96\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}(\log n/n)^{1/4},\qquad c_{\mathrm{bias}}=0.1,$
where the plug-in variance proxy is
$\widehat{V}_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}).$
Intervals are truncated to $[0,1]$. Coverage is computed as
$\mathrm{Coverage}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}.$
5.3 Results
Table 2 summarizes the results for all five graphons. The LOO smoother achieves a comparable MSE to the classical ZLZ smoother. Coverage for both confidence interval constructions is close to the nominal $95\%$ level, with the Empirical Bernstein intervals slightly more conservative.
Graphon $\mathrm{MSE}_{0}^{\mathrm{LOO}}$ $\mathrm{MSE}_{0}^{\mathrm{ZLZ}}$ $\mathrm{Cov}_{\mathrm{Bousquet}}$ $\mathrm{Cov}_{\mathrm{Normal}}$
Smooth 0.0027 0.0030 0.999 0.995
Block 0.0021 0.0027 1.000 1.000
Wiggle 0.0031 0.0030 0.944 0.885
Rank-one 0.0029 0.0030 0.999 0.996
Spiky 0.0021 0.0033 0.977 0.977
Table 2: Row-wise MSE and empirical coverage for five graphons at $n=200$.
6 Discussion and Extensions
This paper develops a principled framework for inference on individual edge probabilities in graphon models based on a simple leave-one-out modification of neighborhood smoothing. By constructing LOO neighborhoods that do not depend on the column being predicted, we restore conditional independence of the centered edge variables and thereby make classical concentration and normal-approximation tools available. The resulting methodology delivers entrywise confidence intervals and an honest cross-validation scheme for data-driven tuning, while retaining the optimal row-wise mean-squared error rate known for classical neighborhood smoothing.
The LOO construction is conceptually simple and quite general, which makes it a promising building block for extensions beyond the basic undirected, unweighted, static setting considered here. we outline several directions for future work.
  1. 1.
    Directed and weighted graphons. For directed networks, one could replace the Zhang et al. (2017) distance by an asymmetric distance that incorporates both in and out neighborhood information, and then apply the same LOO decoupling on the relevant columns. For weighted networks, a similar strategy could be pursued using appropriate moment-based distances built from $A$ and $A^{2}$, with boundedness or tail assumptions tailored to the edge-weight distribution.
  2. 2.
    Extension to sparse regimes. For models of the form $P_{ij}=\rho_{n}W(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$, the effective variance of $U_{ij}^{(-j)}$ shrinks with $\rho_{n}$. Normal approximation would require renormalization by $\sqrt{h_{n}\rho_{n}}$, and empirical Bernstein bounds must account for rare-event behavior. While the leave-one-out decoupling remains valid, the concentration and Berry-Esseen analysis would require substantially different arguments.
  3. 3.
    Degree correction and heterogeneous sparsity. Many real networks exhibit highly heterogeneous degree distributions. Extending the present analysis to degree corrected graphon or blockmodel formulations would require incorporating node-specific scale factors into both the distance and the bias analysis. The LOO idea remains applicable, but the variance proxy $V_{ij}$ and the bias bounds would need to be adapted to handle varying effective sample sizes across nodes.
  4. 4.
    Adaptive and simultaneous confidence bands. While we have focused on pointwise confidence intervals for individual edge probabilities, the same machinery can be used to construct simultaneous bands over subsets of rows or columns, for example by combining the Berry-Esseen bounds with high-dimensional Gaussian approximation results. It would also be natural to let the bias cushion constant $c_{\mathrm{bias}}$ be selected in a data-driven way, for instance by cross-validation or stability criteria, in order to balance coverage and tightness of the intervals.
  5. 5.
    Applications to real networks. Finally, it would be valuable to apply the proposed methodology to real data sets, such as social, citation, or biological networks, in order to assess its practical performance and interpretability. In particular, the ability to attach uncertainty quantification to individual edge probabilities or to row-wise summaries could provide new insights in applied settings where only a single large network is observed.
Identifiability limitations.
The within-block $L^{2}$ identifiability condition in Assumption 3.5 is structurally stronger than conditions used solely for MSE analysis. It is required here to translate rowwise $L^{2}$ closeness into entrywise bias control. Extending the inference theory to settings with weaker or purely combinatorial identifiability (e.g., high-$K$ block models) remains an open problem.
The main limitation of the current approach is the requirement of piecewise Lipschitz smoothness for the underlying graphon. Extensions to rougher kernels, highly discontinuous structures, or models with sharp community boundaries would likely require modified distance metrics and a more delicate bias analysis. Addressing these challenges, while retaining tractable inference procedures, remains an important open problem.