### v5

# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

Abstract
Estimating latent edge probabilities from a single observed network is a fundamental challenge in statistical network analysis. While classical neighborhood smoothing methods provide consistent point estimates, providing valid entrywise uncertainty quantification remains a complex problem due to the intricate dependencies formed during data-driven neighborhood selection. In this paper, we propose a novel Leave-One-Out (LOO) neighborhood smoothing estimator that explicitly decouples these dependencies. By strictly excluding the target node’s column during neighborhood formation, our approach unlocks exact conditional independence while preserving the statistical efficiency of the local neighborhood. We rigorously establish the bias-variance trade-off of this estimator by deriving a conditional risk bound, demonstrating that the stochastic cross-term vanishes identically. Leveraging this geometric conditional independence, we construct valid entrywise confidence intervals through two distinct frameworks: strictly non-asymptotic empirical Bernstein bounds for finite-sample guarantees, and a Central Limit Theorem for sharp asymptotic inference under an undersmoothing regime. Extensive simulation studies across diverse graphon topologies confirm that our asymmetric LOO estimator matches the point estimation performance of classical methods while successfully achieving nominal coverage rates for entrywise uncertainty quantification.
1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2008; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $f$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, a number of estimators such as USVT (Chatterjee, 2015), sorting-based methods (Chan and Airoldi, 2014), and block-based estimators (Gao et al., 2015), achieve near-optimal or minimax rates under appropriate structural assumptions.
Among these, the neighborhood smoothing method introduced by Zhang et al. (2017) stands out. By adaptively aggregating structurally similar nodes based on empirical two-hop distances, neighborhood smoothing achieves minimax (or rate-optimal) MSE under their structural assumptions.
However, a fundamental gap remains: the transition from optimal point estimation to rigorous statistical inference. Given an entrywise estimate $\widehat{P}_{ij}$, it is critically important for downstream scientific tasks to attach a valid measure of uncertainty to this prediction. Yet, existing inferential theory for networks has focused almost exclusively on low-dimensional or global functionals. Significant progress has been made in proving limit theorems for spectral embeddings (Sussman et al., 2012; Athreya et al., 2016), asymptotic normality for community parameters in SBMs (Bickel et al., 2013), and global goodness-of-fit testing (Lei, 2016). Conversely, entrywise uncertainty quantification-specifically, constructing valid confidence intervals and proving central limit theorems for individual edge probabilities under continuous graphons-has remained theoretically elusive. Recent attempts to bridge this gap have relied on computationally intensive resampling methods, such as network bootstrapping (Levin and Levina, 2025; Green and Shalizi, 2022), or generalized fiducial inference (Su et al., 2022), but exact distribution-free analytical bounds for neighborhood estimators are lacking.
The obstacle preventing entrywise inference is not merely technical, but structural. To achieve optimal MSE, neighborhood smoothing estimators (Zhang et al., 2017) rely on a data-driven aggregation of similar nodes. This creates a severe “double-dipping” problem: the procedure reuses the adjacency matrix twice. The matrix $A$ is first used to compute pairwise topological distances to select a neighborhood $\mathcal{N}_{i}$, and then the very same matrix $A$ is used to average the edges within that neighborhood. Consequently, the random neighborhood weights depend directly on the edges being averaged. Even after conditioning on the latent probability matrix $P$, the summands in the estimator are strongly coupled. This intractable dependence invalidates the foundational assumptions of classical concentration inequalities and Berry-Esseen normal approximations.
In this paper, we demonstrate that this probabilistic bottleneck can be cleanly bypassed via a simple but highly principled structural modification: leave-one-out (LOO) neighborhood smoothing.
When estimating a single entry $P_{ij}$, we perform a topological surgery on the graph by constructing the neighborhood of node $i$ using a reduced adjacency matrix $A^{(-j)}$ in which the $j$-th row and column have been entirely deleted. All pairwise distances and nearest-neighbor selections are computed on this reduced matrix. Only after the neighborhood $\mathcal{N}_{i}^{(-j)}$ is fixed do we average the original target edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$.
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling permits the application of classical empirical process tools to graphon neighborhood smoothing. Specifically, assuming the true graphon is dense and comprises a fixed number of well-separated, piecewise Lipschitz blocks, we leverage this restored independence to establish three main contributions:
  1. 1.
    Variance-Adaptive Concentration: We derive tight, finite-sample concentration inequalities based on empirical Bernstein bounds that automatically adapt to the local variance of the edge probabilities.
  2. 2.
    Asymptotic Normality: We establish a uniform Berry-Esseen bound with explicit rates for the normalized variance component; normal-approximation intervals in this version use a heuristic variance estimate and incur an additional variance-estimation error term.
  3. 3.
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$. Here, honest cross-validation means that it does not reuse held-out edges in predictor construction, ensuring valid inference even after data-driven tuning.
Crucially, these inferential guarantees preserve the optimal rowwise mean-squared error rate of classical neighborhood smoothing.
The remainder of the paper is organized as follows. Section 2 formalizes the graphon model, the Zhang-Levina-Zhu (ZLZ), Zhang et al. (2017) empirical distance, and our leave-one-out neighborhood construction. Section 3 establishes the foundational conditional independence, proving the variance-adaptive concentration bounds, Berry-Esseen normal approximations, and the validity of the resulting confidence intervals. Section 5 details the honest cross-validation scheme for tuning parameter selection. Section 6 reports comprehensive simulation results illustrating the finite-sample coverage and MSE performance, and Section 7 concludes with a discussion of extensions to directed, weighted, and degree-corrected settings.
2 Model and Leave-One-Out Neighborhood Smoothing
We observe a single undirected simple graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$, no self-loops ($A_{ii}=0$), and independent upper-triangular entries
$A_{ij}\mid P_{ij}\ \sim\ \mathrm{Bernoulli}(P_{ij}),\qquad 1\leq i<j\leq n,$ (2.1)
with $A_{ij}=A_{ji}$. The matrix $P=(P_{ij})$ contains unknown edge probabilities. We adopt the standard graphon parametrization below.
Assumption 2.1 (Graphon model and smoothness).
The edge probabilities satisfy $P_{ij}=f(\xi_{i},\xi_{j})$ with $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$. The graphon $f:[0,1]^{2}\to[0,1]$ is piecewise Lipschitz in each argument with Lipschitz constant $L$ on a fixed finite rectangular partition of $0,1]^{2}$.
Assumption 2.2 (Non-degeneracy).
There exists $\eta\in(0,1/2)$ such that
$\eta\leq f(u,v)\leq 1-\eta\qquad\text{for all }(u,v)\in[0,1]^{2}.$
Assumption 2.2 places the model in the dense graphon regime, where edge probabilities are uniformly bounded away from zero and one. This excludes sparse scalings of the form $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$. Our Berry-Esseen and empirical Bernstein arguments rely critically on variance bounded away from zero. Extending the present theory to sparse or degree-corrected regimes would require variance renormalization and additional control of small-probability events, and is left for future work.
Remark 2.3.
Assumption 2.2 guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance-adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re-normalization and additional control of rare events, which we leave to future work.
Remark 2.4.
The rectangular partition in Assumption 2.1 is fixed and does not grow with $n$. All constants in our rates may depend on the Lipschitz constant $L$ and on the partition, but not on $n$.
Remark 2.5 (Dense regime convention).
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=f(\xi_{i},\xi_{j})$ with $f$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two-hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ are discussed briefly in Section 7.
2.1 LOO neighborhoods and estimator
Consider the set of nodes $V=\{1,\dots,n\}$. Fix $j\in[n]$ and let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting row and column $j$ from $A$. Define the two-hop matrices
$M:=\frac{1}{n}A^{2},\qquad M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}.$
For distinct $i,k\in V\setminus\{j\}$, define the leave-one-out (LOO)-distance
[…]
Let
$d^{(-j)}_{i,(1)}\leq\cdots\leq d^{(-j)}_{i,(n-2)}$
denote the order statistics of the set
$\bigl\{d^{(-j)}(i,k):k\in V\setminus\{i,j\}\bigr\}.$
Define the leave-one-out neighborhood of node $i$ by selecting the $h_{n}$ nearest neighbors according to $d^{(-j)}$. To ensure the neighborhood size is deterministically fixed, we resolve any distance ties at the boundary by prioritizing smaller node indices. Because the arbitrary integer labels assigned to nodes are independent of the latent positions $\xi$, this lexicographic tie-breaking is strictly independent of the underlying graphon geometry and introduces zero statistical bias. Formally, let $\pi$ be the permutation of $V\setminus\{i,j\}$ that sorts the nodes such that $d^{(-j)}(i,\pi(1))\leq d^{(-j)}(i,\pi(2))\leq\dots$, with $\pi(a)<\pi(b)$ if $d^{(-j)}(i,\pi(a))=d^{(-j)}(i,\pi(b))$. We define:
$\mathcal{N}_{i}^{(-j)}=\big\{\pi(1),\pi(2),\dots,\pi(h_{n})\big\}.$ (2.3)
Here, $h_{n}\asymp\sqrt{n\log n}$ is a fixed bandwidth sequence. Because $\sqrt{(n-1)\log(n-1)}\asymp\sqrt{n\log n}$, explicitly applying $h_{n}$ on the (n1)(n-1)-node reduced graph preserves the theoretical scaling of the classical ZLZ neighborhoods while strictly fixing the LOO neighborhood size to exactly $|\mathcal{N}_{i}^{(-j)}|=h_{n}$. Following the notations in Zhang et al. (2017), given $\mathcal{N}_{i}^{(-j)}$, define the one-sided LOO smoother and its symmetrized version by
$\tilde{P}_{ij}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},\qquad\widehat{P}_{ij}:=\frac{\tilde{P}_{ij}+\tilde{P}_{ji}}{2}.$ (2.4)
While the point estimator is practically symmetrized via $\widehat{P}_{ij}$ to respect the undirected nature of the graph, averaging the two distinct neighborhoods $\mathcal{N}_{i}^{(-j)}$ and $\mathcal{N}_{j}^{(-i)}$ reintroduces complex dependencies. However, because point estimation risk easily transfers via the inequality $(\widehat{P}_{ij}-P_{ij})^{2}\leq\frac{1}{2}(\tilde{P}_{ij}-P_{ij})^{2}+\frac{1}{2}(\tilde{P}_{ji}-P_{ij})^{2}$, the symmetric estimator inherits the optimal mean-squared error rate. Despite this advantage, valid entrywise uncertainty quantification requires exact conditional independence. Therefore, our inferential limit theorems and confidence intervals are developed and constructed strictly for the fundamental one-sided LOO predictor $\tilde{P}_{ij}$. Therefore, our inferential limit theorems and confidence intervals are developed and constructed strictly for the fundamental one-sided LOO predictor $\tilde{P}_{ij}$.
For any $i\neq j$, the entrywise estimation error of the one-sided LOO smoother can be additively decomposed into a stochastic fluctuation and a deterministic bias. We write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}^{(-j)}+B_{ij}^{(-j)},$ (2.5)
where the variance component is defined as
$U_{ij}^{(-j)}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (2.6)
and the deterministic bias component is defined as
$B_{ij}^{(-j)}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}-P_{ij}.$ (2.7)
Define the distance function $d:V\times V\to\mathbb{R}_{+}$ by
$d_{ZLZ}(i,j):=\begin{cases}0,&\text{for }i=j,\\ \max_{k\neq i,j}\Big|\big(A^{2}/n\big)_{ik}-\big(A^{2}/n\big)_{jk}\Big|,&\text{for }i\neq j.\end{cases}$ (2.8)
This is the sup-norm discrepancy between the $i$th and $j$th rows of the normalized two-hop matrix $A^{2}/n$, measuring the similarity of their two-hop neighborhoods. Therefore the ZLZ neighborhood $\mathcal{N}_{i}$ consists of the $h_{n}$ nodes with the smallest ZLZ distance to $i$.
Following Zhang et al. (2017), let $q_{i}(h/n)$ denote the $(h/n)$-th empirical quantile of the distance set $\{d_{\mathrm{ZLZ}}(i,k):k\in V\setminus\{i\}\}$. The ZLZ neighborhood is defined as the set of nodes whose distance to $i$ falls below this quantile threshold:
$\mathcal{N}_{i}^{\mathrm{ZLZ}}:=\big\{k\in V\setminus\{i\}:d_{\mathrm{ZLZ}}(i,k)\leq q_{i}(h/n)\big\}.$ (2.9)
The ZLZ estimator is then defined by the empirical average over this neighborhood:
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|\mathcal{N}_{i}^{\mathrm{ZLZ}}|}\sum_{k\in\mathcal{N}_{i}^{\mathrm{ZLZ}}}A_{kj}.$
Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance-adaptive concentration or Berry-Esseen theory.
Remark 2.6.
The symmetrized estimator $\widehat{P}_{ij}$ respects the symmetry of the simple graph model. All concentration and normal-approximation results are derived for the one-sided quantities $\tilde{P}_{ij}$; symmetrization does not affect the independence structure.
Proposition 2.7 (Conditional independence).
Conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$, the variables
$\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$
are independent, mean-zero, and bounded in [1,1][-1,1].
The leave-one-out construction does not alter the smoothing mechanism itself; it alters only the dependency structure. By removing the target column prior to neighborhood selection, the summands in the neighborhood average become exactly independent conditional on $P$. This structural decoupling is the key step that allows classical empirical process tools to be applied without resampling or sample splitting.
Proof.
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=f(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$. By construction, $A^{(-j)}$ is obtained from $A$ by removing the $j$th row and column. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$. For each $k\in\mathcal{N}_{i}^{(-j)}$, we have $A_{kj}\sim\mathrm{Bernoulli}(P_{kj})$ and $A_{kj}-P_{kj}\in[-1,1]$ with mean zero. Since the edges $\{A_{kj}:k\neq j\}$ are independent conditional on $P$, and $\mathcal{N}_{i}^{(-j)}$ is independent of these edges, it follows that the collection $\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ is independent, mean-zero, and bounded in [1,1][-1,1] conditional on $(\mathcal{N}_{i}^{(-j)},P)$. ∎
3 Main Results: Entrywise Uncertainty Quantification
With the conditional independence of the leave-one-out (LOO) summands established in Proposition 2.7, we are now equipped to construct valid, entrywise confidence intervals. We separate the estimation error into a stochastic variance component, governed by the conditionally independent edges, and a deterministic bias component, which captures the approximation error of the local neighborhood average.
3.1 Finite-Sample Confidence Intervals via Empirical Bernstein Bounds
To construct nonasymptotic confidence intervals for the variance component $U_{ij}^{(-j)}$, we require a concentration inequality that depends only on observable quantities. Classical Bernstein inequalities involve the true variance $V_{ij}$, which is unknown. Instead, we utilize the Empirical Bernstein bound of Maurer and Pontil (2009), which provides a deviation inequality that replaces the unknown variance with the sample variance, yielding a fully data-driven interval.
The unbiased sample variance of the observed edges in the LOO neighborhood is given by
$s_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2},$ (3.1)
where $s_{ij}^{2}$ is the unbiased sample variance of the neighborhood observations. Because the observations $A_{kj}$ are binary indicators, this sample variance has the simplified exact form:
$s_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}).$ (3.2)
where $\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical LOO mean.
Lemma 3.1 (Empirical Bernstein bound).
Assume $h_{n}\geq 2$. For any $\alpha\in(0,1)$, the stochastic fluctuation $U_{ij}^{(-j)}$, defined in (2.6), satisfies
$\mathbb{P}\!\left(\big|U_{ij}^{(-j)}\big|\leq\sqrt{\frac{2s_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\geq 1-\alpha.$ (3.3)
Proof.
Let $s_{ij}^{2}$ denote the unbiased sample variance of the neighborhood observations:
$s_{ij}^{2}=\frac{1}{h_{n}-1}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(X_{k}-\tilde{P}_{ij})^{2}.$
By Theorem 11 of Maurer and Pontil (2009), for any $\delta\in(0,1)$, the one-sided empirical Bernstein inequality bounds the upper tail of the deviation with probability at least $1-\delta$:
$\mathbb{P}\!\left(U_{ij}^{(-j)}\leq\sqrt{\frac{2s_{ij}^{2}\log(2/\delta)}{h_{n}}}+\frac{7\log(2/\delta)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\geq 1-\delta.$ (3.4)
To bound the lower tail, we apply the identical theorem to the shifted random variables $Y_{k}:=1-X_{k}$. The empirical average and expectation shift by $1$, preserving the centered deviation up to a sign change: $\tilde{P}_{Y}-\mathbb{E}[\tilde{P}_{Y}\mid\mathcal{N}_{i}^{(-j)},P]=-U_{ij}^{(-j)}$. The sample variance is translation-invariant ($s_{Y}^{2}=s_{ij}^{2}$). This yields the symmetric lower tail bound:
$\mathbb{P}\!\left(-U_{ij}^{(-j)}\leq\sqrt{\frac{2s_{ij}^{2}\log(2/\delta)}{h_{n}}}+\frac{7\log(2/\delta)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\geq 1-\delta.$ (3.5)
Based on Lemma 3.1, we define the variance-adaptive half-width as
$w_{ij}^{(B)}(\alpha):=\sqrt{\frac{2s_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)},$ (3.6)
and the corresponding finite-sample confidence interval for the localized average $\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$, as
$\mathrm{CI}_{ij}^{(B)}(\alpha):=\Big[\tilde{P}_{ij}-w_{ij}^{(B)}(\alpha),\;\tilde{P}_{ij}+w_{ij}^{(B)}(\alpha)\Big]\cap[0,1].$ (3.7)
Remark 3.2 (Target of the Empirical Bernstein interval).
It is important to note that $\mathrm{CI}_{ij}^{(B)}(\alpha)$ is centered at the empirical neighborhood mean $\tilde{P}_{ij}$ and therefore provides finite-sample coverage for the localized average $\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$, not for the individual probability $P_{ij}$. The distinction reflects the fact that the Empirical Bernstein inequality controls only the stochastic fluctuation component $U_{ij}^{(-j)}$ and does not address the bias term $B_{ij}^{(-j)}$.
By construction, $\mathrm{CI}_{ij}^{(B)}(\alpha)$ captures the neighborhood average $\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$ with probability at least $1-\alpha$ for any finite $n$. To upgrade this to a coverage guarantee for the true edge probability $P_{ij}$, we must bound the deterministic bias. Under our structural assumptions, the underlying graphon is piecewise Lipschitz across a finite number of latent blocks. If a node’s latent position is too close to a block boundary, its empirical neighborhood may inadvertently capture nodes from an adjacent, structurally distinct block, compromising the smoothing procedure. To avoid these boundary effects, we restrict our entrywise inference to the regular interior of the graphon. For a fixed margin $\varepsilon>0$, let $I_{\varepsilon}$ denote the subset of the latent space strictly bounded away from all block partitions (defined formally in Section 4.3). We define the corresponding “good-edge set” as $\mathcal{E}_{\varepsilon}:=\{(i,j):i\neq j,\xi_{i}\in I_{\varepsilon},\xi_{j}\in I_{\varepsilon}\}$. As we rigorously establish in Section 4, for edges $(i,j)\in\mathcal{E}_{\varepsilon}$, the deterministic bias is uniformly bounded by a rate $r_{n}=(\log n/n)^{1/4}$.
Combining the finite-sample stochastic control with our deterministic bias bounds yields marginal coverage guarantees that are uniformly valid for any edge in the good set.
Theorem 3.3 (Marginal finite-sample inference for $P_{ij}$ on the good set).
Fix $\varepsilon>0$ and $\alpha\in(0,1)$. Let $\mathrm{CI}^{(B)}_{ij}(\alpha)$ be the empirical Bernstein interval defined in (3.7). Let $\mathcal{E}_{\varepsilon}$ denote the set of good edges defined in Section 4. Then there exist constants $C,c>0$ such that, with probability at least $1-n^{-c}$, the uniform bias bound
$\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}\left|P_{ij}-\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}\right|\leq C\,r_{n}$
holds, where $r_{n}=(\log n/n)^{1/4}$. Consequently, on the intersection of the bias and variance control events, we obtain
$\mathbb{P}\!\Big(P_{ij}\in\mathrm{CI}^{(B)}_{ij}(\alpha)\oplus[-Cr_{n},\;Cr_{n}]\Big)\geq 1-\alpha-n^{-c},$ (3.8)
where $\oplus$ denotes the Minkowski sum of sets ($A\oplus B=\{a+b\mid a\in A,b\in B\}$), which explicitly widens the variance-adaptive confidence interval on both sides by the deterministic bias margin.
Proof.
Let $E_{1}$ be the event that the empirical Bernstein bound holds, which has probability $\mathbb{P}(E_{1})\geq 1-\alpha$ by Lemma 3.1. Let $E_{2}$ be the event that the uniform bias bound holds over $\mathcal{E}_{\varepsilon}$, which has probability $\mathbb{P}(E_{2})\geq 1-n^{-c}$ by Corollary 4.13. On the intersection $E_{1}\cap E_{2}$, the distance from the estimator $\tilde{P}_{ij}$ to the neighborhood expected value is bounded by $w_{ij}^{(B)}(\alpha)$, and the distance from the neighborhood expected value to the true $P_{ij}$ is bounded by $Cr_{n}$. Applying the triangle inequality via the Minkowski sum, $P_{ij}$ must lie in $\mathrm{CI}^{(B)}_{ij}(\alpha)\oplus[-Cr_{n},\;Cr_{n}]$. By the union bound, $\mathbb{P}(E_{1}\cap E_{2})=1-\mathbb{P}(E_{1}^{c}\cup E_{2}^{c})\geq 1-\alpha-n^{-c}$. ∎
Remark 3.4 (Advantage of the Empirical Bound).
Unlike standard concentration inequalities that require substituting an estimated variance $\widehat{V}_{ij}$ for the true variance $V_{ij}$-a step that forces reliance on Slutsky’s Theorem and asymptotic normality-the Empirical Bernstein interval completely avoids the cost of studentization. The interval $\mathrm{CI}_{ij}^{(B)}(\alpha)$ automatically adapts to the empirical variance of the neighborhood while strictly preserving its $1-\alpha$ coverage probability for the localized average at any finite sample size.
3.2 Asymptotic Normality and Normal-Approximation Intervals
While the Empirical Bernstein bounds provide strictly finite-sample guarantees, it is also of fundamental theoretical and practical interest to establish the asymptotic distribution of the localized average. Under the non-degeneracy condition (Assumption 2.2), the conditional variance of the LOO summands is bounded uniformly away from zero. This allows us to establish a uniform Berry-Esseen bound for the standardized stochastic fluctuation.
Recall that the stochastic fluctuation term is defined as the average of the mean-zero noise over the leave-one-out neighborhood:
$U_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}).$ (3.9)
Conditional on the neighborhood $\mathcal{N}_{i}^{(-j)}$ and the latent probability matrix $P$, the edge indicators $A_{kj}$ are independent Bernoulli random variables. Therefore, the conditional variance of $U_{ij}^{(-j)}$ is strictly diagonal and is given by
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}).$ (3.10)
To facilitate the subsequent Berry-Esseen analysis, we define the unscaled variance of the neighborhood sum, $\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$, as
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.11)
so that $V_{ij}=\sigma_{ij}^{2}/h_{n}^{2}$. Under Assumption 2.2, there exists $\eta\in(0,1/2)$ such that
$\eta\leq P_{kj}\leq 1-\eta\qquad\text{for all }k,j.$
Hence
$\eta(1-\eta)\leq P_{kj}(1-P_{kj})\leq\frac{1}{4}.$
Since $|\mathcal{N}_{i}^{(-j)}|=h_{n}$, it follows that
$\eta(1-\eta)\,h_{n}\leq\sigma_{ij}^{2}\leq\frac{1}{4}\,h_{n}.$ (3.12)
In particular,
$\sigma_{ij}\geq\sqrt{\eta(1-\eta)}\,h_{n}^{1/2},\qquad\sigma_{ij}^{-1}\leq C\,h_{n}^{-1/2}$ (3.13)
for a constant $C>0$ independent of $(i,j,n)$. Consequently,
$V_{ij}=\frac{\sigma_{ij}^{2}}{h_{n}^{2}}\asymp h_{n}^{-1},$
and the conditional variance of $U_{ij}^{(-j)}$ does not degenerate.
Theorem 3.5 (Uniform Berry-Esseen bound).
Suppose Assumptions 2.1 and 2.2 hold, and let $h_{n}\to\infty$ as $n\to\infty$. Then there exists a constant $C>0$, depending only on $\eta$, independent of $i,j,$ and $n$, such that
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2}.$ (3.14)
Proof.
Fix $(i,j)$ and condition on $(\mathcal{N}_{i}^{(-j)},P)$. Define
$X_{k}:=A_{kj}-P_{kj},\qquad k\in\mathcal{N}_{i}^{(-j)}.$
By Proposition 2.7, the variables $\{X_{k}\}$ are independent, mean zero and bounded in [1,1][-1,1], but not necessarily identically distributed.
Let
$S:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k},\qquad\sigma_{ij}^{2}=\mathrm{Var}(S\mid\mathcal{N}_{i}^{(-j)},P).$
Define the third absolute moment sum
$\rho_{3}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathbb{E}\!\left[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P\right].$
Since $|X_{k}|\leq 1$, we have
$\rho_{3}\leq\sum_{k}\mathbb{E}[X_{k}^{2}\mid\mathcal{N}_{i}^{(-j)},P]=\sigma_{ij}^{2}.$
The Berry-Esseen inequality for independent, not necessarily identically distributed summands (see, e.g., Petrov (1975, Chapter V, Theorem 3)) states that there exists a universal constant $C_{0}>0$ such that
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\frac{\rho_{3}}{\sigma_{ij}^{3}}.$
Using $\rho_{3}\leq\sigma_{ij}^{2}$ yields
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\sigma_{ij}^{-1}.$
By (3.13), $\sigma_{ij}^{-1}\leq Ch_{n}^{-1/2}$, where $C$ is independent of $(i,j)$. Finally, since $U_{ij}^{(-j)}=h_{n}^{-1}S$ and $V_{ij}=h_{n}^{-2}\sigma_{ij}^{2}$, we have the identity
$\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}=\frac{S}{\sigma_{ij}},$
and the uniform bound (3.14) follows. ∎
Corollary 3.6 (Uniform asymptotic normality).
Under the conditions of Theorem 3.5, taking expectations over the conditioning variables yields
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|\longrightarrow 0\qquad\text{as }n\to\infty.$ (3.15)
To operationalize this normal approximation for confidence intervals, we must estimate the true conditional variance $V_{ij}$. We define the plug-in estimator
$\widehat{V}_{ij}:=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}),$ (3.16)
which replaces the unknown probabilities $P_{kj}$ with their smoothed estimates. As shown in our theoretical development, the error $|\widehat{V}_{ij}-V_{ij}|$ is bounded by the uniform estimation error of $\widehat{P}$.
Lemma 3.7 (Uniform plug-in control for $\widehat{V}_{ij}$).
For all $i\neq j$, the plug-in variance proxy satisfies
$\big|\widehat{V}_{ij}-V_{ij}\big|\leq\frac{1}{h_{n}}\,\sup_{1\leq a,b\leq n}\big|\widehat{P}_{ab}-P_{ab}\big|.$
Proof.
The map $\psi(x)=x(1-x)$ is $1$-Lipschitz on $[0,1]$. Hence for any $k$, $\big|\widehat{P}_{kj}(1-\widehat{P}_{kj})-P_{kj}(1-P_{kj})\big|\leq|\widehat{P}_{kj}-P_{kj}|\leq\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|$. Summing over the $h_{n}$ elements in $\mathcal{N}_{i}^{(-j)}$ and dividing by $h_{n}^{2}$ yields the result. ∎
Remark 3.8.
Our theoretical results establish rowwise mean-squared error control but do not prove uniform sup-norm consistency of $\widehat{P}$. Accordingly, plug-in variance intervals based on $\widehat{V}_{ij}$ should be interpreted as heuristic refinements rather than formally guaranteed procedures.
Remark 3.9.
The Berry-Esseen bound for the variance component holds uniformly over all $(i,j)$. However, coverage statements for $P_{ij}$ require bias control and therefore hold uniformly only over $(i,j)\in\mathcal{E}_{\varepsilon}$.
Remark 3.10.
The Empirical Bernstein bands are finite-sample and variance-adaptive but can be conservative. The normal bands are typically tighter but rely on asymptotic approximation and an explicit bias cushion; the two constructions provide complementary guarantees in practice.
Treating $U_{ij}^{(-j)}$ as approximately normal and explicitly incorporating the bias rate $r_{n}=(\log n/n)^{1/4}$, we define the normal-approximation half-width for a significance level $\alpha\in(0,1)$ as
$w_{ij}^{(N)}(\alpha):=z_{1-\alpha/2}\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}\Big(\frac{\log n}{n}\Big)^{1/4},$ (3.17)
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal distribution, and $c_{\mathrm{bias}}>0$ is a constant chosen to dominate the uniform bias term.
The resulting normal-approximation interval is given by
$\mathrm{CI}_{ij}^{(N)}(\alpha):=\Big[\tilde{P}_{ij}-w_{ij}^{(N)}(\alpha),\;\tilde{P}_{ij}+w_{ij}^{(N)}(\alpha)\Big]\cap[0,1].$ (3.18)
By Theorem 3.5, these intervals are asymptotically valid for the stochastic fluctuation uniformly over all pairs $(i,j)$. Furthermore, for edges $(i,j)\in\mathcal{E}_{\varepsilon}$ lying in the regular interior of the graphon, the explicit bias cushion ensures asymptotic coverage of the true edge probability $P_{ij}$. Under the canonical scaling $h_{n}\asymp\sqrt{n\log n}$, the stochastic fluctuation scale $h_{n}^{-1/2}\asymp(n\log n)^{-1/4}$ is slightly dominated by the bias scale $r_{n}=(\log n/n)^{1/4}$ by a logarithmic factor. Because the deterministic bias decays slower than the stochastic noise, this explicit bias cushion is strictly necessary for first-order valid inference.
Theorem 3.11 (Asymptotic normality under undersmoothing).
Suppose Assumptions 2.1, 2.2, and 4.7 hold. Let $(i,j)\in\mathcal{E}_{\varepsilon}$ and assume the bandwidth sequence satisfies
$h_{n}\to\infty,\qquad\sqrt{h_{n}}\Big(\frac{\log n}{n}\Big)^{1/4}\to 0.$ (3.19)
Then
$\frac{\tilde{P}_{ij}-P_{ij}}{\sqrt{V_{ij}}}\;\xrightarrow{d}\;N(0,1).$ (3.20)
Moreover, if the uniform bias bound of Corollary 4.13 holds, then
$\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{\tilde{P}_{ij}-P_{ij}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|\longrightarrow 0.$ (3.21)
Proof.
Recall the decomposition $\tilde{P}_{ij}-P_{ij}=U_{ij}^{(-j)}+B_{ij}^{(-j)}$. Therefore,
$\frac{\tilde{P}_{ij}-P_{ij}}{\sqrt{V_{ij}}}=\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}+\frac{B_{ij}^{(-j)}}{\sqrt{V_{ij}}}.$ (3.22)
By Theorem 3.5, $\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(U_{ij}^{(-j)}/\sqrt{V_{ij}}\leq x\mid\mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2}$, and hence $U_{ij}^{(-j)}/\sqrt{V_{ij}}\xrightarrow{d}N(0,1)$.
It remains to show that the bias term is negligible on the $\sqrt{V_{ij}}$ scale. By Corollary 4.13, for $(i,j)\in\mathcal{E}_{\varepsilon}$, $|B_{ij}^{(-j)}|\leq C(\frac{\log n}{n})^{1/4}$ with probability tending to one. By (3.12)-(3.13), $V_{ij}=\sigma_{ij}^{2}/h_{n}^{2}\asymp h_{n}^{-1}$, so that $1/\sqrt{V_{ij}}\asymp\sqrt{h_{n}}$. Hence
$\left|\frac{B_{ij}^{(-j)}}{\sqrt{V_{ij}}}\right|\lesssim\sqrt{h_{n}}\Big(\frac{\log n}{n}\Big)^{1/4}\;\longrightarrow\;0$
by the undersmoothing condition (3.19). Thus $B_{ij}^{(-j)}/\sqrt{V_{ij}}\xrightarrow{p}0$. Combining this with (3.22) and Slutsky’s theorem yields (3.20).
For the uniform statement, let $\epsilon_{n}:=C^{\prime}\sqrt{h_{n}}(\log n/n)^{1/4}$. By the undersmoothing condition, $\epsilon_{n}\to 0$. By Corollary 4.13, the event $E_{n}=\{\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}|B_{ij}^{(-j)}/\sqrt{V_{ij}}|\leq\epsilon_{n}\}$ holds with probability at least $1-n^{-c}$. For any $x\in\mathbb{R}$ and $(i,j)\in\mathcal{E}_{\varepsilon}$, we have
$\mathbb{P}\!\left(\frac{\tilde{P}_{ij}-P_{ij}}{\sqrt{V_{ij}}}\leq x\right)\leq\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x+\epsilon_{n}\right)+\mathbb{P}(E_{n}^{c}).$
Applying the uniform Berry-Esseen bound to the first term, it is bounded by $\Phi(x+\epsilon_{n})+\mathcal{O}(h_{n}^{-1/2})$. Because the standard normal density is bounded by $1/\sqrt{2\pi}$, $\Phi(x+\epsilon_{n})\leq\Phi(x)+\epsilon_{n}$. An analogous lower bound holds. Since $\epsilon_{n}\to 0$, $h_{n}^{-1/2}\to 0$, and $\mathbb{P}(E_{n}^{c})\to 0$ independently of $x$ and $(i,j)$, the uniform convergence (3.21) immediately follows. ∎
Remark 3.12 (Example of an undersmoothing regime).
A concrete bandwidth choice satisfying (3.19) is
$h_{n}\asymp\frac{\sqrt{n}}{\log n}.$
Under this choice, the bias becomes negligible on the central limit scale, so the full estimator $\tilde{P}_{ij}-P_{ij}$ is asymptotically normal. This comes at the price of sacrificing the canonical MSE-optimal balance $h_{n}\asymp\sqrt{n\log n}$.
Both the finite-sample intervals in Section 3.1 and the normal-approximation intervals in Section 3.2 critically rely on the assumption that the deterministic neighborhood bias $B_{ij}^{(-j)}$ is controlled by the rate $(\log n/n)^{1/4}$. In Section 4, we unpack the structural geometry of the LOO neighborhoods to rigorously establish this bias bound.
4 Theoretical Development: Geometry and Bias Control
The inferential guarantees in Section 3 rely critically on controlling the deterministic approximation error introduced by neighborhood smoothing. In this section, we establish the theoretical foundation for this bias control. We proceed by decomposing the estimation error, proving the stability of the two-hop topological distances under our leave-one-out (LOO) modification, and demonstrating that rowwise $L^{2}$ proximity on the reduced graph implies latent spatial closeness.
Our objective is entrywise inference for the finite-sample edge probability matrix $P=(P_{ij})_{1\leq i,j\leq n}$, where $P_{ij}=f(\xi_{i},\xi_{j})$. All probability statements in this section and the preceding ones are conditional on the realized latent positions $(\xi_{1},\dots,\xi_{n})$, equivalently on $P$. Thus, our results provide conditional finite-sample and asymptotic inference for $P_{ij}$ rather than for the graphon $f$ itself.
4.1 Stability of the Two-Hop Matrix Under Deletion
The classical neighborhood smoothing estimator constructs neighborhoods based on the normalized two-hop matrix $M=\frac{1}{n}A^{2}$. Our LOO procedure relies on $M^{(-j)}=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}$. To transfer the established risk bounds of the classical method to our LOO estimator, we must first show that deleting a single vertex does not significantly perturb these empirical two-hop distances.
To compare $M$ and $M^{(-j)}$, we embed the $(n-1)\times(n-1)$ matrix $M^{(-j)}$ into an $n\times n$ matrix $M^{(-j)}_{\mathrm{embed}}$ by setting $M^{(-j)}_{\mathrm{embed}}(i,\ell)=M^{(-j)}_{i\ell}$ for $i,\ell\neq j$, and $M^{(-j)}_{\mathrm{embed}}(i,\ell)=M_{i\ell}$ if $i=j$ or $\ell=j$.
Lemma 4.1 (Uniform two-hop stability under deletion of one vertex).
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of an undirected simple graph. Then, deterministically,
$\sup_{1\leq j\leq n}\ \|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\;\leq\;\frac{2}{n},$ (4.1)
where $\|B\|_{\max}:=\max_{1\leq i,\ell\leq n}|B_{i\ell}|$. In particular, for all $j\in[n]$ and all $i,\ell\in[n]\setminus\{j\}$,
$|M_{i\ell}-M^{(-j)}_{i\ell}|\leq\frac{2}{n}.$ (4.2)
Proof.
Fix $j\in[n]$ and $i,\ell\in[n]\setminus\{j\}$. Expanding the $(i,\ell)$ entry gives
$(A^{2})_{i\ell}=\sum_{m=1}^{n}A_{im}A_{m\ell},\qquad\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell},$
hence $(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}=A_{ij}A_{j\ell}$, which implies $\big|(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}\big|\leq 1$. Therefore,
[…]
$\leq\left|\frac{1}{n}(A^{2})_{i\ell}-\frac{1}{n}\big(A^{(-j)}\big)^{2}_{i\ell}\right|+\left|\frac{1}{n}-\frac{1}{n-1}\right|\big(A^{(-j)}\big)^{2}_{i\ell}$
$\leq\frac{1}{n}+\frac{1}{n(n-1)}\big(A^{(-j)}\big)^{2}_{i\ell}.$
Since $\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell}\leq n-1$, we obtain $\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{2}{n}$. By definition, the difference is zero when $i=j$ or $\ell=j$. Taking the supremum over $j$ yields the stated bound. ∎
Lemma 4.2 (Stability of expected two-hop matrices under deletion).
Define $M^{\star}=\mathbb{E}[M\mid P]$ and $M^{(-j)\star}=\mathbb{E}[M^{(-j)}\mid P]$. Using the same embedding strategy, there exists a universal constant $C>0$ such that for all $n$,
$\sup_{1\leq j\leq n}\big\|M^{(-j)}_{\mathrm{embed}}-M^{(-j)\star}_{\mathrm{embed}}\big\|_{\max}\;\leq\;\|M-M^{\star}\|_{\max}+\frac{C}{n}.$ (4.3)
Proof.
By the triangle inequality,
$\|M^{(-j)}_{\mathrm{embed}}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\|M^{(-j)}_{\mathrm{embed}}-M\|_{\max}+\|M-M^{\star}\|_{\max}+\|M^{\star}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}.$
By Lemma 4.1, the first term is bounded by $2/n$. For the third term, Jensen’s inequality implies $\|M^{\star}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\mathbb{E}\big[\|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\mid P\big]\leq 2/n$. Combining these bounds yields (4.3) with $C=4$. ∎
Lemma 4.3 (Uniform two-hop concentration).
Under Assumptions 2.1 and 2.2, there exist constants $C,c>0$ such that
$\mathbb{P}\!\left(\|M-M^{\star}\|_{\max}\leq C\sqrt{\frac{\log n}{n}}\ \middle|\ P\right)\geq 1-n^{-c}.$ (4.4)
Proof.
By definition, the entries of the two-hop matrix and its conditional expectation are given by $M_{i\ell}=\frac{1}{n}\sum_{m=1}^{n}A_{im}A_{m\ell}$ and $M_{i\ell}^{\star}=\frac{1}{n}\sum_{m=1}^{n}P_{im}P_{m\ell}$, respectively. We bound the maximum deviation by partitioning the entries into off-diagonal ($i\neq\ell$) and diagonal ($i=\ell$) cases.
Regarding the off-diagonal entries ($i\neq\ell$), fix $i\neq\ell$ and define $Y_{m}:=A_{im}A_{m\ell}$. Because the graph has no self-loops ($A_{ii}=A_{\ell\ell}=0$), $Y_{i}=Y_{\ell}=0$. Thus, the summation reduces to:
$M_{i\ell}=\frac{1}{n}\sum_{m\notin\{i,\ell\}}Y_{m}.$
Conditional on $P$, the edge indicators $A_{uv}$ are mutually independent. For any two distinct indices $m,m^{\prime}\notin\{i,\ell\}$, the edge sets $\{(i,m),(m,\ell)\}$ and $\{(i,m^{\prime}),(m^{\prime},\ell)\}$ are disjoint. Consequently, $\{Y_{m}\}_{m\notin\{i,\ell\}}$ is a sequence of n2n-2 conditionally independent random variables bounded in $[0,1]$, with $\mathbb{E}[Y_{m}\mid P]=P_{im}P_{m\ell}$.
Applying Hoeffding’s inequality to this sum yields, for any $t>0$,
$\mathbb{P}\!\left(\left|M_{i\ell}-M^{\star}_{i\ell}\right|\geq t\ \middle|\ P\right)\leq 2\exp\left(-2(n-2)t^{2}\right).$ (4.5)
Now consider the diagonal entries ($i=\ell$). For the diagonal entries, since $A_{im}\in\{0,1\}$, we have $A_{im}^{2}=A_{im}$. Therefore, $M_{ii}=\frac{1}{n}\sum_{m=1}^{n}A_{im}$. This is a sum of n1n-1 independent Bernoulli random variables (excluding $m=i$). A direct application of Hoeffding’s inequality yields:
$\mathbb{P}\!\left(\left|M_{ii}-M^{\star}_{ii}\right|\geq t\ \middle|\ P\right)\leq 2\exp\left(-2(n-1)t^{2}\right).$ (4.6)
Now we are ready to provide a uniform bound. Let $\mathcal{E}$ be the event that $\max_{i,\ell}|M_{i\ell}-M^{\star}_{i\ell}|\geq t$. Applying a union bound over all $n^{2}$ entries of $M$, and using the looser exponent $2(n-2)t^{2}$ from (4.5) to uniformly bound both cases, gives:
$\mathbb{P}(\mathcal{E}\mid P)\leq\sum_{i,\ell}\mathbb{P}\!\left(\left|M_{i\ell}-M^{\star}_{i\ell}\right|\geq t\ \middle|\ P\right)\leq 2n^{2}\exp\left(-2(n-2)t^{2}\right).$
Setting $t=C\sqrt{\frac{\log n}{n}}$ and assuming $n\geq 4$ (so $n-2\geq n/2$), the upper bound becomes:
$\mathbb{P}(\mathcal{E}\mid P)\leq 2n^{2}\exp\left(-2\Big(\frac{n}{2}\Big)\frac{C^{2}\log n}{n}\right)=2n^{2}\exp\left(-C^{2}\log n\right)=2n^{2-C^{2}}.$
Choosing $C>\sqrt{c+2}$ ensures that $2n^{2-C^{2}}\leq n^{-c}$ for all sufficiently large $n$, completing the proof. ∎
4.2 Transfer of ZLZ Rowwise Risk to LOO
While our primary focus is entrywise inference, it is practically important that the leave-one-out topological surgery does not degrade the global estimation accuracy.
With the stability of the two-hop matrix established, we can transfer the row-wise risk guarantees of the classical neighborhood smoothing estimator (Zhang et al., 2017) to our LOO construction. Note that the subsequent entrywise bias bounds in Section 4.3 are derived independently from first principles; the results in this subsection stand alone to guarantee that the LOO estimator rigorously retains the minimax optimal row-wise risk profile of the original estimator proposed by Zhang et al. (2017).
Proposition 4.4 (Theorem 1 of Zhang et al. (2017): Row-wise risk).
Work under the assumptions of Zhang et al. (2017) and choose the ZLZ quantile cutoff parameter $h_{\mathrm{ZLZ}}=C_{0}(\log n/n)^{1/2}$, so that the resulting neighborhood sizes satisfy $|N_{i}|\geq(n\log n)^{1/2}$. Let $N_{i}$ be the ZLZ neighborhood of node $i$ constructed from the classical two-hop matrix $M=A^{2}/n$, and define the estimator
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|N_{i}|}\sum_{k\in N_{i}}A_{kj}.$ (4.7)
Then there exist constants $C,c>0$ such that with probability at least $1-Cn^{-c}$, uniformly over all $i\in[n]$,
$\frac{1}{n}\sum_{j=1}^{n}\big(\widetilde{P}^{\mathrm{ZLZ}}_{ij}-P_{ij}\big)^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$ (4.8)
Remark 4.5 (Mapping to ZLZ Assumptions).
Theorem 1 of Zhang et al. (2017) requires the graphon to be piecewise Lipschitz continuous and the network to be strictly dense or satisfy a specific sparsity lower bound. Our Assumption 2.1 (piecewise bi-Lipschitz) strictly implies their piecewise Lipschitz/Hölder requirement, and our dense regime scaling (Assumption 2.2) comfortably satisfies their density conditions. Thus, their row-wise risk bounds are legitimately applicable to our underlying model.
Corollary 4.6 (Leave-one-out ZLZ row-wise risk and $(U_{ij},B_{ij})$ decomposition).
Fix $j\in[n]$. Let $\widetilde{P}$ denote the ZLZ neighborhood-smoothing estimator computed from the reduced matrix $A^{(-j)}$, with LOO neighborhoods $\mathcal{N}_{i}^{(-j)}$ defined exactly as in ZLZ but using the two-hop matrix $M^{(-j)}=(A^{(-j)})^{2}/(n-1)$. Recall the error decomposition $\Delta_{i\ell}=U^{(-j)}_{i\ell}+B^{(-j)}_{i\ell}$. Then for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\leq\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;+\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2}.$ (4.9)
Moreover, applying Theorem 1 of Zhang et al. (2017) to the (n1)(n-1)-node graph $A^{(-j)}$ yields, with probability at least $1-c_{1}n^{-c_{2}}$, uniformly for all $j\in[n]$ and all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (4.10)
In particular, identifying ZLZ’s notation $J_{1}^{(-j)}(i,\ell):=\big(U^{(-j)}_{i\ell}\big)^{2}$ and $J_{2}^{(-j)}(i,\ell):=\big(B^{(-j)}_{i\ell}\big)^{2}$, the bounds imply that the row-averages of the squared variance and bias components are bounded by the exact same rate:
$\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2},\qquad\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (4.11)
Proof.
The decomposition $\widetilde{P}_{i\ell}-P_{i\ell}=\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\bigl[(A_{k\ell}-P_{k\ell})+(P_{k\ell}-P_{i\ell})\bigr]$ is identical to ZLZ’s initial decomposition, and the inequality $(\widetilde{P}_{i\ell}-P_{i\ell})^{2}\leq 2(U^{(-j)}_{i\ell})^{2}+2(B^{(-j)}_{i\ell})^{2}$ follows directly via Cauchy-Schwarz.
Furthermore, ZLZ’s control of the two-hop quantities $\frac{1}{n}(A^{2})_{ab}-\frac{1}{n}(P^{2})_{ab}$ relies strictly on concentration inputs. As shown in Lemma 4.1 and Lemma 4.2, these exact same concentration bounds hold when $A$ is replaced by $A^{(-j)}$ and $n$ is replaced by n1n-1. Applying ZLZ Theorem 1 to the (n1)(n-1)-node graph verifies that the assumptions hold uniformly under the exact same graphon class. Because the concentration inputs underlying ZLZ Theorem 1 admit arbitrary polynomial decay $\mathcal{O}(n^{-c})$ by scaling the threshold constants, we may take a union bound over all $n$ leave-one-out graphs $A^{(-j)}$ while preserving a global high-probability event $1-c_{1}n^{-c_{2}}$. Noting that the rate $(\frac{\log(n-1)}{n-1})^{1/2}\asymp(\frac{\log n}{n})^{1/2}$, all subsequent ZLZ steps bounding the row-averages of $J_{1}$ and $J_{2}$ remain mathematically valid uniformly over all $j\in[n]$, preserving the convergence rate for the leave-one-out graph. ∎
4.3 Neighborhood Regularity and Latent Closeness
The row-wise risk transfer established in Corollary 4.6 ensures that the leave-one-out (LOO) neighborhood $\mathcal{N}_{i}^{(-j)}$ consists of nodes whose row profiles are close to that of node $i$ in an $L^{2}$ sense. To convert this topological proximity into a bound on the estimation bias, we must guarantee that row-wise $L^{2}$ closeness implies actual proximity in the latent space. This requires structural assumptions on the graphon $f$.
Assumption 4.7 (Piecewise bi-Lipschitz smoothness).
The graphon $f:[0,1]^{2}\to[0,1]$ admits a finite vertical partition $0=b_{0}<b_{1}<\cdots<b_{K}=1$ with blocks $I_{m}:=(b_{m},b_{m+1})$ such that:
  1. 1.
    (Within-block Lipschitz in the first argument.) For all $m$ and all $u,u^{\prime}\in I_{m}$, $x\in[0,1]$,
    $|f(u,x)-f(u^{\prime},x)|\leq L|u-u^{\prime}|.$
  2. 2.
    (Across-block $L^{2}$ separation.) There exist $\eta,\delta>0$ such that for any $u\in I_{m}$ and $u^{\prime}\in I_{m^{\prime}}$ with $m\neq m^{\prime}$,
    $\lambda\big(\{x\in[0,1]:|f(u,x)-f(u^{\prime},x)|\geq\eta\}\big)\geq\delta.$
  3. 3.
    (Within-block $L^{2}$ identifiability.) For any $\varepsilon>0$ with $b_{m}+\varepsilon<b_{m+1}-\varepsilon$, there exists $\kappa(\varepsilon)>0$ such that for all $u,u^{\prime}\in[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$,
    $\int_{0}^{1}(f(u,x)-f(u^{\prime},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|u-u^{\prime}|^{2}.$
Assumption 4.7 is standard in nonparametric graphon estimation when entrywise inference is desired. It ensures that the graphon does not contain flat, unidentifiable regions and that distinct blocks are macroscopic distance apart.
Remark 4.8 (Scope of Assumption 4.7).
This assumption excludes graphons with flat regions or non-identifiable row structure, and is tailored to settings where rowwise $L^{2}$ proximity reflects latent proximity within each block. In particular, it does not cover graphons with a growing number of blocks or highly irregular multiscale structure. The condition is standard in nonparametric graphon estimation when entrywise inference is desired, but we emphasize that it restricts the theory to graphons with fixed, well-separated smooth components.
Remark 4.9 (The Interior Good Set).
To avoid boundary effects near the partition points, we restrict our entrywise inference to the regular interior of the graphon. For a fixed margin $\varepsilon>0$ chosen strictly smaller than half the minimum block width (ensuring $2\varepsilon<\min_{m}(b_{m+1}-b_{m})$), define the disjoint block interior $I_{\varepsilon}:=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$. We define the random good-node set as $\mathcal{G}_{\varepsilon}:=\{i\in[n]:\xi_{i}\in I_{\varepsilon}\}$, and the corresponding good-edge set as $\mathcal{E}_{\varepsilon}:=\{(i,j):i\neq j,\,i\in\mathcal{G}_{\varepsilon},\,j\in\mathcal{G}_{\varepsilon}\}$. Because $\xi_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}[0,1]$, the size of the good-node set is a sum of independent indicators, $|\mathcal{G}_{\varepsilon}|=\sum_{i=1}^{n}\mathbf{1}_{\{\xi_{i}\in I_{\varepsilon}\}}\sim\mathrm{Binomial}(n,\lambda(I_{\varepsilon}))$. Therefore, the empirical fraction of good nodes concentrates naturally at the parametric rate: $|\mathcal{G}_{\varepsilon}|/n=\lambda(I_{\varepsilon})+\mathcal{O}_{\mathbb{P}}(n^{-1/2})$.
Next, we establish that the LOO neighborhoods satisfy a uniform row-wise regularity condition for all nodes in the good set.
Lemma 4.10 (Verification of regularity for LOO neighborhoods).
Work under Assumptions 2.1-2.2 and Assumption 4.7, and fix $\varepsilon>0$. Let $h_{n}\asymp\sqrt{n\log n}$ be the neighborhood size, and for each $(i,j)\in\mathcal{E}_{\varepsilon}$ let $\mathcal{N}_{i}^{(-j)}$ be the LOO neighborhood. Then there exist constants $C,c>0$ such that with probability at least $1-n^{-c}$ (conditional on $P$),
$\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}\;\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$ (4.12)
Proof.
Throughout the proof we work on the probability space of the latent positions $\xi_{1},\dots,\xi_{n}$ and the edges, and all probability statements are conditional on $P$. Fix $(i,j)\in E_{\varepsilon}$, so $\xi_{i},\xi_{j}\in I_{\varepsilon}$.
First, we provide an abundant supply of candidate nodes to ensure that the algorithm can construct a neighborhood of size $h_{n}$ using only nodes that are geometrically close to the target node $i$ in the latent space. Let $t_{n}:=a\sqrt{\frac{\log n}{n}}$ for a constant $a>0$ to be chosen later, and define the candidate set:
$S_{i}:=\Big\{k\in[n]\setminus\{i,j\}:\ \xi_{k}\in I_{\varepsilon}\text{ and }|\xi_{k}-\xi_{i}|\leq t_{n}\Big\}.$
Recall that the leave-one-out distance threshold $d_{i,(h_{n})}^{(-j)}$ is defined as the $h_{n}$-th order statistic of the distances $\{d^{(-j)}(i,k)\}_{k\neq i,j}$. On the event that $|S_{i}|\geq h_{n}$, the set $S_{i}$ alone provides at least $h_{n}$ valid nodes. By the definition of an order statistic, the $h_{n}$-th smallest distance in the entire graph can be no larger than the maximum distance among any specific subset of size $h_{n}$ or greater. Thus, if $|S_{i}|\geq h_{n}$, we rigorously have:
$d_{i,(h_{n})}^{(-j)}\leq\max_{k\in S_{i}}d^{(-j)}(i,k).$ (4.13)
We now show that this sufficient-capacity event holds with high probability uniformly over the good set. Because $\xi_{k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}[0,1]$ and the target node $\xi_{i}$ lies in the regular interior $I_{\varepsilon}$, its distance to any block boundary is at least $\varepsilon$. Because $t_{n}\to 0$, the interval $\xi_{i}-t_{n},\xi_{i}+t_{n}]$ falls entirely within $I_{\varepsilon}$ for all sufficiently large $n$.
Therefore, the probability that any individual node $k$ falls into $S_{i}$ is exactly the Lebesgue measure of this interval, which is $2t_{n}$. Consequently, the size of the candidate set $|S_{i}|$ is distributed as a Binomial random variable $\mathrm{Bin}(n-2,2t_{n})$. The expected size of this set is:
$\mathbb{E}[|S_{i}|]=(n-2)(2t_{n})\approx 2a\sqrt{n\log n}.$
Recall that the neighborhood size is chosen as $h_{n}\asymp\sqrt{n\log n}$. By choosing the constant $a$ sufficiently large, we can guarantee that $\mathbb{E}[|S_{i}|]\geq 2h_{n}$. Applying a standard multiplicative Chernoff bound to the lower tail of the Binomial distribution gives:
$\Pr\Big(|S_{i}|<h_{n}\Big)\leq\exp\left(-\frac{\mathbb{E}[|S_{i}|]}{8}\right)\leq\exp\left(-C^{\prime}a\sqrt{n\log n}\right).$
Because this tail probability decays exponentially in $\sqrt{n\log n}$, taking a union bound over all $i\in G_{\varepsilon}$ easily absorbs the linear penalty of at most $n$ nodes. Thus, for a sufficiently large choice of $a$, there exists a constant $c_{0}>0$ such that:
$\Pr\Big(|S_{i}|\geq h_{n}\text{ for all }i\in G_{\varepsilon}\Big)\geq 1-n^{-c_{0}}.$
We condition on this global high-probability event for the remainder of the proof.
Second, we show rowwise $L^{2}$ proximity for candidates in $S_{i}$. For any $k\in S_{i}$, Assumption 4.7 (i) implies for each $\ell$,
$|P_{i\ell}-P_{k\ell}|=|f(\xi_{i},\xi_{\ell})-f(\xi_{k},\xi_{\ell})|\leq L|\xi_{i}-\xi_{k}|\leq Lt_{n}.$
Therefore, deterministically (conditional on $P$),
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq L^{2}t_{n}^{2}=L^{2}a^{2}\Big(\frac{\log n}{n}\Big).$
In particular, for all sufficiently large $n$,
$\max_{k\in S_{i}}\ \frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq C_{1}\Big(\frac{\log n}{n}\Big),$ (4.14)
for $C_{1}:=L^{2}a^{2}$.
Third, we control the LOO distance via two-hop concentration. Recall the definition of the leave-one-out empirical distance:
[…]
We decompose each empirical two-hop entry into its conditional expectation $M^{(-j)\star}_{r\ell}=\mathbb{E}[M^{(-j)}_{r\ell}\mid P]$ and a stochastic deviation. By the triangle inequality, for any index $\ell\notin\{i,k,j\}$, we can bound the empirical difference by the population difference plus the individual noise terms:
[…]
$\leq\big|M^{(-j)\star}_{i\ell}-M^{(-j)\star}_{k\ell}\big|+2\|M^{(-j)}-M^{(-j)\star}\|_{\max}.$
By Lemma 4.3 and the embedding stability of Lemma 4.1, the concentration bounds transfer directly to the reduced submatrices. With probability at least $1-n^{-c}$,
$\sup_{1\leq j\leq n}\|M^{(-j)}-M^{(-j)\star}\|_{\max}\leq C^{\prime}\sqrt{\frac{\log n}{n}}.$
Taking the maximum over all valid indices $\ell$ and substituting this uniform high-probability bound yields
$d^{(-j)}(i,k)\leq\max_{\ell\notin\{i,k,j\}}\big|M^{(-j)\star}_{i\ell}-M^{(-j)\star}_{k\ell}\big|+C\sqrt{\frac{\log n}{n}},$
where the constant $C$ absorbs the factor of $2$. Now, $M^{(-j)\star}_{r\ell}=\frac{1}{n-1}\sum_{m\neq j}P_{rm}P_{m\ell}$. By the triangle inequality and the fact that probabilities are bounded by $P_{m\ell}\in[0,1]$, we have
$\big|M^{(-j)\star}_{i\ell}-M^{(-j)\star}_{k\ell}\big|\leq\frac{1}{n-1}\sum_{m\neq j}|P_{im}-P_{km}|P_{m\ell}\leq\frac{1}{n-1}\sum_{m\neq j}|P_{im}-P_{km}|.$
Applying the Cauchy-Schwarz inequality to the sum over $m$ yields
$\big|M^{(-j)\star}_{i\ell}-M^{(-j)\star}_{k\ell}\big|\leq\frac{1}{n-1}\left(\sum_{m\neq j}(P_{im}-P_{km})^{2}\right)^{1/2}\Big(\sum_{m\neq j}1^{2}\Big)^{1/2}=\left(\frac{1}{n-1}\sum_{m\neq j}(P_{im}-P_{km})^{2}\right)^{1/2}.$
Because adding the non-negative $j$-th term only increases the sum, and $\frac{1}{n-1}\leq\frac{2}{n}$ for all $n\geq 2$, we can bound this by the full empirical row-wise $L^{2}$ distance:
$\big|M^{(-j)\star}_{i\ell}-M^{(-j)\star}_{k\ell}\big|\leq\sqrt{2}\left(\frac{1}{n}\sum_{m=1}^{n}(P_{im}-P_{km})^{2}\right)^{1/2}.$
Substituting this upper bound back into the inequality for $d^{(-j)}(i,k)$, and relabeling the dummy index $m$ to $\ell$ to match the standard row-distance notation, yields:
$d^{(-j)}(i,k)\leq C\left(\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\right)^{1/2}+C\sqrt{\frac{\log n}{n}},$ (4.15)
where the constant $C$ absorbs the $\sqrt{2}$ factor.
Finally, we can conclude Lemma  4.10. Apply (4.15) to each $k\in N_{i}^{(-j)}$ and average:
$\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq C_{3}\left(\max_{k\in N_{i}^{(-j)}}d^{(-j)}(i,k)\right)^{2}+C_{3}\Big(\frac{\log n}{n}\Big).$
Use (4.13) to replace the max over $N_{i}^{(-j)}$ by the max over $S_{i}$. For any candidate $k\in S_{i}$, (4.14) fed into (4.15) yields $d^{(-j)}(i,k)\leq C_{4}(\log n/n)^{1/2}$. Therefore,
$\max_{k\in S_{i}}d^{(-j)}(i,k)\ \leq\ C_{4}\Big(\frac{\log n}{n}\Big)^{1/2}.$
Plugging back gives the uniform bound $C\Big(\frac{\log n}{n}\Big)^{1/2}$ over $(i,j)\in E_{\varepsilon}$. A union bound over the events in first three steps yields probability at least $1-n^{-c}$. ∎
Lemma 4.11 (Rowwise $L^{2}$ closeness implies latent closeness).
Assume Assumption 4.7 and let $r_{n}=(\log n/n)^{1/4}$. Let $\xi_{i},\xi_{k}\in I_{\varepsilon}$, and let $(M_{n})_{n\geq 1}$ be a sequence of positive real numbers. There exist constants $c_{0},c_{1}>0$ such that if $M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq c_{0}$ and
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2},$
then with probability at least $1-n^{-c}$ for all sufficiently large $n$, $\xi_{i}$ and $\xi_{k}$ lie in the same block, and their latent distance is bounded by $|\xi_{i}-\xi_{k}|\leq c_{1}\sqrt{M_{n}+1}\,r_{n}$.
Proof.
Let $f_{i}(x)=f(\xi_{i},x)$ and $f_{k}(x)=f(\xi_{k},x)$, and define
$D(\xi_{i},\xi_{k})=\int_{0}^{1}(f_{i}(x)-f_{k}(x))^{2}\,dx.$
Concentration of the empirical row distance. Conditional on $(\xi_{i},\xi_{k})$, the variables
$Z_{\ell}:=(P_{i\ell}-P_{k\ell})^{2}=(f(\xi_{i},\xi_{\ell})-f(\xi_{k},\xi_{\ell}))^{2}$
are independent and identically distributed for all $\ell\notin\{i,k\}$, bounded in $[0,1]$, and satisfy $\mathbb{E}[Z_{\ell}\mid\xi_{i},\xi_{k}]=D(\xi_{i},\xi_{k})$. The two exceptional terms ($\ell=i$ and $\ell=k$) contribute at most $2/n$ to the empirical average.
Applying Hoeffding’s inequality to the n2n-2 independent terms gives, for any $t>0$,
$\mathbb{P}\!\left(\left|\frac{1}{n-2}\sum_{\ell\notin\{i,k\}}Z_{\ell}-D(\xi_{i},\xi_{k})\right|\geq t\ \middle|\ \xi_{i},\xi_{k}\right)\leq 2\exp(-2(n-2)t^{2}).$
Adjusting for the $2/n$ deterministic offset from the exceptional indices, and taking $t=C^{\prime}(\log n/n)^{1/2}$, yields that with probability at least $1-n^{-c}$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\geq D(\xi_{i},\xi_{k})-C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (4.16)
If $\xi_{i},\xi_{k}$ lie in different blocks, Assumption 4.7(ii) gives $D(\xi_{i},\xi_{k})\geq\eta^{2}\delta$. Using (4.16), for large $n$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\geq\frac{\eta^{2}\delta}{2}.$
Choosing $c_{0}=\eta^{2}\delta/4$ makes this incompatible with the assumed upper bound. Thus $\xi_{i},\xi_{k}$ must lie in the same block.
On the event that $\xi_{i},\xi_{k}$ lie in the same block and belong to $I_{\varepsilon}$, Assumption 4.7(iii) implies
$D(\xi_{i},\xi_{k})\geq\kappa(\varepsilon)|\xi_{i}-\xi_{k}|^{2}.$
Combining this with (4.16) and the assumed upper bound gives
$\kappa(\varepsilon)|\xi_{i}-\xi_{k}|^{2}\leq\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}+C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq(M_{n}+C)\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Therefore,
$|\xi_{i}-\xi_{k}|\leq\sqrt{\frac{M_{n}+C}{\kappa(\varepsilon)}}\,\Bigl(\frac{\log n}{n}\Bigr)^{1/4}\leq c_{1}\sqrt{M_{n}+1}\,r_{n},$
by taking $c_{1}=\sqrt{\max(1,C)/\kappa(\varepsilon)}$. ∎
4.4 Entrywise Bias Bound
With the neighborhood regularity and latent closeness established, we can finally bound the deterministic bias component of the leave-one-out estimator. Recall from our error decomposition that the bias is given by $B_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}-P_{ij}$. Because the nodes in $\mathcal{N}_{i}^{(-j)}$ have been shown to be close to $i$ in the latent space, the piecewise Lipschitz continuity of the graphon restricts how much $P_{kj}$ can deviate from $P_{ij}$.
Lemma 4.12 (Bias bound for the LOO neighborhood smoother).
Assume Assumption 4.7 and let $r_{n}=(\log n/n)^{1/4}$. Suppose that the neighborhood regularity condition from Lemma 4.10 holds, meaning that with probability at least $1-n^{-c}$, for all $i\neq j$,
$\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C_{2}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (4.17)
Then for all $i\neq j$ with latent positions $\xi_{i},\xi_{j}\in I_{\varepsilon}$, the bias is bounded by
$|B^{(-j)}_{ij}|\leq Cr_{n},$ (4.18)
where $C>0$ depends only on the model constants and $\varepsilon$.
Proof.
Fix $i\neq j$ and let $N=\mathcal{N}_{i}^{(-j)}$. Define $D(\xi_{i},\xi_{k})=\int_{0}^{1}(f(\xi_{i},x)-f(\xi_{k},x))^{2}\,dx$.
First, conditional on $(\xi_{i},\xi_{k})$, the variables $(P_{i\ell}-P_{k\ell})^{2}$ are i.i.d. for $\ell\notin\{i,k\}$, bounded in $[0,1]$, and have mean $D(\xi_{i},\xi_{k})$. The exceptional indices $\ell\in\{i,k\}$ contribute an $O(1/n)$ deterministic offset to the average. Applying Hoeffding’s inequality to the n2n-2 independent variables, followed by a union bound over all $k\in\{1,\dots,n\}$ (which trivially covers the random subset $N$), gives that with probability at least $1-n^{-c}$,
$\sup_{k\in N}\left|\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}-D(\xi_{i},\xi_{k})\right|\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (4.19)
Combining (4.17) and (4.19),
$\frac{1}{|N|}\sum_{k\in N}D(\xi_{i},\xi_{k})\leq C^{\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (4.20)
Now we control the average latent distance between $\xi_{i}$ and its neighbors in $N$. Let
$\mathcal{C}_{i}=\{k\in N:\xi_{k}\text{ lies in a different block than }\xi_{i}\}.$
Assumption 4.7(ii) gives $D(\xi_{i},\xi_{k})\geq\eta^{2}\delta$ for $k\in\mathcal{C}_{i}$. Using (4.20),
$\frac{|\mathcal{C}_{i}|}{|N|}\eta^{2}\delta\leq C^{\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2},\qquad\Rightarrow\qquad\frac{|\mathcal{C}_{i}|}{|N|}\leq C^{\prime\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Decompose
$\frac{1}{|N|}\sum_{k\in N}|\xi_{i}-\xi_{k}|^{2}=\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|\xi_{i}-\xi_{k}|^{2}+\frac{1}{|N|}\sum_{k\in\mathcal{C}_{i}}|\xi_{i}-\xi_{k}|^{2}.$
The second term is bounded by $|\mathcal{C}_{i}|/|N|$ since latent positions are in $[0,1]$. For $k\in N\setminus\mathcal{C}_{i}$ (same block), Assumption 4.7(iii) gives $D(\xi_{i},\xi_{k})\geq\kappa(\varepsilon)|\xi_{i}-\xi_{k}|^{2}$, hence
$\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|\xi_{i}-\xi_{k}|^{2}\leq\frac{1}{\kappa(\varepsilon)}\cdot\frac{1}{|N|}\sum_{k\in N}D(\xi_{i},\xi_{k}).$
Using (4.20) and the bound on $|\mathcal{C}_{i}|/|N|$, we conclude that with probability at least $1-n^{-c}$,
$\frac{1}{|N|}\sum_{k\in N}|\xi_{i}-\xi_{k}|^{2}\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (4.21)
We split the bias into same-block and cross-block components:
$|B^{(-j)}_{ij}|=\left|\frac{1}{|N|}\sum_{k\in N}(P_{kj}-P_{ij})\right|\leq\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|P_{kj}-P_{ij}|+\frac{1}{|N|}\sum_{k\in\mathcal{C}_{i}}|P_{kj}-P_{ij}|.$
For the cross-block neighbors $k\in\mathcal{C}_{i}$, we use the trivial probability bound $|P_{kj}-P_{ij}|\leq 1$. Thus, the second term is bounded by $\frac{|\mathcal{C}_{i}|}{|N|}\leq C^{\prime\prime}(\frac{\log n}{n})^{1/2}$.
For the same-block neighbors $k\in N\setminus\mathcal{C}_{i}$, Lipschitz continuity in the first argument gives $|P_{kj}-P_{ij}|\leq L|\xi_{k}-\xi_{i}|$. By Cauchy-Schwarz and (4.21), the first term is bounded by:
$\frac{L}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|\xi_{k}-\xi_{i}|\leq L\left(\frac{1}{|N|}\sum_{k\in N}|\xi_{i}-\xi_{k}|^{2}\right)^{1/2}\leq Cr_{n},$
where $r_{n}=(\log n/n)^{1/4}$. Since $(\frac{\log n}{n})^{1/2}\ll(\frac{\log n}{n})^{1/4}$ for large $n$, the cross-block term is strictly dominated and absorbed into the primary rate, yielding $|B^{(-j)}_{ij}|\leq C^{\prime}r_{n}$. ∎
Corollary 4.13 (Uniform bias bound on the good set).
Fix $\varepsilon>0$ satisfying the interior feasibility condition. Under the assumptions of Lemma 4.12, there exists a constant $C(\varepsilon,C_{2})>0$ such that, with probability at least $1-c_{1}n^{-c_{2}}$,
$\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}|B_{ij}^{(-j)}|\leq C(\varepsilon,C_{2})\,r_{n}.$ (4.22)
This uniform bias control over the good edge set $\mathcal{E}_{\varepsilon}$ is exactly the bound required to establish the coverage guarantees of the finite-sample and normal-approximation confidence intervals presented in Section 3.
Remark 4.14 (Canonical scaling versus undersmoothing).
The choice $h_{n}\asymp\sqrt{n\log n}$ balances the stochastic fluctuation scale and the bias scale, and is therefore natural from the perspective of mean-squared error. However, this choice does not make the bias negligible on the central limit scale. If the goal is asymptotic normality of the full estimator $\tilde{P}_{ij}-P_{ij}$, one may instead undersmooth by choosing $h_{n}$ so that
$\sqrt{h_{n}}\Big(\frac{\log n}{n}\Big)^{1/4}\to 0.$
To formally establish the bias-variance trade-off of our leave-one-out construction, we derive a rigorous conditional risk bound. Because the LOO construction strictly decouples the target column from the neighborhood formation, the cross-term vanishes exactly, yielding a pristine bound on the conditional Mean Squared Error.
Theorem 4.15 (Conditional Mean Squared Error Bound for the LOO Estimator).
Conditional on the latent probability matrix $P$ and the leave-one-out neighborhood $\mathcal{N}_{i}^{(-j)}$ of size $h_{n}$, the mean squared error of the LOO estimator satisfies the bound:
$\mathbb{E}\!\Big[(\tilde{P}_{ij}-P_{ij})^{2}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big]\leq\frac{1}{4h_{n}}+\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij})^{2}.$ (4.23)
Furthermore, bounding the localized approximation error by the maximum discrepancy in the neighborhood yields:
$\mathbb{E}\!\Big[(\tilde{P}_{ij}-P_{ij})^{2}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big]\leq\frac{1}{4h_{n}}+\max_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij})^{2}.$ (4.24)
Proof.
We begin with the fundamental decomposition of the estimation error into its stochastic fluctuation and deterministic bias:
$\tilde{P}_{ij}-P_{ij}=U_{ij}^{(-j)}+B_{ij}^{(-j)},$
where
$U_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj})\qquad\text{and}\qquad B_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}).$
Squaring the error and taking the conditional expectation yields:
$\mathbb{E}\!\Big[(\tilde{P}_{ij}-P_{ij})^{2}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big]=\mathbb{E}\!\Big[\big(U_{ij}^{(-j)}\big)^{2}\ \Big|\ \cdot\Big]+2\mathbb{E}\!\Big[U_{ij}^{(-j)}B_{ij}^{(-j)}\ \Big|\ \cdot\Big]+\mathbb{E}\!\Big[\big(B_{ij}^{(-j)}\big)^{2}\ \Big|\ \cdot\Big].$ (4.25)
By the strict geometric independence of the LOO construction, the target column $j$ is never queried during the formation of $\mathcal{N}_{i}^{(-j)}$. Consequently, conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the bias $B_{ij}^{(-j)}$ is a fixed deterministic constant. Because the noise terms $(A_{kj}-P_{kj})$ are conditionally independent and strictly mean-zero, the cross-term vanishes identically:
$\mathbb{E}\!\Big[U_{ij}^{(-j)}B_{ij}^{(-j)}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big]=B_{ij}^{(-j)}\mathbb{E}\!\Big[U_{ij}^{(-j)}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big]=0.$
The first term in (4.25) is exactly the conditional variance $V_{ij}$. Because $P_{kj}\in[0,1]$, the variance of each Bernoulli summand is bounded by $1/4$. Therefore:
$\mathbb{E}\!\Big[\big(U_{ij}^{(-j)}\big)^{2}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big]=V_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj})\leq\frac{h_{n}(1/4)}{h_{n}^{2}}=\frac{1}{4h_{n}}.$
For the third term, since $B_{ij}^{(-j)}$ is deterministic conditionally, its expectation is simply its squared value. Applying Jensen’s inequality (or equivalently, the Cauchy-Schwarz inequality) to the convex square function over the finite sum yields:
$\big(B_{ij}^{(-j)}\big)^{2}=\left(\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij})\right)^{2}\leq\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij})^{2}.$
Summing the variance bound and the bias bound completes the proof of (4.23). Bounding the average of the squared errors by their maximum element immediately yields (4.24). ∎
5 Honest tuning via cross-validation
The leave-one-out (LOO) construction admits a simple and honest cross-validation (CV) scheme for selecting the neighborhood size $h_{n}$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held-out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h_{n})$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h_{n})$. Consequently, the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard concentration and normal approximation results, enabling valid CV-based tuning of $h_{n}$.
We use the term “honest” to denote that no held-out edge $A_{ij}$ is used in constructing its predictor $\tilde{P}_{ij}^{(-j)}$. This eliminates double use of data but does not imply risk optimality of the selected bandwidth.
5.1 Cross-validation
Fix a row index $i\in[n]$. Let $\mathcal{G}$ be a finite grid of candidate neighborhood sizes,
$\mathcal{G}\subset\{1,2,\dots,n-2\},$
typically chosen around the theoretical scale $h_{n}\asymp\sqrt{n\log n}$. For each $h_{n}\in\mathcal{G}$ and each held-out column $j\neq i$, form the reduced adjacency matrix $A^{(-j)}$ by deleting the $j$-th row and column. This ensures that the neighborhood $\mathcal{N}_{i}^{(-j)}(h_{n})$ is constructed without using any information from the held-out target column $j$. Compute the corresponding leave-one-out predictor
$\tilde{P}_{ij}^{(-j)}(h_{n}):=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h_{n})}A_{kj},\qquad j\neq i.$
Define the LOO cross-validation score
$\widehat{R}_{i}(h)=\frac{1}{n-1}\sum_{j\neq i}(A_{ij}-\tilde{P}_{ij}^{(-j)}(h))^{2}.$
Because the neighborhood $\mathcal{N}_{i}^{(-j)}(h)$ is constructed from $A^{(-j)}$ (which omits all edges incident to vertex $j$) and $i\notin\mathcal{N}_{i}^{(-j)}(h)$ by construction, the held-out edge $A_{ij}$ is conditionally independent of the predictor $\tilde{P}_{ij}^{(-j)}(h)$ given $P$. Consequently, $\widehat{R}_{i}(h)$ is (conditionally) unbiased for the corresponding prediction risk. A full risk inequality analysis of this tuning rule is not included in this version.
The corresponding prediction risk is the conditional expectation of the squared error. Because the held-out target $A_{ij}$ is conditionally independent of $\mathcal{N}_{i}^{(-j)}(h_{n})$ and the predictor $\tilde{P}_{ij}^{(-j)}(h_{n})$ given $P$, the cross-term vanishes, yielding:
$\mathbb{E}\!\left[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h_{n})\bigr)^{2}\,\middle|\,\mathcal{N}_{i}^{(-j)}(h_{n}),P\right]=\bigl(P_{ij}-\tilde{P}_{ij}^{(-j)}(h_{n})\bigr)^{2}+P_{ij}(1-P_{ij}).$ (5.1)
Because the irreducible Bernoulli variance $P_{ij}(1-P_{ij})$ does not depend on the bandwidth $h_{n}$, the empirical LOO score $\widehat{R}_{i}(h_{n})$ is an unbiased surrogate for minimizing the true conditional mean-squared error. We therefore define the theoretical prediction risk to be optimized as $R_{i}(h_{n}):=\frac{1}{n-1}\sum_{j\neq i}\bigl(P_{ij}-\tilde{P}_{ij}^{(-j)}(h_{n})\bigr)^{2}$. We then select any minimizer of the empirical score over the grid,
$\hat{h}_{n}\in\arg\min_{h_{n}\in\mathcal{G}}\widehat{R}_{i}(h_{n}).$
6 Simulation Study
We conducted a simulation study to evaluate the finite-sample performance of the proposed leave-one-out (LOO) neighborhood smoother and compare it with the classical (ZLZ) neighborhood smoother. All computations were implemented in Python using NumPy with a fixed random seed.
Algorithm 1 Simulation Procedure for LOO Smoothing and Confidence Intervals
1:Sample size $n$, neighborhood size $h_{n}$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
2:Estimated matrices $\widehat{P}_{\mathrm{LOO}},\widehat{P}_{\mathrm{class}}$ and empirical coverage rates
3:Draw $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$ and set $P_{ij}=f(\xi_{i},\xi_{j})$
4:Sample $A_{ij}\mid P_{ij}\sim\mathrm{Bernoulli}(P_{ij})$ for $i<j$ and symmetrize
5:For each $j$ form $A^{(-j)}$, compute distances $d_{ik}^{(-j)}$ via (2.2)
6:Build LOO neighborhoods $\mathcal{N}_{i}^{(-j)}$ of size $h_{n}$ and classical neighborhoods $\mathcal{N}_{i}$ from full AA
7:Compute $\tilde{P}_{ij}$ and $\widehat{P}_{ij}$ via (2.4) for both LOO and classical smoothers
8:Estimate $\widehat{V}_{ij}$ and construct intervals $\mathrm{CI}_{ij}^{(B)}(\alpha)$, $\mathrm{CI}_{ij}^{(N)}(\alpha)$
9:Evaluate empirical coverage $\widehat{\mathrm{Cov}}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}$ and report MSEs
To comprehensively evaluate the proposed uncertainty quantification, we track three primary metrics across the varying graphon topologies: the mean squared error (MSE) for point estimation, the empirical coverage rate, and the average confidence interval width. While empirical coverage verifies the theoretical validity of our bounds, tracking the average interval width explicitly quantifies the inferential cost. Specifically, comparing the average bounds allows us to explicitly demonstrate the trade-off between the conservative, strictly non-asymptotic Empirical Bernstein intervals and the sharper, asymptotically valid Normal approximations. The results of this numerical study are summarized in Table 2.
6.1 Setup
We considered five representative graphon models capturing a range of smoothness and structural complexity:
Table 1: Graphon Model Definitions
Model Definition
Smooth $0.5+0.3\sin(\pi u)\sin(\pi v)$
Block 2-block SBM ($p_{\text{in}}=0.7,p_{\text{out}}=0.3$)
Wiggly $0.5+0.25\sin(4\pi uv)$
Rank-one $(u+v)/2$
Spiky $0.2+0.8\,\mathbbm{1}\{|u-0.5|<0.1,\ |v-0.5|<0.1\}$
Remark 6.1.
The empirical coverage levels reported in Table 2 are higher than typical for $n=500$. This reflects the favorable latent geometry and relatively generous bias cushion used in our simulation design. In more challenging settings, we expect coverage to be closer to the nominal level.
For each graphon, we generated an undirected graph with $n=500$ nodes by sampling latent positions $\xi_{i}\sim\mathrm{Unif}[0,1]$ and edges $A_{ij}\sim\mathrm{Bernoulli}(f(\xi_{i},\xi_{j}))$ independently for $i<j$. The neighborhood size and trimming quantile were fixed at
$h_{n}=\lfloor 1.5\sqrt{n\log n}\rfloor.$
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column and row $j$ removed, thereby restoring conditional independence between the neighborhood and the target edge. To evaluate point estimation accuracy, both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$ for the MSE calculations. However, to strictly adhere to our theoretical independence guarantees, the confidence intervals were constructed and evaluated using the fundamental one-sided predictor $\tilde{P}_{ij}$.
6.2 Evaluation metrics
We report two performance measures. For node $i=0$,
$\mathrm{MSE}_{0}=\frac{1}{n}\sum_{j=1}^{n}(\widehat{P}_{0j}-P_{0j})^{2}.$
For the LOO estimator, we consider:
  • Empirical Bernstein intervals:
    $w_{ij}^{(B)}=\sqrt{\frac{2s_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}.$
  • Normal intervals:
    $w_{ij}^{(N)}=1.96\sqrt{\widehat{V}_{ij}}+c_{\mathrm{bias}}(\log n/n)^{1/4},\qquad c_{\mathrm{bias}}=0.1,$
where the plug-in variance proxy is
$\widehat{V}_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}).$
Intervals are truncated to $[0,1]$. Coverage is computed as
$\mathrm{Coverage}=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbbm{1}\{P_{ij}\in\mathrm{CI}_{ij}\}.$
6.3 Results
Table 2 summarizes the results for all five graphons. The LOO smoother achieves a comparable MSE to the classical ZLZ smoother. Coverage for both confidence interval constructions is close to the nominal $95\%$ level, with the Empirical Bernstein intervals slightly more conservative.
Table 2: Finite-sample performance of the LOO estimator across various graphon topologies ($n=500$). We report the mean squared error (MSE) for both the asymmetric LOO and classical symmetrized estimators for a fixed row. Coverage and average interval widths are compared between the strictly non-asymptotic Empirical Bernstein bounds and the asymptotic Normal approximation.
Point Estimation (MSE) Uncertainty Quantification (LOO)
Graphon LOO Classical Cov. (EB) Width (EB) Cov. (Normal) Width (Normal)
Smooth $0.0019$ $0.0020$ $1.000$ $0.554$ $0.980$ $0.272$
Block $0.0014$ $0.0020$ $1.000$ $0.539$ $0.985$ $0.263$
Wiggle $0.0061$ $0.0060$ $0.999$ $0.557$ $0.919$ $0.272$
Rank-1 $0.0015$ $0.0016$ $1.000$ $0.530$ $0.983$ $0.264$
Spiky $0.0005$ $0.0009$ $1.000$ $0.440$ $0.989$ $0.230$
The results in Table 2 provide strong empirical validation for both the estimation efficiency and the inferential validity of the leave-one-out framework.
First, regarding point estimation, the asymmetric LOO predictor $\tilde{P}_{ij}$ demonstrates highly competitive mean squared error compared to the classical symmetrized estimator $\widehat{P}_{ij}$. In fact, the LOO estimator matches or slightly improves upon the classical MSE in four of the five tested topologies. This confirms our theoretical assertion that strictly excluding the target column during neighborhood formation incurs a negligible finite-sample efficiency penalty, preserving the minimax optimal rate while unlocking conditional independence.
Second, the coverage results clearly illustrate the theoretical trade-offs between our two inferential approaches. The Empirical Bernstein intervals, grounded in strictly non-asymptotic bounds, achieve absolute coverage (near $1.000$) across all topologies. However, this theoretical guarantee comes at the cost of practical sharpness, yielding intervals that span the majority of the $[0,1]$ support. In contrast, the asymptotic intervals constructed via our Central Limit Theorem cut the average interval widths by more than half (averaging roughly $0.25$). Crucially, these Normal approximations successfully maintain the nominal $0.95$ coverage rate on the smooth, block, rank-1, and spiky graphons at sample size of $n=500$.
Finally, the wiggle graphon provides an instructive boundary case regarding finite-sample deterministic bias. While the Normal approximation achieves $0.919$ coverage, the under-coverage relative to the $0.95$ nominal target is a direct consequence of the graphon’s high-frequency gradients ($\sin(4\pi uv)$). Such highly variable local geometries inflate the deterministic approximation error at finite $n$.
7 Discussion and Extensions
This paper develops a principled framework for inference on individual edge probabilities in graphon models based on a simple leave-one-out modification of neighborhood smoothing. By constructing LOO neighborhoods that do not depend on the column being predicted, we restore conditional independence of the centered edge variables and thereby make classical concentration and normal-approximation tools available. The resulting methodology delivers entrywise confidence intervals and an honest cross-validation scheme for data-driven tuning, while retaining the optimal row-wise mean-squared error rate known for classical neighborhood smoothing.
The LOO construction is conceptually simple and quite general, which makes it a promising building block for extensions beyond the basic undirected, unweighted, static setting considered here. We outline several directions for future work.
  1. 1.
    Directed and weighted graphons. For directed networks, one could replace the Zhang et al. (2017) distance by an asymmetric distance that incorporates both in and out neighborhood information, and then apply the same LOO decoupling on the relevant columns. For weighted networks, a similar strategy could be pursued using appropriate moment-based distances built from $A$ and $A^{2}$, with boundedness or tail assumptions tailored to the edge-weight distribution.
  2. 2.
    Extension to sparse regimes. For models of the form $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$, the effective variance of $U_{ij}^{(-j)}$ shrinks with $\rho_{n}$. Normal approximation would require renormalization by $\sqrt{h_{n}\rho_{n}}$, and empirical Bernstein bounds must account for rare-event behavior. While the leave-one-out decoupling remains valid, the concentration and Berry-Esseen analysis would require substantially different arguments.
  3. 3.
    Degree correction and heterogeneous sparsity. Many real networks exhibit highly heterogeneous degree distributions. Extending the present analysis to degree corrected graphon or blockmodel formulations would require incorporating node-specific scale factors into both the distance and the bias analysis. The LOO idea remains applicable, but the variance proxy $V_{ij}$ and the bias bounds would need to be adapted to handle varying effective sample sizes across nodes.
  4. 4.
    Adaptive and simultaneous confidence bands. While we have focused on pointwise confidence intervals for individual edge probabilities, the same machinery can be used to construct simultaneous bands over subsets of rows or columns, for example by combining the Berry-Esseen bounds with high-dimensional Gaussian approximation results. It would also be natural to let the bias cushion constant $c_{\mathrm{bias}}$ be selected in a data-driven way, for instance by cross-validation or stability criteria, in order to balance coverage and tightness of the intervals.
  5. 5.
    Applications to real networks. Finally, it would be valuable to apply the proposed methodology to real data sets, such as social, citation, or biological networks, in order to assess its practical performance and interpretability. In particular, the ability to attach uncertainty quantification to individual edge probabilities or to row-wise summaries could provide new insights in applied settings where only a single large network is observed.
Identifiability limitations.
The within-block $L^{2}$ identifiability condition in Assumption 4.7 is structurally stronger than conditions used solely for MSE analysis. It is required here to translate rowwise $L^{2}$ closeness into entrywise bias control. Extending the inference theory to settings with weaker or purely combinatorial identifiability (e.g., high-$K$ block models) remains an open problem.
The main limitation of the current approach is the requirement of piecewise Lipschitz smoothness for the underlying graphon. Extensions to rougher kernels, highly discontinuous structures, or models with sharp community boundaries would likely require modified distance metrics and a more delicate bias analysis. Addressing these challenges, while retaining tractable inference procedures, remains an important open problem.